A Complete Bibliography of ACM Transactions on Computational Logic

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

18 October 2023 Version 1.77

Title word cross-reference

1 [CC19]. 2 [MMZ21]. _{mod} [BS09]. α [Bas21]. αβ [Bas21]. λ [KS13b, MP03]. λδ [Gui09]. Λμ [Sau10, Van18a]. μ [GLSG15, MM07]. n! [AI03]. O [ACOW22]. O(m log n) [GJKW17]. ω [BBS07, CHH09]. p [Kra21]. π [TM10]. qMμ [MM07]. × [HK21].

-Calculus [GLSG15, KS13b, MM07, TM10, MP03, Sau10]. -Minimal [ACOW22]. -regular [BBS07, CHH09]. -Relations [Bas21]. -Renaming [Bas21]. -Safe [CC19]. -Sequents [MMZ21]. -time [Kra21].

2003 [Kol05]. 2008 [KM10]. 2SAT [Sub04].

3-valued [SG07].

abduction [LS07, LY07]. Abductive [KKLJU14, MSS14, SDSS13, ACG^{+}08]. abelian [MOG05]. Abilities [BJP19]. ability [GS09]. abnormality [CEG05]. Abstract [BG03a, BG08a, BD07b, FFP15, HCGT18, MZ17, TZ02, TZ04, GRS05, Gur00, MM02]. abstract-state [Gur00]. Abstraction [CGWW15, CV10, LB07, LLM^{+}07, SG07]. abstraction-refinement [CV10, SG07]. Abstractions [GRV21, NNN11, YRSW07]. Access [BBGS20]. accessing [ACW12]. Accumulative [BCHK14]. Ackermann [LOW16]. Ackermann-Complete [LOW16]. Acquisition [DN21]. ACT [GF18]. ACT-R [GF18]. Action [Kuz21, PS17, BG02]. actions [DST13, ILNR09]. Active [BCF21]. Activities [AG19]. Actor [AFMGMM16]. Actor-Based [AFMGMM16]. Actual [Bas21]. Acyclic [EEH17a]. Adding [DLM21, PZ20]. Addition [CLLM19]. additive [HV05]. adjoint [DST13]. Adjoints [GCK^{+}22]. Admissibility [CFMP19, WZ08]. Affine [AR02, Les18, Try16, CM06]. Again [CLRT21]. against [BBS07]. agent [ACG^{+}08, DNS00, Lia05]. Agents [BJP19, DL04, DKS06]. aggregate [CSN05]. aggregates [Fer11]. Algebra [AO19, BBH^{+}14, DST13, FS15b, NSST18, AE09, BP11, BG02, DMS06, Koz00, YFDJ09]. Algebra-coalgebra [BBH^{+}14]. Algebraic [CMMZ23, Bou09, TZ02, VGD06, VGD07]. Algebras [ACMU21, CFP^{+}19, RV01, TZ04]. Algorithm [GJKW17, Rab18, BK10, BBH^{+}14, CM06]. Algorithmic [ACW12, CLSZ15, EEH17a]. Algorithmically [CP22]. Algorithms [Kra21, Lib23, BG03a, BG06, BG07a, BG07b, BG08a, BG11a, Gur00]. Alphabets [Tan13, NSV04]. Alternating [BBDM23, BFF^{+}21, CDG15, GKR18, JL11, LW08, KV01, Laz11]. Alternating-Time [CDG15, GKR18]. Alternation [BKV15, KS17]. Alternative [WWT15]. Amalgamations [AG23]. among [CSN05]. Analysis [AFMGMM16, ACJ^{+}18, BF13, KRW12, NV19, Tak23, AS02, ACW12, JK09, MP09]. Analytic [GGOP09]. Analyticity [LZ19]. Annotated [MSS14, SPSS11, SSS12, SBSM13, URS10]. Announcements [BvDG22]. Answer [AZZ14, BJN20, DSTW13, GS13b, OSTW21, WEZL23, WWSL16, EFW07, HVV08, Sak05, SI08, SBTM06]. Answer-Set [BJN20, WWSL16, OSTW21]. Answering [AB20, Fon15, LMTV19, SDSS13, CDL08]. answers [LY07]. Anti [CK20].

Anti-unification [CK20]. Application [BCF21, CLCZ10]. Applications [BHM15, CC19, CG18, DHS16, DST13, IS04, VH05]. Applying [BJN20]. Approach [Bd12, CLSZ15, DSTW13, DDDW22, HS15, LDJ^{+}18, LOS19, MW15, MZ17, MBRS20, DPR08, EFL^{+}04, LLM^{+}07, LF01, SYY^{+}03]. Approximate [BHM15, SSS12, Rat04]. approximating [AD14]. Approximation [BCF21]. Arbitrary [FP17]. Architecture [GF18]. Argumentation [CES14, FFP15, MMS23]. Arithmetic [BT23, CLZ23, CGI^{+}18, KHST12, KPS22, Sch05, VM09, BB14, BM10, BJW05, CT06, Kla08]. arithmetical [HLT07]. array [ACW12]. array-accessing [ACW12]. Arrays [GGKN23, MNPS20]. Aspect [Tak23]. Aspects [Les18]. Assertional [LL15]. Assignments [EG22]. Assumptions [KPS22]. Asymptotic [VZ19]. Asynchronous [BvDG22, GS13a]. ATL [CDG15]. atoms [DFO06]. Attacker [BCEO20]. Autoepistemic [CMVT12, DEPT11, KR02]. Automata [BFF^{+}21, BLM^{+}23, BGMZ14, BCOS21, CHS14, CHS15, CHO17, CCD15, CLLM19, FMS16, FKL15, HMOS17, HK15a, HSTW20, MBRS20, QS19, Sei18, Tan13, BDR08, BBD03, DL09, JL11, Kla08, KV01, LW08, Laz11, Tan14, Tri09]. Automated [ALGN18, CCCK16, CKK17, Gul18, SKGST09]. Automatic [AR04, KRS05, LDJ^{+}18]. Axiomatic [DDDW22, SH07]. Axiomatisation [ACF^{+}22]. axiomatizability [FD14]. Axiomatizations [BMS13]. axiomatizing [BG02]. axioms [BGL12, IS06].

Back [CLRT21, GHO02]. Backdoors [FS15a]. Background [BBT^{+}21]. bad [BMS^{+}11]. Barbs [BGM14]. base [AFIL09, DEPT11]. Based [AFMGMM16, BCEO20, BF13, BH23, CES14, FKL15, HLM13, LNH21, MW15, MZ17, Try16, WWT15, AR04, ABRS09, BG03b, ET06, GGOS09, GS09, LLM^{+}07, LSS04, MMS23, MBRS20, MV04, MN12, Rei01, SG07, Zha06]. Bases [AKS13, GSM^{+}22, MBN12, TSH15, AAD11, EFST05, Lib00a]. Basic [Kon21, Mie04, BM08]. BDD [BH23]. BDD-Based [BH23]. Be [ALN16]. behavior [NNN11]. Behavioral [CR15, BG11a]. Behaviors [BKM^{+}23]. behind [LM02]. Being [APNT22]. Belief [Ara23, BZWS18, CD14, CPRW16, DSTW13, FL20, GSM^{+}22, Lib16, PKW12, Lia05]. Beliefs [SFL17]. Bendix [KV05a]. Bernays [EIP20]. Beth [AJM^{+}23]. between [BFW03, FF20, GHO02, OSTW21]. Beyond [OSTW21]. BI [LWG13]. Big [Dag22]. Big-step [Dag22]. Bimodal [HK15b]. Binary [ACF^{+}22, CLLM19, NV19]. Bisimulation [FDY14, GJKW17, BG02, LT09, TM10]. blind [SV08]. Boolean [Ari07, BV02, CSS10, GHPW21, HP03, KRW12, LWG13]. bound [AI03]. Bounded [AF23, EEH17b, GKN^{+}20, GN23, PZ22, BB14, BSH10, BM10, BK02, DG07, GPW10, SS07]. Bounds [AGM13, BKV15, BBMP23, BIK^{+}21, Kla08, VZ19, AD14, BSH10, JK09, KS13b, SP09]. Branching [CLLM19, GJKW17, GC12, LS15, KV05b, ZHD14]. branching-time [ZHD14]. Branchwidth [LOS19]. broadcast [LvdMR00]. Brzozowski [BBH^{+}14]. Büchi [BK12, Tri09]. Budget [NT17]. Buss [BSH10].

Cai [PSS18]. Calculi [CGPT22, DHS16, GS13b, HCGT18, LZ19, LP19, BO02, GGOP09, MOG05, Vor01]. Calculus [AVD^{+}19, DMP18, GLSG15, GGS20, Kha22, LNH21, MN12, NS15, SPAV15, AJ05, ET06, GL10, GGOS09, JK09, KS13b, MM07, MP03, OPS07, PQ07, Rei01, Sal03, SP11, Sau10, TM10, VV07, ZHD14]. camera [BGLC04]. Can [HSTW20]. canonical [BD07b, HP05]. Canonisation [GN23]. capture [BG03a, BG08a, Gur00]. capturing [Lib01]. cardinality [CSS10, Sze11]. Case [NV19, GN11]. Cases [VP20]. Categories [Van18b]. Category [AMMU19, HPS22]. causes [CHK08, CHK10]. CCS [AFIL09]. Central [BCKL19]. Certainty [YSG09]. Certified [PDHR14, CS08]. Chains [BHM15, ASSB00, CDKM13]. Change [AG19, ACOS17, BZWS18, DSTW13, PKW12]. channel [BBS07]. Characterisation [Van18a]. Characterisations [BBMP23]. Characterization [GMR12, KS17, WWT15, DCHA03, Kon09]. Characterizations [CR15, EFW07, YRSW07]. characterizing [SYY^{+}03]. Checking [ALGN18, BP20, BCKL19, BGS16, BMM^{+}19, CFMP19, DHKP17, EvdHK^{+}20, EY12, FCGQ20, FKL15, GHO^{+}20, GRV21, MPZ23, MBRS20, MMPV14, Tri09, VP20, YLYF14, ASSB00, BMS^{+}11, Bou09, BDR08, BBD03, GGV02, LLM^{+}07, LW10]. Choiceless [PSS18]. CHR [DGM12, GM09a]. Church [KS13b]. Circle [DD15]. Circles [Soj16]. Circuits [DW22, Kra21]. Circular [AL23]. circumscription [CEG05]. Class [EIP20]. Classes [CGT17, GKN^{+}20, GHO^{+}20, GNT04, Pau06]. classical [AL10]. Classification [CM17, CKK17, HMR05]. Classifications [CES14]. Classifiers [NSM18]. Clausal [FDP01]. clauses [SV08]. Clique [HS15]. Clique-Width [HS15]. Closed [CP22, Les18]. Closure [CR20, YSG09]. CNF [DKS17]. Co [BK12]. coalgebra [BBH^{+}14]. Coalgebraic [BMS13, KPS22, SV18]. Coalitional [EG22]. code [CS08]. Cognitive [GF18, HKO19]. Coherence [LW15]. Coincide [EGT16]. collapse [BM10].

Collapsibility [CMMZ23]. Collapsible [BCH^{+}21, BCOS21, HMOS17]. combination [DEPT11, GNZ08]. combinations [BGR14]. Combinatorics [BM17]. combinatory [DZ13]. Common [SA13]. communication [AFIL09]. Commutative [HTAC19]. commuting [CKRT08]. compact [Ber04, DPR08, Lib00a]. Comparison [BMM^{+}19]. Comparisons [WZ05]. Compilability [Lib00a, LS07]. Complement [GN12, MP03]. Complete [BCEO20, BMS13, GRV21, LOW16, MPZ23, SDW14, AMMO09, Cra07, DCHA03, GHK08, KV05a, Lut04, Bau14]. Completeness [CGPT22, DD16, Dra16, HK21, SV18, Bou09, NT10]. Complexes [CC19]. Complexities [KRH13, LS15]. Complexity [ABM21, Asp15, AO19, BBC^{+}16, BGL13, BJP17, Bul11, CEG05, CGL16, CMMZ23, CM17, CKRT08, CMVT12, CES14, DHL12, DKS17, DT10, DKDV22, FFP15, FGT22, GJL15, HKVV18, Kha22, Lib06, Lib23, LMS17, MM12, MMS23, NC12, VZ19, Wro19, BMR09, BG09, Che14, CK00, Dal09a, ET01, EFL^{+}04, EFW07, FFF05, GMS12, HLT07, JK09, KHST12, LT03, VM09, SV08]. Components [AKNZ17]. Composing [BDFM23]. Composition [ABS^{+}23, AG19, ACF^{+}22]. Compositional [NSM18, GM09a, NT10]. compositionality [Rab07]. Comprehending [Cha06]. comprehensive [GNZ08]. compressed [GGSS11]. Compression [EEH17a]. computability [Bau14, Jap06a, Jap06b, TZ02]. Computable [BCP17, Sim15, BG01, DG07]. Computation [NSST18, AS02, BMM12, GM02, TZ04, WZ05]. Computational [ABEZIP03, Asp15, AVD^{+}19, Kou21, ASP09, BMR09, Dal09a, Ser01]. Computationally [BCEO20, GHK08]. computed [LY07]. Computing [BCD13, CJL13, GJKW17, HCMS13]. Concept [DHS16]. Conceptual [AKRZ14, TSH15]. concerning [CFS10]. concrete [BD07a, Lut04, TZ04]. Concurrency [AFMGMM16, BP20, DHL12, DHS16]. Concurrent [ALGN18, CKPP12, EG22, FS15b, SDW14, BMR06, CDH11, DGM04, NNN11]. condensing [GRS05]. conditional [FHK00, GGOS09, Luk01, OPS07]. conditions [GK08]. Configuration [ADGV14]. configurations [LvdMR00]. Confluence [FMZV15, Bou09, CNNR03, GNT04, KS13b]. Confluent [Sim15]. congruence [LT09]. Conjunction [DD16]. Conjunctive [CDL08, CM17, GKNS19]. Connectedness [KNPHZ13]. connectives [Fer11]. Connectivity [SSV23]. Conservative [Wro19, Bul11]. consistency [BCM09]. Consistent [Fon15]. Constant [GK18, IS06, DG07]. Constant-Depth [GK18, IS06]. Constrained [FKN17]. Constraint [BF13, BMT13, BJP17, DMM17, GF18, WDB13, Wro19, AR04, BGLC04, BMR06, BMRS10, BCM09, Bul11, CSS10, DGM04, DPR08, KV05a, MR03, YSG09]. constraint-solving [DPR08]. Constraints [BCF21, CMMZ23, NT17, BCM09, BV02, BFW03, CDL08, CLS07, HP03, Lib03, Luk01, Rat04, Rat06, Sze11]. Construction [LW15]. Constructive [CDG15]. Containment [AK15, BBGS20, GKNS19, CDL08]. Context [Dal09a]. Contexts [BGM14]. Contextual [HPS22, NPP08]. continuation [GM09b]. continuity [BK02]. Continuous [BFHH17, BLM^{+}23, SJV12, ASSB00, CDKM13, FR10]. continuous-time [ASSB00, CDKM13, FR10]. contract [KKW10]. contract-signing [KKW10].

Contractions [San17]. control [BGLC04, Moy09]. Convergence [Lyn05]. Convergent [Rat04]. Converse [GLSG15]. Convolution [DHS16]. Cookbook [AKRZ14]. cooperation [MN12]. Coordination [SI08]. copying [SV08]. Coq [HCMS13]. Coq/SSReflect [HCMS13]. Correct [APNT22]. Correction [BG08a]. Correctness [Dra16, GKNS19, LL15, Cha06, DGM04, KT03]. Cost [ACJ^{+}18]. Costs [ST15]. Counter [FCGQ20, BLM^{+}23]. counterexample [CV10]. counterexample-guided [CV10]. counterexamples [SG07]. Counting [ABM21, CW16, HK15b, KSS22, IS06, Kon09, Lib00b, Sch05]. Covering [PZ20]. Craig [AJM^{+}23, CGS10]. criteria [Vor01]. Criterion [Ara23]. Cryptographic [CCCK16, CLCZ10, SV08]. CSL [KM10]. CSP [BCGH23, CHN20, DKS17]. CSPs [BMV22]. CTL [SG07, ZHD14]. Cube [CC19]. Curve [NC12]. cycles [CLCZ10]. Cyclicity [ZG14].

Data [ACOS17, AKRZ14, DHL12, DKDV22, QS19, SDW14, BDM^{+}11, JL11, Laz11, Tan14, YSG09]. Databases [ABM21, AMN^{+}03, GHK08]. datalog [BK10, CFS10, GPW10, ACF05, AKNZ17, BBGS20, GGV02]. Decidability [BHM15, DFO06, Fig12, Kou21, MOK15, VH05, BDT10]. Decidable [CGT17, PR23, EŠ10]. decide [ST14]. Deciding [BG02, CLCZ10, CNNR03, GLSG15, KKW10]. Decision [BG22, FFF15, LZ19, MPZ23, NV19, BJW05, CV10, GS09, RV01]. Declarative [SJV12, GL13]. Decomposed [PS17]. Decomposing [GJL15]. decomposition [BG11b, MV04]. decomposition-based [MV04]. Deconstructions [CM17]. decorated [BFV04]. Deduction [Gul18, MMZ21]. Deductive [MO12, GGV02]. deep [BG09]. default [FHK00, Lib07]. Defeasible [BCM^{+}21, ABGM01, BAGM10]. Definability [GN23, PSS18, BLN07]. definable [BS09, EH01, Lib03]. Defining [Pau06]. Definitions [AJM^{+}23, MPZ23, WDB13, Avi03, DBM01, DT08]. Degree [AF23, AD14, DG07, GL10, Lyn05]. delay [DG07]. Dense [GHO^{+}20]. Dependence [DHL12, DK12, HLM13]. Dependencies [BB20]. Dependent [HPS22, LW15, SBTM06]. Depth [BKV15, EEH17b, GK18, PZ22, BSH10, IS06]. Depth-Bounded [EEH17b]. Description [ACOS17, AKRZ14, AJM^{+}23, BBK^{+}20, BCM^{+}21, DNR02, KRH13, BGL12, CDL08, EILS11, Lut04, ST14, WZ08]. descriptions [Ber04]. Descriptive [VZ19]. Detecting [CGT17]. detection [AGP07]. Determinacy [BKONR23]. Deterministic [AL04, HK15a, Mam16, Sei18]. Diagonal [CGL16]. Diagrams [NV19, KS13b]. Dichotomies [HLPW20]. Dichotomy [Fon15]. Differential [Kaw09, Pla15, Pla17]. Differentiation [TW23]. Diffie [CKRT08]. Diffusion [SBSM13]. Digraphs [DMM17]. Dimension [CP22, HLT07]. directed [Sto05]. Discrete [ACOW22, NC12]. Discrete-Time [ACOW22]. Disjunction [BF13, Sto05, FFF05]. disjunctions [JNS^{+}06]. Disjunctive [FS15a, EŠ10, WZ05]. Display [CGPT22, CR16, CLRT21]. Displayed [GP23]. Distance [FGT22]. Distributed [FLY22, GS13a]. Distributing [AKNZ17]. distribution [HP03]. Distributions [CR15]. distributive [SS07]. distributivity [SV10]. DL [WWT15]. DL-Lite [WWT15]. DLV [LPF^{+}06]. Do [HSTW20]. Does [Pra13]. Domain [SBTM06, DMS06]. Domain-dependent [SBTM06]. Domains [SJV12, AR04, GRS05, HW10, Lut04]. Dominance [KKLJU14]. Downward [Fig12]. DPLL [BGL13, Lib06]. Driven [AKS13, FFF15]. Dual [HTAC19, Kra21].

Dualities [CFMP19]. Duality [FF20, BBH^{+}14]. duration [CDKM13]. Durations [BDR08]. Dynamic [BB20, FS15b, LNH21, Sch15, Tak23, BERS04, FD14, GMS12, GK08, KB11, SYY^{+}03]. Dynamical [ACOW22, BHM15]. Dynamics [GSM^{+}22].

Eager [BT23]. Editorial [ALP02, AKS01, Apt05, BMR09]. Effect [PS17]. Effectful [GMS20]. Effective [BTV16, CP22, KS17, BJW05, KHST12]. Effects [DLM21, WT03]. Efficient [APNT22, CGS10, KKLJU14, Rat06, BGV01]. Efficiently [GLSG15]. Eilenberg [AMMU19]. Elementary [MOK15, AJ05, CM06]. Eliminating [Avi03]. Elimination [FGT22, GGS20, Cra07]. Eliminations [VBH23]. Email [SA13]. Embedding [DEPT11]. Embeddings [CG18]. emptiness [Tri09]. encryption [CKRT08]. Enough [APNT22]. enriched [AL10]. Entailment [MPZ23]. entailments [Ari07]. Enumeration [KS13a]. environment [BD07a, PQ07]. Epistemic [CKPP12, CFP^{+}19, HV18, VP20]. Equality [BT23, MM12, Pit20, BGR14]. Equation [KRW12]. equational [AFIL09]. Equations [San17]. Equilibrium [GHPW21, SJV12]. Equivalence [ALGN18, BJK22, BJN20, BMS13, CCCK16, CC19, GJKW17, OSTW21, Soj16, GJ12, HP05, Mur05, Pau06, SH06]. Equivalences [CSN05, EFW07]. equivalent [LPV01]. Erotetic [CKK17]. Erratum [CHS15, CHK10, VGD07]. erroneous [YSG09]. Eternity [Hes05]. Euclidean [KNPHZ13]. Evaluation [FFF15, MBN12, RS14]. Evaluation-Driven [FFF15]. Event [SPAV15]. evolving [EFST05]. Examples [Lib16]. Exchanges [SA13]. execution [DLS01]. Existential [BGS16, LMTV19, PZ22, Che14]. Expansion [BB20, GKN^{+}20]. exponentials [BG11b]. exponentiation [CKRT08]. Expressions [GN12]. Expressive [DD16, DGM12, GL13]. Expressiveness [BMM^{+}19, FCGQ20, GHPW21, ZZ17, BDT10]. EXPSPACE [HK21]. EXPSPACE-Completeness [HK21]. EXPTIME [AE09]. Extended [BB17, BH23, LF01]. Extending [CPV09, MNPS20, Tan14]. extensible [BBC02]. Extension [Tak23, BG08a]. Extensional [CHRW13, SH06]. Extensions [LS15, GGOS09, KB11]. exteriors [MO12]. External [CJL13]. Extraction [Bd12, CKK17].

fail [Lib07]. fail-safe [Lib07]. Fails [CD14]. failure [DNR02, RW05]. Fair [GS13a]. Fall [DMP18]. Fast [LMTV19, MO06, BK10]. Faster [AF23, DHKP17]. FDNC [EŠ10]. feasibility [IS04]. feature [Mie04]. Field [ZG14]. Field-Sensitive [ZG14]. Fields [AVD^{+}19, BJK22]. FIFO [GGKT12]. Finding [DKS17]. Finitary [CHH09, Mur05]. Finite [ACF^{+}22, ACMU21, AB20, Asp15, BJK22, BDFM23, BBC^{+}16, BKV15, EEH17a, GLSG15, KT18, NSV04, NSST18, RS14, AR04, AFIL09, EH01, FD14, GPW10]. Finite-State [NSST18, EH01]. Finite-Variable [BKV15]. Finitely [CGT17]. First [AZZ14, DG12, DG07, EGT16, FGT22, FHK00, GKN^{+}20, GHK08, GJL15, HK15b, Kar13, KT15, MM12, PR23, PZ22, SSV23, TOD20, WDB13, Avi03, Bau14, GS00, GS02, Lib03, Rat04, Sch05]. First-Order [AZZ14, EGT16, FGT22, GKN^{+}20, GJL15, HK15b, Kar13, KT15, MM12, PZ22, WDB13, DG12, DG07, FHK00, GHK08, SSV23, TOD20, Avi03, Bau14, GS00, GS02, Lib03, Rat04, Sch05]. Fixed [Bae12, GGS20, LT03, DGK04, HW10].

Fixed-parameter [LT03]. Fixed-point [GGS20]. Fixpoint [BCF21, CGWW15, SV18, SSS12, VGD06, VGD07]. Flat [SV18, SV08]. Flow [NT17, JK09]. Flows [ABS^{+}23]. Floyd [AMMO09]. Floyd-Hoare [AMMO09]. FO [BS09, GHO^{+}20]. Focalization [Sim14]. Focusing [Bur20]. Forests [BDFM23]. Formal [Gui22, Kra15, GMS12, Gui09]. Formalizing [SPNS14]. formally [ADGR07]. formats [BFV04]. Forms [RMV17, HP05]. Formula [HLM13, AI03]. Formula/Literal [HLM13]. Formulae [MBRS20, Ari07]. Formulas [BLRC^{+}19, DKS17, DU21, TW23, AD14, GS00, LR06, Sub04]. Formulation [CMMZ23]. forth [GHO02]. Forward [FF20]. Foundational [CS08]. foundations [GL13]. Founded [BVD16, CR20, EILS11, WZ05]. Fragment [HRW17, HLPW20, KT18, PR23, AJ05, Bau14]. Fragments [BKMV^{+}17, CMVT12, CPRW16, AL04, BMS^{+}11]. Frames [Sha22]. Framework [MMS23, TSH15, ACG^{+}08, AAD11, AMMO09, BD07a, CV10, CS08, GNZ08, LA13, Lia05, MM02, RS13, She08, SG07]. Frameworks [FFP15, BCM04]. Free [CCDD20, DLM21, BGR14, HV05]. Freedom [KE19]. freeze [DL09]. Frege [BB17, BSH10, GK18, IS06]. Frontier [DKDV22]. full [ST14]. fully [LA13]. Function [CGT17, AE09, EŠ10]. functionals [BK02]. Functions [CGI^{+}18, NSM18, AE09, Avi03, BGV01, Pau06]. Fürer [PSS18]. fusion [Lia05]. Fuzziness [BKM^{+}23]. Fuzzy [FPS14, SJV12].

G3i [FFF15]. Game [BBDM23, CS20, FMS16, GKR18, Pla15, DH02, MN12, SG07]. game-based [MN12, SG07]. Game-Theoretic [GKR18]. Games [BPT14, BMM15, BCH^{+}21, CD14, CKPP12, EG22, GHPW21, LPR18, MW15, Pla17, AL04, BSH10, CHH09, CDH11]. General [BGM14, ZZ17, AMMO09, SYY^{+}03]. Generalized [AMMU19, Kon21, Kon23, MBN12, SBSM13]. Generalizing [Ara23, BCM09]. Generating [BH23]. Generation [CKK17, AR04, CGS10]. generators [AL04]. Generic [BCD13, MT05]. Geometry [Ham18, Dal09b]. Glabbeek [Ham18]. Global [KPS22]. Goal [EG22, Sto05]. goal-directed [Sto05]. Gödel [Gul18]. Good [BBDM23, CS20, BMS^{+}11]. Good-for-Game [BBDM23]. governed [ASP09]. GPU [KI17]. Graded [BMM12]. Grammars [EEH17a]. Graph [ACOS17, BM17, CM17, GHO^{+}20, Tan13, TW23, DZ13]. Graph-Structured [ACOS17]. Graphs [GN23, KSS22, Sei18, Lyn05, Moy09, Sze11, Tri09]. Greatest [Bae12]. Ground [CGT17, GL14, Bou09, VH05]. Guarded [BTV16, HLPW20, KT18, MPZ23, DLS01, GHO02, HVV08]. Guards [KT18]. Guest [BMR09]. guided [CV10].

Hájek [BM08]. Halpern [BKMV^{+}17]. Handling [BF13, GF18]. Happen [AFMGMM16]. Hard [Fon15, HSTW20, SUWC16]. Hardness [BBMP23]. Harrington [CGL16]. Having [Sim15]. Heads [DGM12]. heap [YRSW07]. Hellman [CKRT08]. Help [Pra13]. Herbrand [McK13]. Hercules [LOW16]. Heterogeneous [DKS06, Pit20]. Hierarchies [DK12, Han18]. Hierarchy [BKV15, CM17, GL14, KT15, KS15, KS17, Kon09]. Higher [AVD^{+}19, BCOS21, CHRW13, JR15, Kar13, KS15, LV12, MP03, Pie09, CPV09, Dal09b, IS04, MM02]. Higher-Order [AVD^{+}19, CHRW13, JR15, Kar13, LV12, BCOS21, MP03, Pie09, CPV09, Dal09b, MM02]. highly [GJ12]. Hoare [AMMO09, CK00, KI17, Koz00, Mam16]. Hodges [BBDM23]. Homological [PDHR14]. homology [HCMS13]. Homotopies [KS15]. Homotopy [Soj16]. Horn [BKMV^{+}17, HRW17, KRH13, Lib00a, MBN12, MO12]. Hughes [Ham18]. Hybrid [AKS13, Pla17]. Hydra [LOW16]. hypersequent [MOG05]. Hypersequents [CG18]. Hypotheses [KKLJU14]. Hypothesis [LSS04]. Hypothesis-based [LSS04].

Idempotent [CK20]. Identified [KSS22]. Identifying [KKLJU14]. Identity [Van18b, vdBG12, ST14]. II [BG07a, Jap06b]. III [BG07b]. Image [PDHR14]. IMLL [MO06]. Immerman [PSS18]. Imperfect [BMM^{+}21]. Implementation [SSS12, MV04]. Implicit [GMR12, BMR09]. implies [Cha06]. Improved [BB14]. improvement [BB14]. Inadequacy [BG01]. Inclusion [FKL15, GRV21, Han18, HKMV20, BAGM10]. Inclusions [AJM^{+}23]. Incomplete [ABM21, BLRC^{+}19, YSG09]. incompleteness [Sal03]. increasing [AS02]. Incremental [CGI^{+}18, DLS01]. independence [Lib03]. Index [FMS16]. indexed [LB07]. indexing [Pie09]. Indistinguishability [BCEO20, Kou21]. Induction [FKN17, Sak05]. Inductions [BVD16]. Inductive [MPZ23, WDB13, DBM01, DT08, LF01]. inequality [Rat06]. inexpressibility [BK10]. Inference [AGM13, BBT^{+}21, Kou21, Wro19, ZG14, BD07b, BG09, CM06, GM02, MO12, VV07]. Inferring [GK08]. Infinite [CHS14, CHS15, Tan13, BMT13, NSV04]. Inflationary [DGK04]. infons [GN11]. Information [ABS^{+}23, BBT^{+}21, BMM^{+}21, FL20, KBP23, NT17, KB11]. initial [LvdMR00]. Inner [DD15]. Inputs [ABS^{+}23, BCGH23, CHN20]. Inspired [VBH23]. Instances [DKS17]. Instantiation [EP13]. Integer [CLZ23].

Integers [BLRC^{+}19, BJW05]. Integrity [BCF21]. intelligence [Cha06]. intelligence-related [Cha06]. Interaction [BFW03, Ham18, Sei18, ACG^{+}08, BG11b, Gug07]. Interactions [MW15]. Interactive [Bd12, Kra15, BG06, BG07a, BG07b]. interiors [MO12]. Interpolants [AJM^{+}23, CGS10]. interpolating [BGR14]. Interpolation [BTV16, GGKN23, BGR14]. Interpretation [MZ17]. Interpretation-Based [MZ17]. Interpretations [GKN^{+}20, BBC02, MP09]. Intersection [CP22, DU21, GN12, Van18a, DCHA03, DCDGT10]. intersection-type [DCHA03]. Interval [BBK^{+}20, BGLC04, BMM^{+}19, BMP23, BKMV^{+}17, DHS16, HGS07]. Interval-Rigid [BBK^{+}20]. intractability [CFS10]. Intruder [BCD13]. intuitionism [LM02]. Intuitionistic [AR02, BBDFD20, FFF12, FF20, GGS20, Kon23, Kra15, DST13, FFF05]. Invariant [EEH17b, GS00]. Invariants [ACOW22, BG01]. Isabelle [Pau00]. Isomorphism [TW23, VZ19]. isomorphisms [DCDGT10]. issue [BMR09, GHJP03, KM10, Kol05]. Iterated [GHPW21, Lib23]. iteration [HS10]. IV [BG11b].

Jordan [NC12]. judgments [MT05]. Justification [SFL17]. Justified [SFL17].

K4 [HK21]. Karp [BM10]. Kernels [KI17]. key [CKRT08, CLCZ10]. kinds [Cra07]. Kleene [DMS06, Koz00]. KLM [BCM^{+}21, GGOP09]. KLM-style [BCM^{+}21]. Knowledge [AKS13, BCD13, BBT^{+}21, CKK17, DN21, GSM^{+}22, Kra15, LvdMR00, MBN12, SA13, TSH15, TOD20, DEPT11, DNR02, EFL^{+}04, EFST05, LPF^{+}06, Lib00a, Rei01, SBTM06]. knowledge-base [DEPT11]. knowledge-based [Rei01]. knowledge-state [EFL^{+}04]. Knuth [KV05a]. Kripke [DU21, Sha22].

Labeled [CLRT21]. Labelled [HCGT18]. Labels [BGM14]. lambda [AJ05, PQ07, Sal03, Les18]. lambda-calculus [PQ07]. Language [BMS13, GRV21, CPV09, GGV02]. Languages [AMMU19, BL16, EEH17a, Wro19, BLN07, BS09, CDH10, GL13, GHK08, GMS12, Mur05]. Lattice [HK15a, BMT13]. lattices [SS07]. law [Lyn05]. Laws [PZ22]. Lax [Han18, HKMV20, Rab15]. Layer [FMZV15]. Layers [NDH19]. Learning [NSM18]. Least [Bae12, KS13b]. left [AFIL09, MR03]. Length [GGKN23, Sub04]. LF [HP05, UCB11]. LICS [GHJP03, Kol05]. Light [AR02]. like [Sub04]. Limit [BCKL19]. limitation [Cha06]. limitations [Rab07]. Limited [BBGS20]. Limiting [BL16]. Limits [CR16]. Linear [APNT22, Bae12, BBDFD20, BF13, BMV22, BLRC^{+}19, FPS14, GP23, HK15b, Les18, MMZ21, ST15, Tak23, YLYF14, BBS07, BJW05, CDKM13, Dal09a, Dal09b, GGV02, HV05, KRS05, KV05b]. Linear-Logic [BF13]. Linear-Time [YLYF14, BBS07]. Linearizability [SDW14]. Linearization [CGI^{+}18]. Linking [Bur20]. Lipton [BM10]. Liquid [MNPS20]. lists [DPR08]. LITE [GGV02, WWT15]. Literal [HLM13]. Livelock [KE19]. Livelock-Freedom [KE19]. Liveness [EFH14, HSTW20]. Living [AJM^{+}23]. Local [CLZ23, GSM^{+}22, LW15, PS17, BB14, GM02, Lib00b, Lib01, Sze11]. Local-Effect [PS17]. Locality [GS00]. Locally [BVD16]. LoCo [ADGV14]. Logic [ABS^{+}23, APNT22, ADGV14, AL23, Bae12, BP20, BBC^{+}16, BMM^{+}21, BF13, BZWS18, BGS16, BMM^{+}19, BMP23, BKMV^{+}17, CGT17, CDG15, CHRW13, CW16, CR20, CGWW15, CMVT12, CES14, CPRW16, DW22, DD15, DD16, DLM21, DBM01, DHS16, Dra16, DK12, EIP20, EGT16, EG22, FL20, Fer11, FS15a, FP17, FF20, FGT22, FPS14, GS13b, GKR18, GP23, Gul18, GN11, GC12, Han18, HKMV20, HK21, HTAC19, HV18, Kar13, KP18, KI17, Kon21, Kon23, Kra15, KS17, Kuz21, LOW16, LNH21, LMS17, MM12, MMS23, Mam16, MW15, MPZ23, MBRS20, MSS14, PR23, Pla15, RS14, SSV23, SJV12, SSS12, SDSS13, SPNS14, SFL17, Tak23, TOD20, Try16, VP20, WEZL23, WDB13, Zha06, ZZ17, AL10, AM01, ACG^{+}08, AELP01, AGP07, ABGM01, Avi03, BGL12, Bau14, BMM12]. logic [BMRS10, BDM^{+}11, BERS04, BM08, BDP04, BGV01, CDL08, CK00, CM06, Dal09a, DGK04, DEPT11, DL04, DT08, DG12, DST13, EFL^{+}04, EŠ10, EILS11, FD14, FHK00, FR10, GL13, GK08, HV05, IS04, Jap06a, Jap06b, KR02, Koz00, KT03, LF01, Lia05, LPV01, LT03, LSS04, Luk01, MR03, RW05, Sak05, SH07, SKGST09, Sch05, SPSS11, SYY^{+}03, Sze11, Tan14, VDS01, WZ05, ZHD14, AR02, DBM01]. Logic-Based [CES14, MW15, MMS23, MBRS20]. logic-enriched [AL10]. Logical [BDT10, BLN07, BHM15, BFHH17, BCOS21, CR15, GL13, RS13, YRSW07, Kon09, MM02]. Logics [ACOS17, AKRZ14, AJM^{+}23, BBK^{+}20, BBDFD20, BM17, BDFM23, BTV16, BKV15, BGMZ14, BCM^{+}21, CLSZ15, CLRT21, EEH17b, GW22, HK15b, HKVV18, HK21, HCGT18, KB11, KBP23, KSS22, KNPHZ13, KRH13, KPS22, KWS^{+}03, Lib00b, Lib01, MMZ21, MOK15, SV18, AMMO09, BAGM10, BO02, DNR02, ET06, FFF05, GGOP09, GGOS09, GS09, GHO02, Lib07, LSS04, Lut04, MOG05, OPS07, ST14, SP09, TM10, VGD06, VGD07, Vor01, WZ08]. Long [ALN16]. loop [BG01, CPV09, LR06]. Loops [CJL13, She08]. Lower [AGM13, BSH10, BBMP23, BIK^{+}21, AI03, AD14].

LTL [BGL12, BMS^{+}11, DL09, SV10, AL04]. Lukasiewicz [MW15, MOG05].

Machines [FCGQ20, GMS20, BD07a, BG03a, BG08a, Gur00, NSV04]. Made [BK12]. maintenance [FL20]. Majority [BJK22]. Making [GRS05]. MALL [Ham18]. Managing [ACOS17]. many [LR06]. Map [NSST18]. Map-Reduce [NSST18]. Maps [PKW12]. Markov [ASSB00, BHM15, CV10, CDKM13]. marriage [WT03]. matching [GGSS11]. mathematics [AL10]. MaxDiff [GGKN23]. Maximally [ALN16]. MaxSAT [FMSV23]. May [AFMGMM16, ALN16]. May-Happen-in-Parallel [AFMGMM16]. Meaning [Bas21]. measuring [AELP01]. Mechanizing [Pau00, UCB11]. Mediated [HLPW20]. Meets [AO19]. MELL [LS15]. Memory [DDDW22]. merge [AFIL09]. Mergesort [LL15]. Merging [CPRW16, HRW17, Lib16]. Messages [CCDD20]. Meta [Dag22]. Meta-theory [Dag22]. metalogical [BCM04, CS08]. metaquerying [ABEZIP03]. metatheory [UCB11]. Method [CDG15, MBN12, MP09]. Methods [BCEO20, NT10, Vor01]. Metric [BBK^{+}20, LOW16, FR10, KWS^{+}03, TZ04]. Minimal [ACOW22, Wro19, DNR02]. minimization [BBH^{+}14, BG03b]. Minimizing [HK15a]. Minimum [RW05]. Minkowski [LPR18]. MITL [MBRS20]. Mixed [Lib23]. MKNF [AKS13]. MLL [MO06]. mobile [DL04]. Modal [AJM^{+}23, BDFM23, DMP18, GJL15, HKMV20, KP18, KPS22, MMZ21, MOK15, NDH19, NNN11, PR23, Pra13, DGK04, DL04, DST13, FFF05, GHO02, Lia05, NPP08, SH07, SP09, TM10, Vor01, WZ08]. Modalities [GCK^{+}22, LP19]. moded [BERS04]. Model [AF23, ASSB00, BP20, BCKL19, BGS16, BMM^{+}19, BBD03, CHN20, CDG15, DHKP17, DSTW13, EvdHK^{+}20, EY12, FCGQ20, GHO^{+}20, KP18, LW15, MBRS20, MMPV14, VP20, WWSL16, YLYF14, ZZ17, AELP01, BMS^{+}11, BDR08, GGV02, JNS^{+}06, LLM^{+}07, LW10, RW05]. Model-Checking [EvdHK^{+}20, MMPV14, YLYF14, ASSB00, BDR08]. Model-Theoretic [DSTW13]. modeling [CLS07, ET06]. Modelling [AKRZ14, HTAC19]. Models [AG23, BCGH23, RS14, vdBG12]. Modular [HCGT18, Rab18, AAD11, Sto05]. Modularisation [LP19]. modularity [VGD06, VGD07]. Modulo [CLZ23, CGI^{+}18, ST15, CGS10]. Monad [ACMU21]. Monadic [BBGS20, EGT16, FP17, GPW10, Mam16, Sze11, KS13a]. monads [WT03]. Monodic [DFK06]. Monotone [BVD16]. Monotonic [BKONR23]. Morgan [HTAC19]. Morphisms [Rab15]. Most [CJL13]. motion [BGLC04]. MSO [EH01]. MTL [FCGQ20]. Mu [DMP18]. Mu-Calculus [DMP18]. multi [Lia05]. multi-agent [Lia05]. Multiagent [HKO19, GS09, LvdMR00]. multiary [SP11]. multicast [BMRS10]. Multicriteria [KKLJU14]. Multigranulation [KP18]. Multiplayer [EG22]. Multiple [DGM12, PKW12]. multiplicative [HV05]. multiplicative-additive [HV05]. Multirelations [FS16]. multisets [DPR08]. Multiteam [GW22]. multivalued [LSS04]. MWeb [AAD11]. mwp} [JK09]. mwp}-bounds [JK09].

named [VV07]. Names

[BBK^{+}20, HTAC19, KHST12]. Narrow [ALN16, BCMS18]. Nash [GHPW21]. Natural [CFMP19, MMZ21, LL15]. Nature [CS20]. Necessary [ACF^{+}22]. Negated [Van18a]. Negation [GKNS19, CSN05, DNR02, RW05, ST14]. negation-as-failure [RW05]. Negotiations [SPNS14]. Nested [CHO17, EP13, Kar13, MMZ21, AE09, CEG05, VV07]. Nets [BFHH17, CC19, Ham18, LNH21, McK13, HV05, MO06]. Network [GSM^{+}22, SBSM13]. NEXP [Lut04]. No [DD16, Sim15]. Nominal [FF18, HTAC19, LV12, DG12]. Nominals [AJM^{+}23]. Non [BLRC^{+}19, CR20, FD14, HTAC19, LP19, Sei18, AS02, GK08]. Non-Commutative [HTAC19]. Non-Deterministic [Sei18]. Non-finite [FD14]. Non-Linear [BLRC^{+}19]. Non-normal [LP19]. non-size-increasing [AS02]. non-suspension [GK08]. Non-well-founded [CR20]. Nonconstructive [Bd12]. Nondeterministic [LWG13, Mam16, BBS07]. Nonelementary [LS15]. nonground [DEPT11]. Nonlinear [CGI^{+}18]. nonmonotone [DT08]. nonmonotonic [BO02, ET01, EFST05, EŠ10, GGOP09, Sak05]. Nonuniform [CSS10]. norm [ASP09]. norm-governed [ASP09]. Normal [JR15, LP19, MMZ21, RMV17]. Normalisation [Van18a]. Normality [FS15a]. Normalizing [Sim15]. normative [Ser01]. normed [GV04]. note [CK00]. NP [BB17, KV05a]. NP-complete [KV05a]. Nullstellensatz [IS06]. Number [AB20, BT23, TW23, ADGR07]. numbers [ACF05, CFS10, Rat06]. Numerical [NV19].

OBDD [BIK^{+}21]. objects [BBC02]. Observation [CD14]. Obtain [ACF^{+}22, Fon15]. Off [CGL16]. Off-Diagonal [CGL16]. Omega [LW10]. Omega- [LW10]. One [BLM^{+}23, CJL13, CM17, FCGQ20, HK15b, PZ22, Sim15, SV08]. One-Counter [FCGQ20, BLM^{+}23]. One-Step [Sim15]. One-Variable [HK15b, SV08]. Ontology [HLPW20, WWT15]. Ontology-Mediated [HLPW20]. Open [AB20, HVV08].

Open-world [AB20]. Operational

[DDDW22, GF18]. operations [BMT13]. operator [VGD06, VGD07]. Operators [ACF^{+}22, BZWS18, BVD16, HLM13, SSV23, DST13]. OPL [vHPP00]. Optimal [CDG15, Sub04, CM06, PQ07]. Optimality [GL10]. Optimization [CKK17, ST15, SBSM13]. optimize [Vor01]. Optimizing [CM06]. Oracle [BCP17]. Order [AZZ14, AVD^{+}19, CHRW13, EEH17b, EGT16, FP17, FGT22, GKN^{+}20, GJL15, HK15b, JR15, Kar13, KT15, LV12, MM12, PR23, PZ22, WDB13, AJ05, Avi03, BKZ20, Bau14, BB14, BCOS21, CPV09, Dal09b, DG12, DG07, FHK00, GHK08, GS00, GS02, KS13a, Lib03, MM02, MP03, Pie09, Rat04, Sal03, SSV23, Sch05, Sze11, Tan14, TOD20]. Order-Invariant [EEH17b, GS00]. Ordered [BL16, BGS16, EvdHK^{+}20, NV19, CNNR03]. orderings [BG08b]. Orders [BIK^{+}21, KRS05]. Ordinary [BG06, BG07a, BG07b]. Other [CLSZ15, BCM09]. Out-of-order [BKZ20]. Outlier [AGP07]. Output [BCP17]. Outputs [ABS^{+}23, BCGH23, CHN20]. Overlooked [LW15].

Paraconsistent [Ari07, CLSZ15]. Parallel [ACF^{+}22, AFMGMM16, ACJ^{+}18, GKNS19, SDSS13, BG03a, BG08a, PQ07]. Parallel-Correctness [GKNS19]. parameter [LT03]. Parameterized [AK15, BGL13, DKS17, FGT22, KE19, LDJ^{+}18, MMS23]. Parametric [AELP01, GCK^{+}22, BDR08]. Parametrised [LMS17]. Parikh [Ara23]. Paris [CGL16]. Parity [BPT14, BCH^{+}21, GK18, AD14, CDH11]. Partial [CD14, BG08b, KT03, TZ04]. Partial-Observation [CD14]. partiality [JNS^{+}06]. Partially [BGS16]. passing [Ber04, GM09b]. Path [Van18b, BFW03].

Paths [DMP18]. pattern [MP03]. PCL [GGOS09]. Pebble [BGMZ14, Tan13]. Pebbling [Nor12]. PELCR [PQ07]. Perfect [BJP19, Kra15]. Permissive [DG12]. Permissive-nominal [DG12]. persistency [ACF05, CFS10]. Persistent [BG11a, HCMS13]. Perspective [GHO^{+}20, LV12]. Petri [BFHH17, CC19, LNH21]. Phase [LWG13, WWSL16]. PHP [Kra21]. Phylogeny [BJP17]. Piece [NSM18]. Piece-Wise [NSM18]. Piecewise [BMV22, GGKT12]. Plane [Try16]. planning [AM01, BGLC04, EFL^{+}04, SBTM06]. PLP [She08]. Point [BMM^{+}19, GGS20]. pointer [HS10]. Points [Bae12, DGK04]. Polymorphic [VV07]. Polynomial [GM02, Kha22, PSS18, AS02, GL10, GNT04]. Polynomial-time [GM02]. polynomially [IS06]. positions [Ser01]. Positive [MM12, SUWC16, Che14]. Possibilistic [SFL17]. Power [CFVZ20, CR16, DGM12, NS15]. Precongruence [BFV04]. Predicate [ET06, LB07]. Predicate-calculus-based [ET06]. Predicates [DLM21, EIP20, FP17, LB07]. predicative [AL10]. preference [GGOS09]. preference-based [GGOS09]. preferential [Ari07]. Preferred [AZZ14]. preorders [DCHA03]. Presburger [Kla08]. Preservation [BTV16]. Preserving [BZWS18, TSH15]. prime [ADGR07]. principle [SH07]. principled [AAD11]. Principles [BCM^{+}21, BB14]. privacy [MN12]. Private [HTAC19]. Probabilistic [BJK22, CFP^{+}19, DH02, DNS00, EY12, FFP15, HGS07, LT09, LLM^{+}07, Luk01, MSS14, SSS12, SDSS13, SPAV15, VP20, BBS07, DKS06, ILNR09, SPSS11]. Probability [BCP17, CR15, TOD20]. Problem [BHM15, DMM17, MMPV14, SJV12, Wro19].

Problems [ABM21, ADGV14, BB17, BJP17, DHL12, FMS16, GGKT12, PSS18, SBSM13, Sha22, ABEZIP03, Bul11, CSS10, ET06, GNT04, WZ08]. Procedural [FKN17, CPV09]. Procedure [FFF15, LZ19, MPZ23, BJW05, RV01]. Procedures [AKS13, BGL13, ABRS09, GS09]. Process [GL14, Ber04, BG02]. Processes [CKPP12, FDY14, CV10, YFDJ09]. Processing [PDHR14]. Processor [BGV01]. producing [Ber04]. Product [Soj16]. Program [Bd12, BG08b, BJN20, DHL12, MZ17, GJ12, MP09, Mur05, Zha06]. program-based [Zha06]. Programming [BMV22, CHRW13, DSTW13, SPNS14, AM01, ACG^{+}08, AGP07, BMR06, BMRS10, DBM01, EFL^{+}04, EFW07, GL13, HVV08, LF01, Luk01, Rei01, SI08]. Programs [AZZ14, BJK22, BZWS18, CGT17, Dra16, FLY22, FS15a, FKN17, GS13b, OSTW21, RS14, SBSM13, SDSS13, WEZL23, WWSL16, ZZ17, ACF05, ACW12, BERS04, BDP04, CFS10, DGM04, DEPT11, DBM01, DNS00, EŠ10, EILS11, Fer11, GK08, HVV08, HS10, LPV01, LT03, LSS04, MR03, RW05, Sak05, SKGST09, SYY^{+}03, VDS01, WZ05]. Progression [PS17]. promise [DT10]. Proof [AO19, BG22, BM08, CGL16, CR20, ET01, Ham18, HV05, Kha22, LL15, McK13, SDW14, TM10, TG18, ADGR07, BG09, DST13, MT05, MO06, Sto05, Vor01]. Proof-complexity [ET01]. Proof-Nets [Ham18]. proof-search [Vor01]. Proofs [ALN16, AL23, BPT14, BB17, Bd12, BCMS18, BH23, BIK^{+}21, CLRT21, GM09b, Kra15, NDH19, BSH10, DT10, SKGST09, VDS01]. Propagation [WDB13, CLS07]. Proper [CGPT22]. Properly [GP23]. Properties [BHM15, BCOS21, CCCK16, DHKP17, FGT22, Van18a, YLYF14, ABEZIP03, BBS07, BCM09, CDKM13, CLCZ10, KKW10, Lib00b, Lib01]. Property [AF23, FFF05, LLM^{+}07].

Proposition [BP11]. Propositional [AL23, BPT14, CPRW16, FFF12, FF20, GGS20, HK15b, HKVV18, HLM13, Jap06a, Jap06b, Van18b, BO02, BGV01, CEG05, CK00, DT10, Fer11, GN11, LA13, LS07]. propositions [DST13]. protocol [MN12]. Protocols [CCCK16, CCD15, CCDD20, AM01, CKRT08, CLCZ10, KKW10, SV08]. Provability [FF20]. prove [Hes05]. prover [OPS07]. Proving [DGM04, FMZV15, LDJ^{+}18, NC12, MR03, Vor01]. PSPACE [Bau14, GMR12, SP09]. PSPACE-complete [Bau14]. public [CKRT08]. Pudlák [BSH10]. Pure [HS10, LZ19]. Pushdown [BMP23, BCH^{+}21, BCOS21, CCD15, CLLM19, FKL15, HMOS17, Kar13].

QBF [BCMS18, BB20, BBMP23]. QCSP [LMM^{+}22]. QoS [BMRS10]. QPTL [BBDM23]. QSQR [MBN12]. Qualitative [BFF^{+}21, CDH11, ILNR09]. Quality [BKM^{+}23]. Quantification [Gui22, SUWC16, KR02]. Quantified [CMMZ23, DMM17, Ari07, BCM09, Rat06]. Quantifier [BKV15, BGR14, DLM21, PZ22, DL09]. Quantifier-Free [DLM21, BGR14]. Quantifiers [HTAC19, AD14, Rat04, Sch05]. Quantitative [CDH10, Les18, MW15, MM07]. Quantum [ALGN18, FDY14, FLY22, Tak23, YLYF14, YFDJ09]. Queries [AKNZ17, BKONR23, CM17, GKNS19, NSST18, BDT10, BG11a, Che14, CSN05, DG07, GS02, KS13a]. Query [AKS13, AB20, BL16, Fon15, LMTV19, SDSS13, BLN07, CDL08, GHK08, GGV02]. Query-Driven [AKS13]. Querying [HLPW20]. queues [RV01].

R [GF18]. Ramsey [CGL16, EIP20, FKL15]. Ramsey-Based [FKL15]. Random [BCP17, CP22, WWSL16, KHST12, Lyn05].

Randomization [CHS14, CHS15]. randomness [HLT07]. Rank [DW22, GN23, SP09]. rank-1 [SP09]. ranked [BLN07]. Rational [BM17, BT23, ST15, Try16]. Rationale [Lib23]. RDF [URS10]. Reachability [CLLM19, DLM21, GGKT12, Tan13, ZG14]. Real [Try16, Rat06]. Realizability [Kon21, Kon23]. Realizers [Bd12]. reals [BJW05]. Reason [FL20]. Reason-maintenance [FL20]. Reasoning [BMM15, BKM^{+}23, BJP19, CMVT12, EFST05, HKO19, ILNR09, KP18, KPS22, LNH21, MM02, MMPV14, Sch15, She08, SFL17, TSH15, TOD20, Ari07, BK10, ET01, FHK00, GGOP09, LPF^{+}06, MN12, NT10, YSG09]. Recall [BJP19]. Recognition [SPAV15]. Reconstruction [Rab18]. Recursion [BCOS21, HMOS17, Mam16, AE09, BG02, Dal09b, Kaw09]. Recursive [BKONR23, EY12, She08]. Recycling [LY07]. Reduce [NSST18]. Reduced [NV19]. Reducible [AG23]. Reduction [BB20, PDHR14, CM06, DZ13, PQ07]. reductions [BGV01]. redundancy [Vor01]. redundant [CLS07]. refinement [BG02, CV10, SG07]. Refinements [NDH19]. Refining [GL14]. Reflective [BCM04]. Reflexive [LMM^{+}22]. Refutation [FF20]. Refutation-search [FF20]. refutations [IS06, Sub04]. Region [Try16]. Region-Based [Try16]. Register [QS19, CLLM19, DL09]. Regular [BS09, GN12, BBS07, CHH09, LW10]. Reiterman [ACMU21]. Related [KT18, Cha06]. Relating [DMP18]. Relation [KKLJU14]. Relational [NSST18, AMN^{+}03, VV07]. Relations [BM17, Bas21, CR15, EFH14, GM02, RS13]. Relative [Nor12]. Relaxation [BMV22]. Relevance [Ara23]. Relevance-Sensitive [Ara23]. Reliable [YSG09]. Removing [CLS07]. Renaming [Bas21]. replacement [CT06]. Representation [AG19, ABGM01, LPF^{+}06]. representations [Lib00a]. Representing [MSS14]. Resolution [BCMS18, BBMP23, BH23, Bur20, FLM^{+}15, FMSV23, Kha22, LNH21, NDH19, Nor12, TW23, DFK06, FDP01, Lib06, Sub04, ZHD14]. Resource [AGM13, BK02, HP03, Moy09]. Resource-bounded [BK02]. Resource-distribution [HP03]. resources [MP09]. Restrictions [AB20]. result [BM10]. Results [GGKN23, MM07, ABGM01, ABRS09, CKRT08, CFS10, DFO06, ET01, Lib06, VGD06, VGD07]. Revision [Ara23, WWT15, Lib00a]. Revisions [Lib23]. revisited [DBM01, FHK00]. Revisiting [KR02]. Rewritability [BKONR23]. Rewrite [GL14, RMV17, ABRS09, CNNR03, GNT04, LY07]. rewrite-based [ABRS09]. Rewriting [FF18, FKN17, Sim15, GK09, KS13b, SKGST09, VH05]. Rigid [BBK^{+}20, EEH17a]. Rings [KE19]. Robust [APNT22, GC12]. Role [AJM^{+}23, BM17, ST14]. Rosser [KS13b]. Rough [KP18, KBP23]. routing [BMRS10]. Rule [CJL13, AR04, AAD11]. rule-based [AR04]. Rules [BF13, CR16, CG18, FFF12, GF18, Kou21, LMTV19, Ber04]. Runtime [BKZ20].

S4 [HK21]. S5 [HK21]. Safe [CC19, AE09, Lib07]. Safety [EFH14, Laz11, LOW16]. sampling [FR10]. SAT [BH23, HS15, LOS19, VBH23]. SAT-Inspired [VBH23]. Satisfaction [BJP17, DMM17, BMT13, Bul11, CSS10]. Satisfiability [CLZ23, CDG15, CGI^{+}18, FCGQ20, HKMV20, KT18, LMS17, Pra13, Sha22, ABRS09, BV02, CGS10, JL11]. satisfy [CHK08, CHK10]. Scenarios [CKK17]. Schaefer [MMS23]. scheduling [BERS04, GK08]. Schemas [DHL12]. Schematic [Gui22]. Schemes [BCOS21, EP13, HMOS17, Mam16].

Schönfinkel [EIP20]. SCIFF [ACG^{+}08]. Search [BB17, BGL13, CLZ23, CKK17, vHPP00, BM08, ET06, FF20, MV04, Sto05, TM10, Vor01]. Second [EGT16, FP17, AJ05, BB14, KS13a, Sze11]. Second-Order [EGT16, FP17, AJ05, BB14, KS13a]. Secrecy [TSH15]. Secrecy-preserving [TSH15]. Security [CCD15, CCDD20, AM01, CKRT08, CLCZ10]. Selection [Bur20]. Self [KE19]. Self-Stabilization [KE19]. Semantic [HKVV18, Tak23, WWT15, Ber04, LA13, MP09, EILS11]. Semantical [EFW07]. Semantics [BBDM23, BFF^{+}21, Dag22, DU21, DKDV22, GF18, GS13b, GKR18, GW22, Han18, HKMV20, LWG13, ZZ17, AAD11, BFV04, Dal09a, DH02, EILS11, GM09a, JNS^{+}06, LT03, LSS04, RW05, VGD06, VGD07, WZ05, EFL^{+}04]. semi [BMT13]. semi-lattice [BMT13]. Semicomplete [DMM17]. sensing [ILNR09, Rei01]. Sensitive [Ara23, ZG14]. Sentences [PZ22]. Separating [DD16]. Separation [DD15, DD16, DLM21, DHS16, EIP20, HCGT18, MPZ23, PZ20]. sequence [Lyn05]. sequences [KHST12]. Sequent [BO02, HCGT18, LZ19, LP19, MOG05, LA13, OPS07, SP11, Vor01]. Sequential [Gur00]. sequentiality [BK02]. Sequents [MMZ21]. Set [AZZ14, BVD16, BJN20, DSTW13, GS13b, KP18, KBP23, Kou21, WWSL16, EFW07, HVV08, OSTW21, SI08, SBTM06]. Set-Inductions [BVD16]. Sets [BGS16, CP22, WEZL23, BMT13, DFO06, DPR08, DG12, Sak05]. Several [BIK^{+}21]. Shallow [RMV17]. Shoham [BKMV^{+}17]. Short [BCMS18]. signed [Ari07]. signing [KKW10]. Simple [BCMS18, HPS22, KP18]. Simplicial [vdBG12]. Simplification [FFF12]. simply [BERS04]. simulate [IS06]. Simulation [BG03b, Hes05, Tri09]. Simulation-based [BG03b]. simulations [GV04]. Simultaneous [Bou09]. single [SV08]. singleton [Cra07, SH06]. situation [Rei01]. Size [BBMP23, NV19, AI03, AS02, GL10, KS13b, Kla08]. size-degree [GL10]. Size-width [BBMP23]. Skolem [Avi03, BHM15]. Slicing [MZ17]. Small [DKS17, FLM^{+}15, Kra21, BG06, BG07a, BG07b]. small-step [BG06, BG07a, BG07b]. SMT [BLRC^{+}19]. Social [SBSM13]. societies [ASP09]. Soft [BMR06, BMRS10]. soft-constraint [BMRS10]. software [Cha06]. Solutions [MSS14, San17]. Solvable [BMV22]. Solve [SBSM13]. Solver [BH23]. solvers [AR04]. Solving [BLRC^{+}19, SJV12, BGLC04, DPR08, ET06, KV05a, Rat04, Rat06]. Some [GK18, IS04]. Sound [AL23, BMS13, Cra07, Kou21, SDW14, AMMO09]. Space [FLM^{+}15]. Spaces [HK21, KNPHZ13, KWS^{+}03]. spatio [GHK08]. spatio-temporal [GHK08]. Spawns [CM17]. Special [CC19, GHJP03, KM10, Kol05, BMR09]. specification [CHK08, CHK10, TZ02]. Specifications [BCHK14, Bou09, Hes05, SV10, TM10]. specified [Lyn05]. Specifying [ASP09]. Spectra [KT15]. Splittable [TG18]. Splitting [VGD06, VGD07]. SSReflect [HCMS13]. ST [BG02]. Stability [LL15]. Stabilization [KE19]. Stable [ZZ17, JNS^{+}06]. standard [OPS07]. State [GMS20, NSST18, BG03a, BG08a, EFL^{+}04, EH01, Gur00, Mur05, NSV04]. Statecharts [LM02]. static [MP09]. Statistical [DHKP17]. Step [Sim15, BG06, BG07a, BG07b, Dag22]. steps [LM02]. STGLAs [Ber04]. Stochastic [CD14, HKO19, BBD03]. Strategic [BKM^{+}23, BJP19, DN21, MW15, GS09]. Strategies [CKPP12, MMPV14, GK09, MV04, vHPP00]. Strategy [BMM^{+}21, CS20, HV18, PDHR14, Sim15, KKW10]. Stratification [BCF21].

Streams [BKZ20, Sau10]. Strength

[EFH14, Nor12, CT06]. Strict

[HKMV20, Van18a, MP03]. string [EH01]. strings [NSV04]. Strong [BJN20]. Strongly [LPV01]. Structural [CR16, KRW12, Sim14, LA13]. structure [BG11b, Gug07]. Structured [ACOS17, GJ12]. Structures [EEH17b, EvdHK^{+}20, NS15, SDW14, DG07, GPW10]. Stuttering [GJKW17]. style [BCM^{+}21, GM09b]. Subatomic [BG22, TG18]. Subcube [FMSV23]. Subgraph [VZ19]. Subset [HK21]. Subsets [DKS17]. Substitution [NS15, Pie09]. Substructural [KT03]. Substructures [BMM15]. Subsystems [GK18]. Successor [PZ20]. Succinctly [MSS14]. Succinctness [EEH17b, GN12]. Sums [FMSV23, Sha22]. Sup [MP09]. Sup-interpretations [MP09]. Super [BDP04, MSS14]. Super-Solutions [MSS14]. Superposition [HW10, VBH23]. Support [CJL13]. suspension [GK08]. Switchability [CMMZ23]. Symbolic [BCEO20, Ber04, FDY14, HMR05]. Symbols [CGT17, EŠ10]. Symmetric [DW22]. Synchronizing [QS19]. Syntactic [CGPT22, DZ13]. syntactical [AS02]. Syntax [BZWS18, MM02]. Syntax-Preserving [BZWS18]. Synthesis [CDG15, GS13a, NSM18]. System [BG22, Gui22, Sim15, BG11b, CHK08, CHK10, Gug07, Gui09, LPF^{+}06]. Systems [ACOW22, ALGN18, BMP23, CG18, CLLM19, CR15, DMP18, EY12, FMZV15, GS13a, GGKT12, GL14, HKO19, KRW12, KBP23, RMV17, Sch15, TG18, YLYF14, BBS07, CNNR03, GNT04, GS09, HMR05, IS06, KB11, LA13, LY07, LvdMR00, VH05].

Tableau

[CDG15, GS13b, GGOS09, GS09, ST14]. Tableau-based [GS09]. Tableaux [FFF12, GGOP09]. tabling [VDS01]. Taming [CLSZ15, FS16]. Tautologies [CGL16]. Team [DKDV22, HKVV18]. Technique [SDW14, VH05]. Techniques [BLRC^{+}19]. Temporal [APNT22, AKRZ14, BBK^{+}20, BBDFD20, BCHK14, BMM^{+}19, BMP23, BKMV^{+}17, CDG15, DHKP17, EG22, FPS14, GKR18, GC12, HK15b, KBP23, LOW16, LMS17, MSS14, SSS12, Tak23, VP20, AELP01, BK10, DFK06, DKS06, FDP01, FR10, GL13, GHK08, SPSS11, ZHD14]. Tense [CLRT21]. Term [PR23, RMV17, CNNR03, GNT04, KS13b, Pie09, RV01, SKGST09, VH05]. Terminating [RS14]. Termination [BERS04, GK09, JR15, VDS01, BG08b, MR03, SKGST09, SYY^{+}03]. Terms [Les18, DG12, GGSS11, SP11]. Testers [AF23]. Testing [BCGH23, CHN20, CDG15, LLM^{+}07]. tests [GM09b, Koz00]. their [ACF05, CFS10, KB11]. Theorem [AG19, AMMU19, McK13, PZ20, ADGR07, BAGM10, OPS07, ACMU21, NC12]. Theoretic [DSTW13, GKR18, HPS22]. Theoretical [NV19]. Theories [AG23, CLZ23, EP13, GJL15, LW15, PS17, ST15, BGR14, CEG05, CGS10, DZ13, DLS01, MO12]. Theory [BCF21, BGM14, CHN20, CR20, GMS20, KBP23, Kra21, Kra15, Rab15, Soj16, AL10, BG11a, Dag22, DST13, FR10, GV04, HP05, Mie04, MT05, NPP08, Ser01]. there [LR06]. Three [DZ13, RS14]. Three-Valued [RS14]. Tight [BK12, VZ19, BM10]. Time [ACOW22, BCGH23, CDG15, FPS14, GKR18, PSS18, YLYF14, AS02, ASSB00, BBS07, CDKM13, FR10, GM02, GGV02, Kra21, KV05b, ZHD14, Lut04]. TIME-complete [Lut04]. Timed [CLLM19, HSTW20, MBRS20, BDR08, DGM04, LW08, Tri09]. Timed-register [CLLM19]. Tool [BCD13]. Topological [EFH14, KNPHZ13, Sal03, vdBG12, FD14, GS02]. Torus [Soj16]. Tournaments [LMM^{+}22]. tower [AD14]. tower-type [AD14]. TPTL [FCGQ20]. trace [BFV04]. Tractability [DKDV22, BMS^{+}11, BMT13]. tradeoffs [GL10]. Transcendental [CGI^{+}18]. transducers [EH01]. transductions [EH01]. Transfer [Kra15, PZ20]. transformation [LF01]. Transformations [BJN20]. Transition [CR15, DMP18, WWSL16, HMR05]. Transitive [CR20, KT18]. transitivity [BV02]. Translating [BK12]. Translation [GF18, SH07]. Tree [BL16, BBGS20, BFF^{+}21, EEH17a, GL14, BS09, BMM12, Sub04]. tree-like [Sub04]. Trees [BG22, BBC^{+}16, CHS14, CHS15, CW16, GLSG15, Kar13, BLN07, JL11, KS13a, KRS05, Mie04, Pie09, Tan14]. Treewidth [Pra13, GPW10]. True [BP20]. Truly [BJP19]. Trust [HKO19]. Twelf [CS08]. Two [ACF^{+}22, BBC^{+}16, BDM^{+}11, CW16, DD15, DD16, KT18, KS17, PR23, Soj16, BK02, EH01, Tan14]. Two-Variable [BBC^{+}16, CW16, DD15, KT18, KS17, BDM^{+}11, Tan14]. two-way [EH01]. Typal [Pit20]. Type [LW15, Rab18, Soj16, AL10, AD14, BFW03, BK02, CM06, DCHA03, HP05, NPP08, VV07]. type-two [BK02]. Typechecking [AMN^{+}03]. Typed [BBC02, FF18]. Types [Asp15, HPS22, MNPS20, Pit20, Van18a, Van18b, vdBG12, DCDGT10, IS04, SH06]. Typing [CCDD20, Sau10].

ugly [BMS^{+}11]. Unbounded [DHKP17]. Uncertain [FL20, SFL17]. uncertainty [ILNR09]. Undecidability [CFS10, LWG13, WZ08, Mur05]. Undecidable [HK15b, Kuz21, VP20, GJ12]. Underlying [EFH14]. Unfolding [JNS^{+}06]. Unicast [BMRS10]. Unification [GGSS11, LV12, CK20, SS07, WZ08]. Unified [BK12, LA13]. Uniform [GMS20, OSTW21, DPR08]. Unifying [DDDW22, DHS16]. Uninterpreted [EIP20, BGV01]. Union [GKNS19]. Unique [San17]. Uniqueness [RMV17]. unit [HV05]. unit-free [HV05]. UNITY [Pau00]. Univalent [KS15]. Universal [BJK22, Gui22, Kra21]. Universes [KS15, LW15]. Unprovability [FF20]. unranked [BLN07]. Unsatisfiable [DKS17]. Until [BL16]. Update [HLM13]. Updates [CFP^{+}19, Zha06]. Upper [AGM13, KS13b]. Usage [AGM13]. Useful [BK12]. Using [ACOS17, APNT22, CFMP19, KKLJU14, ST14, SBSM13, SPNS14, BCGH23, BGV01, Pie09, Van18a].

Vacuity [GC12]. Validity [BBGS20]. value [Ber04]. Valued [BMV22, RS14, SG07]. Values [BCHK14, Tan14]. Variable [BBC^{+}16, BKV15, CW16, DD15, HK15b, KT18, KT15, KS17, Lib03, PR23, BDM^{+}11, SV08, Tan14]. Variables [DD16, Gui22, TW23, CPV09, Hes05]. Variants [KT18]. Variation [AF23]. Varieties [AMMU19]. VASS [LS15]. Vector [CLLM19]. Verifiable [ACG^{+}08]. Verification [APNT22, BCEO20, BKZ20, BHM15, CCCK16, CDKM13, CGI^{+}18, DDDW22, FLY22, KE19, LDJ^{+}18, BGV01, MO06]. verified [ADGR07]. Verifying [AM01, BBS07, FKN17, SV08]. versus [TZ04]. Via [Asp15, BSH10, FKN17, GM02, GL14, Ham18, HP03, KHST12, MO06]. View [BFHH17, HPS22]. Views [BKONR23, AMN^{+}03, BDT10]. Visible [BBT^{+}21, BJN20]. Visibly [BMP23, FKL15]. vs [BMM^{+}19].

way [EH01]. Weak [CFVZ20, DDDW22, Kra21, KV01, BG02, CT06]. Weakness [EFH14]. Web [AAD11, EILS11]. Weighted [AK15, BGMZ14, CHO17]. Well [BVD16, EILS11, GJ12, BG08b, CR20, WZ05]. Well-Founded [BVD16, EILS11, WZ05]. Well-structured [GJ12]. Weyl [AL10]. Where [EGT16, Lib07]. Width [FLM^{+}15, GN23, HS15, BBMP23]. Win [CD14]. winning [CHH09]. Wise [NSM18]. within [CPRW16, HCMS13]. without [AJM^{+}23, MM12]. Witnesses [WEZL23]. witnessing [BB14]. Word [BM17]. Words [QS19, BDM^{+}11, Laz11]. world [AB20]. Worst [NV19]. Worst-Case [NV19].

XML [AMN^{+}03, HGS07]. XPath

[Fig12, JL11].

YAPA [BCD13].

Zeno [LOW16]. Zero [PZ22].

References

Analyti:2011:MPF

[AAD11] Anastasia Analyti, Grigoris Antoniou, and Carlos Viegas Damasio. MWeb: a principled framework for modular Web rule bases and its semantics. ACM Transactions on Computational Logic, 12(2):17:1– 17:??, January 2011. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Amarilli:2020:FOW

[AB20] Antoine Amarilli and Michael Benedikt. Finite open-world query answering with number restrictions. ACM Transactions on Computational Logic, 21(4):27:1–27:73, October 2020. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

                                                     https://dl.acm.org/doi/
                                                     10.1145/3365834.

Angiulli:2003:CPM

[ABEZIP03] Fabrizio Angiulli, Rachel Ben-Eliyahu-Zohary, Giovambattista Ianni, and Luigi Palopoli. Computational properties of metaquerying problems. ACM Transactions on Computational Logic, 4(2): 149–180, April 2003. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Antoniou:2001:RRD

[ABGM01] Grigoris Antoniou, David Billington, Guido Governatori, and Michael J. Maher. Representation results for defeasible logic. ACM Transactions on Computational Logic, 2(2):255–287, 2001. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

http://www.acm.org/pubs/
articles/journals/tocl/
2001-2-2/p255-antoniou/
p255-antoniou.pdf;   http:
//www.acm.org/pubs/citations/
journals/tocl/2001-2-2/
p255-antoniou/.

Arenas:2021:CCP

[ABM21] Marcelo Arenas, Pablo Barceló, and Mikaël Monet. The complexity of counting problems over incomplete databases. ACM Transactions on Computational Logic, 22 (4):21:1–21:52, October 2021. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3461642.

Armando:2009:NRR

[ABRS09] Alessandro Armando, Maria Paola Bonacina, Silvio Ranise, and Stephan Schulz. New results on rewrite-based satisfiability procedures. ACM Transactions on Computational Logic, 10(1):4:1–4:??, January 2009. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Aamer:2023:IOC

[ABS^{+}23] Heba Aamer, Bart Bogaerts, Dimitri Surinx, Eugenia Ternovska, and Jan Van den Bussche. Inputs, outputs, and composition in the logic of information flows. ACM Transactions on Computational Logic, 24(4):33:1–33:??, October 2023. CODEN ???? ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3604553.

Afrati:2005:DPT

[ACF05] Foto Afrati, Stavros Cosmadakis, and Eugénie Foustoucos. Datalog programs and their persistency numbers. ACM Transactions on Computational Logic, 6(3): 481–518, July 2005. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Aceto:2022:TBO

[ACF^{+}22] Luca Aceto, Valentina Castiglioni, Wan Fokkink, Anna Ingólfsdóttir, and Bas Luttik. Are two binary operators necessary to obtain a finite axiomatisation of parallel composition? ACM Transactions on Computational Logic, 23 (4):22:1–22:??, October 2022. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3529535.

Alberti:2008:VAI

[ACG^{+}08] Marco Alberti, Federico Chesani, Marco Gavanelli, Evelina Lamma, Paola Mello, and Paolo Torroni. Verifiable agent interaction in abductive logic programming: The SCIFF framework. ACM Transactions on Computational Logic, 9(4):29:1– 29:??, August 2008. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Albert:2018:PCA

[ACJ^{+}18] Elvira Albert, Jesús Correas, Einar Broch Johnsen, Ka I. Pun, and Guillermo Román-Díez. Parallel cost analysis. ACM Transactions on Computational Logic, 19(4): 31:1–31:??, December 2018. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3274278.

Adamek:2021:RTF

[ACMU21] Jirí Adámek, Liang-Ting Chen, Stefan Milius, and Henning Urbat. Reiterman’s Theorem on finite algebras for a monad. ACM Transactions on Computational Logic, 22 (4):23:1–23:48, October 2021. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3464691.

Ahmetaj:2017:MCG

[ACOS17] Shqiponja Ahmetaj, Diego Calvanese, Magdalena Ortiz, and Mantas Simkus. Managing change in graphstructured data using description logics. ACM Transactions on Computational Logic, 18(4):27:1–27:??, December 2017. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Almagor:2022:MID

[ACOW22] Shaull Almagor, Dmitry Chistikov, Joël Ouaknine, and James Worrell. O-minimal invariants for discrete-time dynamical systems. ACM Transactions on Computational Logic, 23(2):9:1–9:20, April 2022. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3501299.

Alur:2012:AAA

[ACW12] Rajeev Alur, Pavol Cerný, and Scott Weinstein. Algorithmic analysis of arrayaccessing programs. ACM Transactions on Computational Logic, 13(3):27:1–27:??, August 2012. CODEN AT-CLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Atserias:2014:DLB

[AD14] Albert Atserias and Anuj Dawar. Degree lower bounds of tower-type for approximating formulas with parity quantifiers. ACM Transactions on Computational Logic, 15(1): 6:1–6:??, February 2014. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Avigad:2007:FVP

[ADGR07] Jeremy Avigad, Kevin Donnelly, David Gray, and Paul Raff. A formally verified proof of the prime number theorem. ACM Transactions on Computational Logic, 9(1): 2:1–2:23, December 2007. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Aschinger:2014:LLC

[ADGV14] Markus Aschinger, Conrad Drescher, Georg Gottlob, and Heribert Vollmer. LoCo — a logic for configuration problems. ACM Transactions on Computational Logic, 15(3): 20:1–20:??, September 2014. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Arai:2009:NFA

[AE09] Toshiyasu Arai and Naohi Eguchi. A new function algebra of EXPTIME functions by safe nested recursion. ACM Transactions on Computational Logic, 10(4):24:1– 24:??, August 2009. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Alur:2001:PTL

[AELP01] Rajeev Alur, Kousha Etessami, Salvatore La Torre, and Doron Peled. Parametric temporal logic for “model measuring”. ACM Transactions on Computational Logic, 2(3):388–407, 2001. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

http://www.acm.org/pubs/
citations/journals/tocl/
2001-2-3/p388-alur/.

Adler:2023:FPT

[AF23] Isolde Adler and Polly Fahey. Faster property testers in a variation of the bounded degree model. ACM Transactions on Computational Logic, 24(3):25:1–25:??, July 2023. CODEN ???? ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3584948.

Aceto:2009:FEB

[AFIL09] Luca Aceto, Wan Fokkink, Anna Ingolfsdottir, and Bas Luttik. A finite equational base for CCS with left merge and communication merge. ACM Transactions on Computational Logic, 10(1):6:1– 6:??, January 2009. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Albert:2016:MHP

[AFMGMM16] Elvira Albert, Antonio Flores-Montoya, Samir Genaim, and Enrique Martin-Martin. Mayhappen-in-parallel analysis for actor-based concurrency. ACM Transactions on Computational Logic, 17(2):11:1– 11:??, March 2016. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Aameri:2019:RTC

[AG19] Bahar Aameri and Michael Grüninger. A representation theorem for change through composition of activities. ACM Transactions on Computational Logic, 20(4): 20:1–20:??, September 2019. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3329121.

Aameri:2023:RTA

[AG23] Bahar Aameri and Michael Grüninger. Reducible theories and amalgamations of models. ACM Transactions on Computational Logic, 24 (1):9:1–9:??, January 2023. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3565364.

Albert:2013:IRU

[AGM13] Elvira Albert, Samir Genaim, and Abu Naser Masud. On the inference of resource usage upper and lower bounds. ACM Transactions on Computational Logic, 14(3):22:1– 22:??, August 2013. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Angiulli:2007:ODL

[AGP07] Fabrizio Angiulli, Gianluigi Greco, and Luigi Palopoli. Outlier detection by logic programming. ACM Transactions on Computational Logic, 9(1):7:1–7:??, December 2007. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Adler:2003:LBF

[AI03] Micah Adler and Neil Immerman. An n! lower bound on formula size. ACM Transactions on Computational Logic, 4(3):296–314, July 2003. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Aehlig:2005:EFS

[AJ05] Klaus Aehlig and Jan Johannsen. An elementary fragment of second-order lambda calculus. ACM Transactions on Computational Logic, 6(2): 468–480, April 2005. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Artale:2023:LBC

[AJM^{+}23] Alessandro Artale, Jean Christoph Jung, Andrea Mazzullo, Ana Ozaki, and Frank Wolter. Living without Beth and Craig: Definitions and interpolants in description and modal logics with nominals and role inclusions. ACM Transactions on Computational Logic, 24(4):34:1–34:??, October 2023. CODEN ???? ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3597301.

Avni:2015:PWC

[AK15] Guy Avni and Orna Kupferman. Parameterized weighted containment. ACM Transactions on Computational Logic, 16(1):6:1–6:??, March 2015. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Ameloot:2017:DQD

[AKNZ17] Tom J. Ameloot, Bas Ketsman, Frank Neven, and Daniel Zinn. Datalog queries distributing over components. ACM Transactions on Computational Logic, 18(1):5:1– 5:??, April 2017. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Artale:2014:CTC

[AKRZ14] Alessandro Artale, Roman Kontchakov, Vladislav Ryzhikov, and Michael Zakharyaschev. A cookbook for temporal conceptual data modelling with description logics. ACM Transactions on Computational Logic, 15(3):25:1–25:??, September 2014. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Apt:2001:E

[AKS01] Krzysztof R. Apt, Antonis C. Kakas, and Fariba Sadri. Editorial. ACM Transactions on Computational Logic, 2 (4):431, October 2001. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Alferes:2013:QDP

[AKS13] José Júlio Alferes, Matthias Knorr, and Terrance Swift. Query-driven procedures for hybrid MKNF knowledge bases. ACM Transactions on Computational Logic, 14(2): 16:1–16:??, June 2013. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Alur:2004:DGG

[AL04] Rajeev Alur and Salvatore La Torre. Deterministic generators and games for Ltl fragments. ACM Transactions on Computational Logic, 5(1): 1–25, January 2004. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Adams:2010:WPC

[AL10] Robin Adams and Zhaohui Luo. Weyl’s predicative classical mathematics as a logic-enriched type theory. ACM Transactions on Computational Logic, 11(2):11:1– 11:??, January 2010. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Atserias:2023:CYS

[AL23] Albert Atserias and Massimo Lauria. Circular (yet sound) proofs in propositional logic. ACM Transactions on Computational Logic, 24(3):20:1– 20:??, July 2023. CODEN ???? ISSN 1529-3785 (print), 1557-945X (electronic). URL

https://dl.acm.org/doi/
10.1145/3579997.

Ardeshir-Larijani:2018:AEC

[ALGN18] Ebrahim Ardeshir-Larijani, Simon J. Gay, and Rajagopal Nagarajan. Automated equivalence checking of concurrent quantum systems. ACM Transactions on Computational Logic, 19(4): 28:1–28:??, December 2018. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3231597.

Atserias:2016:NPM

[ALN16] Albert Atserias, Massimo Lauria, and Jakob Nordström. Narrow proofs may be maximally long. ACM Transactions on Computational Logic, 17(3):19:1–19:??, July 2016. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Abadi:2002:E

[ALP02] Martín Abadi, Leonid Libkin, and Frank Pfenning. Editorial. ACM Transactions on Computational Logic, 3 (3):335, July 2002. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Aiello:2001:VSP

[AM01] Luigia Carlucci Aiello and Fabio Massacci. Verifying security protocols as planning in logic programming. ACM Transactions on Computational Logic, 2(4):542– 580, October 2001. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Arthan:2009:GFS

[AMMO09] Rob Arthan, Ursula Martin, Erik A. Mathiesen, and Paulo Oliva. A general framework for sound and complete Floyd-Hoare logics. ACM Transactions on Computational Logic, 11(1):7:1–7:??, October 2009. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Adamek:2019:GET

[AMMU19] Jirí Adámek, Stefan Milius, Robert S. R. Myers, and Henning Urbat. Generalized Eilenberg theorem: Varieties of languages in a category. ACM Transactions on Computational Logic, 20 (1):3:1–3:??, February 2019. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3276771.

Alon:2003:TXV

[AMN^{+}03] Noga Alon, Tova Milo, Frank Neven, Dan Suciu, and Victor Vianu. Typechecking XML views of relational databases. ACM Transactions on Computational Logic, 4(3):315–354, July 2003. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Atserias:2019:PCM

[AO19] Albert Atserias and Joanna Ochremiak. Proof complexity meets algebra. ACM Transactions on Computational Logic, 20(1):1:1–1:??, February 2019. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3265985.

Anevlavis:2022:BCE

[APNT22] Tzanis Anevlavis, Matthew Philippe, Daniel Neider, and Paulo Tabuada. Being correct is not enough: Efficient verification using robust linear temporal logic. ACM Transactions on Computational Logic, 23(2):8:1–8:39, April 2022. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

https://dl.acm.org/doi/
10.1145/3491216.

Apt:2005:E

[Apt05]  Krzysztof R. Apt.  Edi-

torial. ACM Transactions on Computational Logic, 6 (4):673, October 2005. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Asperti:2002:ILA

[AR02] Andrea Asperti and Luca Roversi. Intuitionistic Light Affine Logic. ACM Transactions on Computational Logic, 3(1):137–175, January 2002. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Abdennadher:2004:AGR

[AR04] Slim Abdennadher and Christophe Rigotti. Automatic generation of rule-based constraint solvers over finite domains. ACM Transactions on Computational Logic, 5(2): 177–205, April 2004. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Aravanis:2023:GPC

[Ara23] Theofanis Aravanis. Generalizing Parikh’s criterion for relevance-sensitive belief revision. ACM Transactions on Computational Logic, 24 (2):18:1–18:??, April 2023. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3572907.

Arieli:2007:PRP

[Ari07] Ofer Arieli. Paraconsistent reasoning and preferential entailments by signed quantified Boolean formulae. ACM Transactions on Computational Logic, 8(3):18:1–18:??, July 2007. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Aehlig:2002:SAN

[AS02] Klaus Aehlig and Helmut Schwichtenberg. A syntactical analysis of non-size-increasing polynomial time computation. ACM Transactions on Computational Logic, 3(3):

383–401, July 2002. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Artikis:2009:SNG

[ASP09] Alexander Artikis, Marek Sergot, and Jeremy Pitt. Specifying norm-governed computational societies. ACM Transactions on Computational Logic, 10(1):1:1–1:??, January 2009. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Asperti:2015:CCF

[Asp15] Andrea Asperti. Computational complexity via finite types. ACM Transactions on Computational Logic, 16(3): 26:1–26:??, July 2015. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Aziz:2000:MCC

[ASSB00] Adnan Aziz, Kumud Sanwal, Vigyan Singhal, and Robert Brayton. Model-checking continuous-time Markov chains. ACM Transactions on Computational Logic, 1(1):162– 170, July 2000. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/tocl/2000-1-1/
                                        p162-aziz/.

Audrito:2019:HOC

[AVD^{+}19] Giorgio Audrito, Mirko Viroli, Ferruccio Damiani, Danilo Pianini, and Jacob Beal. A higher-order calculus of computational fields. ACM Transactions on Computational Logic, 20(1):5:1–5:??, February 2019. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3285956.

Avigad:2003:EDS

[Avi03] Jeremy Avigad. Eliminating definitions and Skolem functions in first-order logic. ACM Transactions on Computational Logic, 4(3):402–415, July 2003. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Asuncion:2014:PFO

[AZZ14] Vernon Asuncion, Yan Zhang, and Yi Zhou. Preferred firstorder answer set programs. ACM Transactions on Computational Logic, 15(2):11:1– 11:??, April 2014. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Baelde:2012:LGF

[Bae12] David Baelde. Least and greatest fixed points in linear logic. ACM Transactions on Computational Logic, 13(1): 2:1–2:??, January 2012. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Billington:2010:ITD

[BAGM10] David Billington, Grigoris Antoniou, Guido Governatori, and Michael Maher. An inclusion theorem for defeasible logics. ACM Transactions on Computational Logic, 12(1): 6:1–6:??, October 2010. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Basaldella:2021:RAM

[Bas21]  Michele Basaldella.  αβ-

relations and the actual meaning of α-renaming. ACM Transactions on Computational Logic, 22(1):4:1–4:32, January 2021. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3426471.

Bauer:2014:PCF

[Bau14]  Matthew S. Bauer.  A

PSPACE-complete first-order fragment of computability logic. ACM Transactions on Computational Logic, 15(1): 1:1–1:??, February 2014. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Beckmann:2014:IWL

[BB14] Arnold Beckmann and Samuel R. Buss. Improved witnessing and local improvement principles for second-order bounded arithmetic. ACM Transactions on Computational Logic, 15(1):2:1–2:??, February 2014. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Beckmann:2017:NSP

[BB17] Arnold Beckmann and Sam Buss. The NP search problems of Frege and extended Frege proofs. ACM Transactions on Computational Logic, 18(2):11:1–11:??, June 2017. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Beyersdorff:2020:DQD

[BB20] Olaf Beyersdorff and Joshua Blinkhorn. Dynamic QBF dependencies in reduction and expansion. ACM Transactions on Computational Logic, 21(2):8:1–8:27, March 2020. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3355995.

Bono:2002:TIE

[BBC02] Viviana Bono, Michele Bugliesi, and Silvia Crafa. Typed interpretations of extensible objects. ACM Transactions on Computational Logic, 3(4): 562–603, October 2002. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Benaim:2016:CTV

[BBC^{+}16] Saguy Benaim, Michael Benedikt, Witold Charatonik, Emanuel Kieroński, Rastislav Lenhardt, Filip Mazowiecki, and James Worrell. Complexity of twovariable logic on finite trees. ACM Transactions on Computational Logic, 17(4):32:1– 32:??, November 2016. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bryans:2003:MCS

[BBD03] Jeremy Bryans, Howard Bowman, and John Derrick. Model checking stochastic automata. ACM Transactions on Computational Logic, 4(4): 452–492, October 2003. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Balbiani:2020:ILT

[BBDFD20] Philippe Balbiani, Joseph Boudou, Martín Diéguez, and David Fernández-Duque. Intuitionistic linear temporal logics. ACM Transactions on Computational Logic, 21(2): 14:1–14:32, March 2020. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3365833.

Bellier:2023:GGQ

[BBDM23] Dylan Bellier, Massimo Benerecetti, Dario Della Monica, and Fabio Mogavero. Goodfor-game QPTL: an alternating Hodges semantics. ACM Transactions on Computational Logic, 24(1):4:1– 4:??, January 2023. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https://

dl.acm.org/doi/10.1145/
3565365.

Benedikt:2020:MDT

[BBGS20] Michael Benedikt, Pierre Bourhis, Georg Gottlob, and Pierre Senellart. Monadic Datalog, tree validity, and limited access containment. ACM Transactions on Computational Logic, 21(1):6:1– 6:45, January 2020. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bonchi:2014:ACD

[BBH^{+}14] Filippo Bonchi, Marcello M.

Bonsangue, Helle H. Hansen, Prakash Panangaden, Jan J. M. M. Rutten, and Alexandra Silva. Algebra-coalgebra duality in Brzozowski’s minimization algorithm. ACM Transactions on Computational Logic, 15(1):3:1–3:??, February 2014. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Baader:2020:MTD

[BBK^{+}20] Franz Baader, Stefan Borgwardt, Patrick Koopmann, Ana Ozaki, and Veronika Thost. Metric temporal description logics with intervalrigid names. ACM Transactions on Computational Logic, 21(4):30:1–30:46, October 2020. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

https://dl.acm.org/doi/
10.1145/3399443.

Beyersdorff:2023:HCS

[BBMP23] Olaf Beyersdorff, Joshua Blinkhorn, Meena Mahajan, and Tomás Peitl. Hardness characterisations and sizewidth lower bounds for QBF resolution. ACM Transactions on Computational Logic, 24(2):10:1–10:??, April 2023. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3565286.

Baier:2007:VNP

[BBS07] Christel Baier, Nathalie Bertrand, and Philippe Schnoebelen. Verifying nondeterministic probabilistic channel systems against ω-regular linear-time properties. ACM Transactions on Computational Logic, 9(1):5:1–5:??, December 2007. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Benedikt:2021:IVI

[BBT^{+}21] Michael Benedikt, Pierre Bourhis, Balder Ten Cate, Gabrieled Puppis, and Michael Vanden Boom. Inference from visible information and background knowledge. ACM Transactions on Computational Logic, 22(2):13:1–13:69, June 2021. CODEN AT-CLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https://

dl.acm.org/doi/10.1145/
3452919.

Baudet:2013:YGT

[BCD13] Mathieu Baudet, Véronique Cortier, and Stéphanie Delaune. YAPA: a generic tool for computing intruder knowledge. ACM Transactions on Computational Logic, 14(1): 4:1–4:??, February 2013. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bana:2020:VMC

[BCEO20] Gergei Bana, Rohit Chadha, Ajay Kumar Eeralla, and Mitsuhiro Okada. Verification methods for the computationally complete symbolic attacker based on indistinguishability. ACM Transactions on Computational Logic, 21(1):2:1–2:44, January 2020. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bogaerts:2021:SAF

[BCF21] Bart Bogaerts and Luís Cruz-Filipe. Stratification in approximation fixpoint theory and its application to active integrity constraints. ACM Transactions on Computational Logic, 22(1):6:1– 6:19, January 2021. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https://

dl.acm.org/doi/10.1145/
3430750.

Baxter:2023:TUC

[BCGH23] James Baxter, Ana Cavalcanti, Maciej Gazda, and Robert M. Hierons. Testing using CSP models: Time, inputs, and outputs. ACM Transactions on Computational Logic, 24(2):17:1–17:??, April 2023. CODEN AT-CLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https://

dl.acm.org/doi/10.1145/
3572837.

Broadbent:2021:CPP

[BCH^{+}21] Christopher H. Broadbent, Arnaud Carayol, Matthew Hague, Andrzej S. Murawski, C.-H. Luke Ong, and Olivier Serre. Collapsible pushdown parity games. ACM Transactions on Computational Logic, 22(3):16:1–16:51, June 2021. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3457214.

Boker:2014:TSA

[BCHK14] Udi Boker, Krishnendu Chatterjee, Thomas A. Henzinger, and Orna Kupferman. Temporal specifications with accumulative values. ACM Transactions on Computational Logic, 15(4): 27:1–27:??, September 2014. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bortolussi:2019:CLM

[BCKL19] Luca Bortolussi, Luca Cardelli, Marta Kwiatkowska, and Luca Laurenti. Central limit model checking. ACM Transactions on Computational Logic, 20(4):19:1–19:??, September 2019. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3331452.

Basin:2004:RMF

[BCM04] David Basin, Manuel Clavel, and José Meseguer. Reflective metalogical frameworks. ACM Transactions on Computational Logic, 5 (3):528–576, July 2004. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bordeaux:2009:GCO

[BCM09] Lucas Bordeaux, Marco Cadoli, and Toni Mancini. Generalizing consistency and other constraint properties to quantified constraints. ACM Transactions on Computational Logic, 10(3):17:1– 17:??, April 2009. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Britz:2021:PKS

[BCM^{+}21] Katarina Britz, Giovanni Casini, Thomas Meyer, Kody Moodley, Uli Sattler, and Ivan Varzinczak. Principles of KLM-style defeasible description logics. ACM Transactions on Computational Logic, 22(1):1:1–1:46, January 2021. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3420258.

Beyersdorff:2018:SPN

[BCMS18] Olaf Beyersdorff, Leroy Chew, Meena Mahajan, and Anil Shukla. Are short proofs narrow? QBF resolution is not so simple. ACM Transactions on Computational Logic, 19(1): 1:1–1:??, February 2018. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Broadbent:2021:HOR

[BCOS21] Christopher H. Broadbent, Arnaud Carayol, C.-H. Luke Ong, and Olivier Serre. Higher-order recursion schemes and collapsible pushdown automata: Logical properties. ACM Transactions on Computational Logic, 22 (2):12:1–12:37, June 2021. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3452917.

Barmpalias:2017:PCO

[BCP17] George Barmpalias, Douglas Cenzer, and Christopher P. Porter. The probability of a computable output from a random oracle. ACM Transactions on Computational Logic, 18(3):18:1–18:??, August 2017. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Biernacka:2007:CFE

[BD07a] Małgorzata Biernacka and Olivier Danvy. A concrete framework for environment machines. ACM Transactions on Computational Logic, 9(1): 6:1–6:??, December 2007. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bonacina:2007:ACI

[BD07b] Maria Paola Bonacina and Nachum Dershowitz. Abstract canonical inference. ACM Transactions on Computational Logic, 8(1):??, January 2007. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Berardi:2012:IRN

[Bd12] Stefano Berardi and Ugo de’Liguoro. Interactive realizers: a new approach to program extraction from nonconstructive proofs. ACM Transactions on Computational Logic, 13(2):11:1–11:??, April 2012. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bednarczyk:2023:CFF

[BDFM23] Bartosz Bednarczyk, Stéphane Demri, Raul Fervari, and Alessio Mansutti. On composing finite forests with modal logics. ACM Transactions on Computational Logic, 24 (2):12:1–12:??, April 2023. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3569954.

Bojanczyk:2011:TVL

[BDM^{+}11] Mikołaj Bojańczyk, Claire David, Anca Muscholl, Thomas Schwentick, and Luc Segoufin. Two-variable logic on data words. ACM Transactions on Computational Logic, 12(4): 27:1–27:??, July 2011. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Brass:2004:SLP

[BDP04] Stefan Brass, Jürgen Dix, and Teodor C. Przymusinski. Super logic programs. ACM Transactions on Computational Logic, 5(1):129– 176, January 2004. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bruyere:2008:DPM

[BDR08] Véronique Bruyère, Emmanuel Dall’olio, and Jean-François Raskin. Durations and parametric modelchecking in timed automata. ACM Transactions on Computational Logic, 9(2):12:1– 12:??, March 2008. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bailey:2010:LQV

[BDT10] James Bailey, Guozhu Dong, and Anthony Widjaja To. Logical queries over views: Decidability and expressiveness. ACM Transactions on Computational Logic, 11(2): 8:1–8:??, January 2010. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bernardo:2004:SSR

[Ber04] Marco Bernardo. Symbolic semantic rules for producing compact STGLAs from value passing process descriptions. ACM Transactions on Computational Logic, 5(3): 436–469, July 2004. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bossi:2004:TSM

[BERS04] Annalisa Bossi, Sandro Etalle, Sabina Rossi, and Jan-Georg Smaus. Termination of simply moded logic programs with dynamic scheduling. ACM Transactions on Computational Logic, 5(3):470–507, July 2004. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Betz:2013:LLB

[BF13] Hariolf Betz and Thom Frühwirth. Linear-logic based analysis of constraint handling rules with disjunction. ACM Transactions on Computational Logic, 14(1):1:1– 1:??, February 2013. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Berthon:2021:ATA

[BFF^{+}21] Raphaël Berthon, Nathanaël Fijalkow, Emmanuel Filiot, Shibashis Guha, Bastien Maubert, Aniello Murano, Laureline Pinault, Sophie Pinchinat, Sasha Rubin, and Olivier Serre. Alternating tree automata with qualitative semantics. ACM Transactions on Computational Logic, 22 (1):7:1–7:24, January 2021. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3431860.

Blondin:2017:LVC

[BFHH17] Michael Blondin, Alain Finkel, Christoph Haase, and Serge Haddad. The logical view on continuous Petri nets. ACM Transactions on Computational Logic, 18(3):24:1– 24:??, August 2017. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bloom:2004:PFD

[BFV04] Bard Bloom, Wan Fokkink, and Rob J. Van Glabbeek. Precongruence formats for decorated trace semantics. ACM Transactions on Computational Logic, 5(1):26– 78, January 2004. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Buneman:2003:IBP

[BFW03] Peter Buneman, Wenfei Fan, and Scott Weinstein. Interaction between path and type constraints. ACM Transactions on Computational Logic, 4(4):530–577, October 2003. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Blass:2001:ICL

[BG01] Andreas Blass and Yuri Gurevich. Inadequacy of computable loop invariants. ACM Transactions on Computational Logic, 2(1):1–11, 2001. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

http://www.acm.org/pubs/
articles/journals/tocl/
2001-2-1/p1-blass/p1-blass.
pdf;   http://www.acm.org/

                                        pubs/citations/journals/
                                        tocl/2001-2-1/p1-blass/.

Bravetti:2002:DAW

[BG02] Mario Bravetti and Roberto Gorrieri. Deciding and axiomatizing weak ST bisimulation for a process algebra with recursion and action refinement. ACM Transactions on Computational Logic, 3(4): 465–520, October 2002. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Blass:2003:ASM

[BG03a] Andreas Blass and Yuri Gurevich. Abstract state machines capture parallel algorithms. ACM Transactions on Computational Logic, 4(4): 578–651, October 2003. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bustan:2003:SBM

[BG03b] Doron Bustan and Orna Grumberg. Simulation-based minimization. ACM Transactions on Computational Logic, 4(2):181–206, April 2003. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Blass:2006:OIS

[BG06] Andreas Blass and Yuri Gurevich. Ordinary interactive small-step algorithms, I. ACM Transactions on Computational Logic, 7(2):

363–419, April 2006. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Blass:2007:OISa

[BG07a] Andreas Blass and Yuri Gurevich. Ordinary interactive small-step algorithms, II. ACM Transactions on Computational Logic, 8(3): 15:1–15:??, July 2007. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Blass:2007:OISb

[BG07b] Andreas Blass and Yuri Gurevich. Ordinary interactive small-step algorithms, III. ACM Transactions on Computational Logic, 8(3):16:1– 16:??, July 2007. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Blass:2008:ASM

[BG08a] Andreas Blass and Yuri Gurevich. Abstract state machines capture parallel algorithms: Correction and extension. ACM Transactions on Computational Logic, 9(3): 19:1–19:??, June 2008. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Blass:2008:PTW

[BG08b] Andreas Blass and Yuri Gurevich. Program termination and well partial orderings.

ACM Transactions on Computational Logic, 9(3):18:1– 18:??, June 2008. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bruscoli:2009:PCD

[BG09] Paola Bruscoli and Alessio Guglielmi. On the proof complexity of deep inference. ACM Transactions on Computational Logic, 10(2):14:1– 14:??, February 2009. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Blass:2011:PQB

[BG11a] Andreas Blass and Yuri Gurevich. Persistent queries in the behavioral theory of algorithms. ACM Transactions on Computational Logic, 12 (2):16:1–16:??, January 2011. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Burger:2011:SIS

[BG11b] Lutz Straß Burger and Alessio Guglielmi. A system of interaction and structure IV: The exponentials and decomposition. ACM Transactions on Computational Logic, 12(4): 23:1–23:??, July 2011. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Barrett:2022:SPS

[BG22] Chris Barrett and Alessio Guglielmi. A subatomic proof system for decision trees. ACM Transactions on Computational Logic, 23 (4):26:1–26:??, October 2022. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3545116.

Baader:2012:LDL

[BGL12] Franz Baader, Silvio Ghilardi, and Carsten Lutz. LTL over description logic axioms. ACM Transactions on Computational Logic, 13(3):21:1– 21:??, August 2012. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Beyersdorff:2013:PCD

[BGL13] Olaf Beyersdorff, Nicola Galesi, and Massimo Lauria. Parameterized complexity of DPLL search procedures. ACM Transactions on Computational Logic, 14(3): 20:1–20:??, August 2013. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Benhamou:2004:ICS

[BGLC04] Frédéric Benhamou, Frédéric Goualard, Éric Languénou, and Marc Christie. Interval constraint solving for camera control and motion planning. ACM Transactions on Computational Logic, 5(4): 732–767, October 2004. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bonchi:2014:GTB

[BGM14] Filippo Bonchi, Fabio Gadducci, and Giacoma Valentina Monreale. A general theory of barbs, contexts, and labels. ACM Transactions on Computational Logic, 15(4):35:1– 35:??, November 2014. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bollig:2014:PWA

[BGMZ14] Benedikt Bollig, Paul Gastin, Benjamin Monmege, and Marc Zeitoun. Pebble weighted automata and weighted logics. ACM Transactions on Computational Logic, 15(2): 15:1–15:??, April 2014. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bruttomesso:2014:QFI

[BGR14] Roberto Bruttomesso, Silvio Ghilardi, and Silvio Ranise. Quantifier-free interpolation in combinations of equality interpolating theories. ACM Transactions on Computational Logic, 15(1):5:1– 5:??, February 2014. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bova:2016:MCE

[BGS16] Simone Bova, Robert Ganian, and Stefan Szeider. Model checking existential logic on partially ordered sets. ACM Transactions on Computational Logic, 17(2):10:1– 10:??, March 2016. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bryant:2001:PVU

[BGV01] Randal E. Bryant, Steven German, and Miroslav N. Velev. Processor verification using efficient reductions of the logic of uninterpreted functions to propositional logic. ACM Transactions on Computational Logic, 2(1):93–134, 2001. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

http://www.acm.org/pubs/
articles/journals/tocl/
2001-2-1/p93-bryant/p93-
bryant.pdf;   http://www.
acm.org/pubs/citations/
journals/tocl/2001-2-1/
p93-bryant/.

Bryant:2023:GER

[BH23] Randal E. Bryant and Marijn J. H. Heule. Generating extended resolution proofs with a BDD-based SAT solver. ACM Transactions on Computational Logic, 24(4):31:1– 31:??, October 2023. CODEN ???? ISSN 1529-3785 (print), 1557-945X (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3595295.

Biscaia:2015:DAS

[BHM15] M. Biscaia, D. Henriques, and P. Mateus. Decidability of approximate Skolem problem and applications to logical verification of dynamical properties of Markov chains. ACM Transactions on Computational Logic, 16(1):4:1– 4:??, March 2015. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Buss:2021:LBO

[BIK^{+}21] Sam Buss, Dmitry Itsykson, Alexander Knop, Artur Riazanov, and Dmitry Sokolov. Lower bounds on OBDD proofs with several orders. ACM Transactions on Computational Logic, 22 (4):26:1–26:30, October 2021. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3468855.

Barthe:2022:UEM

[BJK22] Gilles Barthe, Charlie Jacomme, and Steve Kremer. Universal equivalence and majority of probabilistic programs over finite fields. ACM Transactions on Computational Logic, 23 (1):5:1–5:42, January 2022. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3487063.

Bomanson:2020:AVS

[BJN20] Jori Bomanson, Tomi Janhunen, and Ilkka Niemelä. Applying visible strong equivalence in answer-set program transformations. ACM Transactions on Computational Logic, 21(4):33:1–33:41, October 2020. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

https://dl.acm.org/doi/
10.1145/3412854.

Bodirsky:2017:CPC

[BJP17] Manuel Bodirsky, Peter Jonsson, and Trung Van Pham. The complexity of phylogeny constraint satisfaction problems. ACM Transactions on Computational Logic, 18(3): 23:1–23:??, August 2017. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bulling:2019:RAS

[BJP19] Nils Bulling, Wojciech Jamroga, and Matei Popovici. Reasoning about strategic abilities: Agents with truly perfect recall. ACM Transactions on Computational Logic, 20(2):10:1–10:??, April 2019. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3309761.

Boigelot:2005:EDP

[BJW05] Bernard Boigelot, Sébastien Jodogne, and Pierre Wolper. An effective decision procedure for linear arithmetic over the integers and reals. ACM Transactions on Computational Logic, 6(3): 614–633, July 2005. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Buss:2002:RBC

[BK02] Samuel R. Buss and Bruce M. Kapron. Resource-bounded continuity and sequentiality for type-two functionals. ACM Transactions on Computational Logic, 3(3): 402–417, July 2002. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bodirsky:2010:FAD

[BK10] Manuel Bodirsky and Jan Kára. A fast algorithm and datalog inexpressibility for temporal reasoning. ACM Transactions on Computational Logic, 11(3):15:1–15:??, May 2010. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Boker:2012:TCB

[BK12] Udi Boker and Orna Kupferman. Translating to Co– Büchi made tight, unified, and useful. ACM Transactions on Computational Logic, 13 (4):29:1–29:??, October 2012. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bouyer:2023:RAQ

[BKM^{+}23] Patricia Bouyer, Orna Kupferman, Nicolas Markey, Bastien Maubert, Aniello Murano, and Giuseppe Perelli. Reasoning about quality and fuzziness of strategic behaviors. ACM Transactions on Computational Logic, 24(3):21:1– 21:??, July 2023. CODEN ???? ISSN 1529-3785 (print), 1557-945X (electronic). URL

https://dl.acm.org/doi/
10.1145/3582498.

Bresolin:2017:HFH

[BKMV^{+}17] Davide Bresolin, Agi Kurucz, Emilio Muñoz-Velasco, Vladislav Ryzhikov, Guido Sciavicco, and Michael Zakharyaschev. Horn fragments of the Halpern–Shoham interval temporal logic. ACM Transactions on Computational Logic, 18(3):22:1–22:??, August 2017. CODEN AT-CLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Benedikt:2023:MDR

[BKONR23] Michael Benedikt, Stanislav Kikot, Piotr Ostropolski-Nalewaja, and Miguel Romero. On monotonic determinacy and rewritability for recursive queries and views. ACM Transactions on Computational Logic, 24(2):16:1– 16:??, April 2023. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3572836.

Berkholz:2015:BQD

[BKV15] Christoph Berkholz, Andreas Krebs, and Oleg Verbitsky. Bounds for the quantifier depth in finite-variable logics: Alternation hierarchy. ACM Transactions on Computational Logic, 16(3):21:1–21:??, July 2015. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Basin:2020:RVO

[BKZ20] David Basin, Felix Klaedtke, and Eugen Zalinescu. Runtime verification over out-of-order streams. ACM Transactions on Computational Logic, 21(1):5:1–5:43, January 2020. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Benedikt:2016:LUO

[BL16] Michael Benedikt and Clemens Ley. Limiting until in ordered tree query languages. ACM Transactions on Computational Logic, 17(2):14:1– 14:??, March 2016. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Blondin:2023:COC

[BLM^{+}23] Michael Blondin, Tim Leys, Filip Mazowiecki, Philip Offtermatt, and Guillermo Pérez. Continuous onecounter automata. ACM Transactions on Computational Logic, 24(1):3:1–3:??, January 2023. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https://

dl.acm.org/doi/10.1145/
3558549.

Benedikt:2007:LDQ

[BLN07] Michael Benedikt, Leonid Libkin, and Frank Neven. Logical definability and query languages over ranked and unranked trees. ACM Transactions on Computational Logic, 8(2):??, April 2007. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Borralleras:2019:IST

[BLRC^{+}19] Cristina Borralleras, Daniel Larraz, Enric Rodríguez-Carbonell, Albert Oliveras, and Albert Rubio. Incomplete SMT techniques for solving non-linear formulas over the integers. ACM Transactions on Computational Logic, 20(4):25:1–25:??, September 2019. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3340923.

Bova:2008:PSH

[BM08] Simone Bova and Franco Montagna. Proof search in Hájek’s basic logic. ACM Transactions on Computational Logic, 9(3):21:1–21:??, June 2008. CODEN AT-CLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Beyersdorff:2010:TKL

[BM10] Olaf Beyersdorff and Sebastian Müller. A tight Karp–Lipton collapse result in bounded arithmetic. ACM Transactions on Computational Logic, 11(4):22:1–22:??, July 2010. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Barcelo:2017:GLR

[BM17] Pablo Barceló and Pablo Muñoz. Graph logics with rational relations: The role of word combinatorics. ACM Transactions on Computational Logic, 18(2):10:1–10:??, June 2017. CODEN AT-CLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bianco:2012:GCT

[BMM12] Alessandro Bianco, Fabio Mogavero, and Aniello Murano. Graded computation tree logic. ACM Transactions on Computational Logic, 13 (3):25:1–25:??, August 2012. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Benerecetti:2015:RAS

[BMM15] Massimo Benerecetti, Fabio Mogavero, and Aniello Murano. Reasoning about substructures and games. ACM Transactions on Computational Logic, 16(3):25:1–25:??, July 2015. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bozzelli:2019:IVP

[BMM^{+}19] Laura Bozzelli, Alberto Molinari, Angelo Montanari, Adriano Peron, and Pietro Sala. Interval vs. point temporal logic model checking: an expressiveness comparison. ACM Transactions on Computational Logic, 20 (1):4:1–4:??, February 2019. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3281028.

Berthon:2021:SLI

[BMM^{+}21] Raphaël Berthon, Bastien Maubert, Aniello Murano, Sasha Rubin, and Moshe Y. Vardi. Strategy logic with imperfect information. ACM Transactions on Computational Logic, 22(1):5:1–5:51, January 2021. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3427955.

Bozzelli:2023:ITL

[BMP23] Laura Bozzelli, Angelo Montanari, and Adriano Peron. Interval temporal logic for visibly pushdown systems. ACM Transactions on Computational Logic, 24(3):24:1– 24:??, July 2023. CODEN ???? ISSN 1529-3785 (print), 1557-945X (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3583756.

Bistarelli:2006:SCC

[BMR06] Stefano Bistarelli, Ugo Montanari, and Francesca Rossi. Soft concurrent constraint programming. ACM Transactions on Computational Logic, 7(3):563–589, July 2006. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Baillot:2009:GES

[BMR09] Patrick Baillot, Jean-Yves Marion, and Simona Ronchi Della Rocca. Guest editorial: Special issue on implicit computational complexity. ACM Transactions on Computational Logic, 10(4):23:1–23:??, August 2009. CODEN AT-CLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bistarelli:2010:UMQ

[BMRS10] Stefano Bistarelli, Ugo Montanari, Francesca Rossi, and Francesco Santini. Unicast and multicast QoS routing with soft-constraint logic programming. ACM Transactions on Computational Logic, 12(1):5:1–5:??, October 2010. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bauland:2011:TMC

[BMS^{+}11] Michael Bauland, Martin Mundhenk, Thomas Schneider, Henning Schnoor, Ilka Schnoor, and Heribert Vollmer. The tractability of model checking for LTL: The good, the bad, and the ugly fragments. ACM Transactions on Computational Logic, 12 (2):13:1–13:??, January 2011. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bonsangue:2013:SCA

[BMS13] Marcello M. Bonsangue, Stefan Milius, and Alexandra Silva. Sound and complete axiomatizations of coalgebraic language equivalence. ACM Transactions on Computational Logic, 14(1):7:1– 7:??, February 2013. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bodirsky:2013:CST

[BMT13] Manuel Bodirsky, H. Dugald Macpherson, and Johan Thapper. Constraint satisfaction tractability from semilattice operations on infinite sets. ACM Transactions on Computational Logic, 14(4): 30:1–30:??, November 2013. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bodirsky:2022:PLV

[BMV22] Manuel Bodirsky, Marcello Mamino, and Caterina Viola. Piecewise linear valued CSPs solvable by linear programming relaxation. ACM Transactions on Computational Logic, 23(1):7:1– 7:35, January 2022. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3488721.

Bonatti:2002:SCP

[BO02] Piero Andrea Bonatti and Nicola Olivetti. Sequent calculi for propositional nonmonotonic logics. ACM Transactions on Computational Logic, 3(2):226–278, April 2002. CODEN AT-CLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bouhoula:2009:SCC

[Bou09] Adel Bouhoula. Simultaneous checking of completeness and ground confluence for algebraic specifications. ACM Transactions on Computational Logic, 10(3):20:1– 20:??, April 2009. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bergstra:2011:PA

[BP11] Jan A. Bergstra and Alban Ponse. Proposition algebra. ACM Transactions on Computational Logic, 12(3): 21:1–21:??, May 2011. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Baldan:2020:MCL

[BP20] Paolo Baldan and Tommaso Padoan. Model checking a logic for true concurrency. ACM Transactions on Computational Logic, 21 (4):34:1–34:49, October 2020. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3412853.

Beckmann:2014:PGP

[BPT14] Arnold Beckmann, Pavel Pudlák, and Neil Thapen. Parity games and propositional proofs. ACM Transactions on Computational Logic, 15(2):17:1–17:??, April 2014. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Benedikt:2009:RTL

[BS09] Michael Benedikt and Luc Segoufin. Regular tree languages definable in FO and in FO_{mod}. ACM Transactions on Computational Logic, 11(1):4:1–4:??, October 2009. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Ben-Sasson:2010:LBB

[BSH10] Eli Ben-Sasson and Prahladh Harsha. Lower bounds for bounded depth Frege proofs via Pudlák–Buss games. ACM Transactions on Computational Logic, 11(3):19:1–19:??, May 2010. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bergstra:2023:EER

[BT23] Jan A. Bergstra and John V. Tucker. Eager equality for rational number arithmetic. ACM Transactions on Computational Logic, 24(3):22:1– 22:??, July 2023. CODEN ???? ISSN 1529-3785 (print), 1557-945X (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3580365.

Benedikt:2016:EIP

[BTV16] Michael Benedikt, Balder Ten Cate, and Michael Vanden Boom. Effective interpolation and preservation in guarded logics. ACM Transactions on Computational Logic, 17(2): 8:1–8:??, March 2016. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bulatov:2011:CCC

[Bul11] Andrei A. Bulatov. Complexity of conservative constraint satisfaction problems. ACM Transactions on Computational Logic, 12(4):24:1– 24:??, July 2011. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Burel:2020:LFR

[Bur20] Guillaume Burel. Linking focusing and resolution with selection. ACM Transactions on Computational Logic, 21(3): 18:1–18:30, May 2020. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3373276.

Bryant:2002:BST

[BV02] Randal E. Bryant and Miroslav N.

Velev. Boolean satisfiability with transitivity constraints. ACM Transactions on Computational Logic, 3(4):604– 627, October 2002. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Bogaerts:2016:WFS

[BVD16] Bart Bogaerts, Joost Vennekens, and Marc Denecker. On well-founded set-inductions and locally monotone operators. ACM Transactions on Computational Logic, 17(4): 27:1–27:??, November 2016. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Balbiani:2022:AA

[BvDG22] Philippe Balbiani, Hans van Ditmarsch, and Saúl Fernández González. Asynchronous announcements. ACM Transactions on Computational Logic, 23(2):10:1–10:38, April 2022. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3481806.

Binnewies:2018:SPB

[BZWS18] Sebastian Binnewies, Zhiqiang Zhuang, Kewen Wang, and Bela Stantic. Syntaxpreserving belief change operators for logic programs. ACM Transactions on Computational Logic, 19(2):12:1– 12:??, June 2018. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Chalopin:2019:SPN

[CC19] Jérémie Chalopin and Victor Chepoi. 1-safe Petri nets and special cube complexes: Equivalence and applications. ACM Transactions on Computational Logic, 20(3):17:1–17:??, July 2019. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3322095.

Chadha:2016:AVE

[CCCK16] Rohit Chadha, Vincent Cheval, Stefan Ciobâca, and Steve Kremer. Automated verification of equivalence properties of cryptographic protocols. ACM Transactions on Computational Logic, 17(4):23:1– 23:??, November 2016. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Chretien:2015:SPP

[CCD15] Rémy Chrétien, Véronique Cortier, and Stéphanie Delaune. From security protocols to pushdown automata. ACM Transactions on Computational Logic, 17(1):3:1– 3:??, December 2015. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Chretien:2020:TMF

[CCDD20] Rémy Chrétien, Véronique Cortier, Antoine Dallon, and Stéphanie Delaune. Typing messages for free in security protocols. ACM Transactions on Computational Logic, 21 (1):1:1–1:52, January 2020. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3343507.

Chatterjee:2014:POS

[CD14] Krishnendu Chatterjee and Laurent Doyen. Partialobservation stochastic games: How to win when belief fails. ACM Transactions on Computational Logic, 15(2):16:1– 16:??, April 2014. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Cerrito:2015:OTM

[CDG15] Serenella Cerrito, Amélie David, and Valentin Goranko. Optimal tableau method for constructive satisfiability testing and model synthesis in the alternating-time temporal logic ATL+. ACM Transactions on Computational Logic, 17(1):4:1–4:??, December 2015. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Chatterjee:2010:QL

[CDH10] Krishnendu Chatterjee, Laurent Doyen, and Thomas A. Henzinger. Quantitative languages. ACM Transactions on Computational Logic, 11(4): 23:1–23:??, July 2010. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Chatterjee:2011:QCP

[CDH11] Krishnendu Chatterjee, Luca De Alfaro, and Thomas A. Henzinger. Qualitative concurrent parity games. ACM Transactions on Computational Logic, 12(4):28:1–28:??, July 2011. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Chen:2013:VLD

[CDKM13] Taolue Chen, Marco Diciolla, Marta Kwiatkowska, and Alexandru Mereacre. Verification of linear duration properties over continuoustime Markov chains. ACM Transactions on Computational Logic, 14(4):33:1–33:??, November 2013. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Calvanese:2008:CQC

[CDL08] Diego Calvanese, Giuseppe De Giacomo, and Maurizio Lenzerini. Conjunctive query containment and answering under description logic constraints. ACM Transactions on Computational Logic, 9(3): 22:1–22:??, June 2008. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Cadoli:2005:CPN

[CEG05] Marco Cadoli, Thomas Eiter, and Georg Gottlob. Complexity of propositional nested circumscription and nested abnormality theories. ACM Transactions on Computational Logic, 6(2):232–272, April 2005. CODEN AT-CLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Creignou:2014:CCL

[CES14] Nadia Creignou, Uwe Egly, and Johannes Schmidt. Complexity classifications for logic-based argumentation. ACM Transactions on Computational Logic, 15(3):19:1– 19:??, September 2014. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Cabrer:2019:CAU

[CFMP19] Leonardo M. Cabrer, Benjamin Freisberg, George Metcalfe, and Hilary A. Priestley. Checking admissibility using natural dualities. ACM Transactions on Computational Logic, 20(1):2:1– 2:??, February 2019. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3275115.

Conradie:2019:PEU

[CFP^{+}19] Willem Conradie, Sabine Frittella, Alessandra Palmigiano, Apostolos Tzimoulis, and Nachoem Wijnberg. Probabilistic epistemic updates on algebras. ACM Transactions on Computational Logic, 20(4): 24:1–24:??, September 2019. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3341725.

Cosmadakis:2010:UIR

[CFS10] Stavros Cosmadakis, Eugenie Foustoucos, and Anastasios Sidiropoulos. Undecidability and intractability results concerning datalog programs and their persistency numbers. ACM Transactions on Computational Logic, 11 (2):12:1–12:??, January 2010. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Carreiro:2020:PW

[CFVZ20] Facundo Carreiro, Alessandro Facchini, Yde Venema, and Fabio Zanasi. The power of the weak. ACM Transactions on Computational Logic, 21 (2):15:1–15:47, March 2020. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3372392.

Ciabattoni:2018:HSR

[CG18] Agata Ciabattoni and Francesco A.

Genco. Hypersequents and systems of rules: Embeddings and applications. ACM Transactions on Computational Logic, 19(2):11:1–11:??, June 2018. CODEN AT-CLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Cimatti:2018:ILS

[CGI^{+}18] Alessandro Cimatti, Alberto Griggio, Ahmed Irfan, Marco Roveri, and Roberto Sebastiani. Incremental linearization for satisfiability and verification modulo nonlinear arithmetic and transcendental functions. ACM Transactions on Computational Logic, 19(3):19:1–19:??, September 2018. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Carlucci:2016:PCP

[CGL16] Lorenzo Carlucci, Nicola Galesi, and Massimo Lauria. On the proof complexity of Paris–Harrington and offdiagonal Ramsey tautologies. ACM Transactions on Computational Logic, 17(4):26:1– 26:??, November 2016. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Chen:2022:SCP

[CGPT22] Jinsheng Chen, Giuseppe Greco, Alessandra Palmigiano, and Apostolos Tzimoulis. Syntactic completeness of proper display calculi. ACM Transactions on Computational Logic, 23 (4):23:1–23:??, October 2022. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3529255.

Cimatti:2010:EGC

[CGS10] Alessandro Cimatti, Alberto Griggio, and Roberto Sebastiani. Efficient generation of Craig interpolants in satisfiability modulo theories. ACM Transactions on Computational Logic, 12(1):7:1– 7:??, October 2010. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Calautti:2017:DDC

[CGT17] Marco Calautti, Sergio Greco, and Irina Trubitsyna. Detecting decidable classes of finitely ground logic programs with function symbols. ACM Transactions on Computational Logic, 18(4):28:1– 28:??, December 2017. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Cranen:2015:AFL

[CGWW15] Sjoerd Cranen, Maciej Gazda, Wieger Wesselink, and Tim A. C. Willemse. Abstraction in fixpoint logic. ACM Transactions on Computational Logic, 16(4):29:1–29:??, November 2015. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Charlesworth:2006:CSC

[Cha06] Arthur Charlesworth. Comprehending software correctness implies comprehending an intelligence-related limitation. ACM Transactions on Computational Logic, 7(3): 590–612, July 2006. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Chen:2014:CEP

[Che14] Hubie Chen. On the complexity of existential positive queries. ACM Transactions on Computational Logic, 15 (1):9:1–9:??, February 2014. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Chatterjee:2009:FWR

[CHH09] Krishnendu Chatterjee, Thomas A. Henzinger, and Florian Horn. Finitary winning in ω-regular games. ACM Transactions on Computational Logic, 11(1): 1:1–1:??, October 2009. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Chockler:2008:WCS

[CHK08] Hana Chockler, Joseph Y. Halpern, and Orna Kupferman. What causes a system to satisfy a specification? ACM Transactions on Computational Logic, 9(3): 20:1–20:??, June 2008. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). See erratum [CHK10].

Chockler:2010:EWC

[CHK10] Hana Chockler, Joseph Y. Halpern, and Orna Kupferman. Erratum for “What causes a system to satisfy a specification?”. ACM Transactions on Computational Logic, 11(4):29:1–29:??, July 2010. CODEN ATCLA8.

ISSN 1529-3785 (print), 1557-945X (electronic). See [CHK08].

Cavalcanti:2020:IOC

[CHN20] Ana Cavalcanti, Robert M.

Hierons, and Sidney Nogueira. Inputs and outputs in CSP: a model and a testing theory. ACM Transactions on Computational Logic, 21(3): 24:1–24:53, May 2020. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3379508.

Chatterjee:2017:NWA

[CHO17] Krishnendu Chatterjee, Thomas A.

Henzinger, and Jan Otop. Nested weighted automata. ACM Transactions on Computational Logic, 18(4):31:1– 31:??, December 2017. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Charalambidis:2013:EHO

[CHRW13] Angelos Charalambidis, Konstantinos Handjopoulos, Panagiotis Rondogiannis, and William W. Wadge. Extensional higher-order logic programming. ACM Transactions on Computational Logic, 14(3):21:1–21:??, August 2013. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Carayol:2014:RAI

[CHS14] Arnaud Carayol, Axel Haddad, and Olivier Serre. Randomization in automata on infinite trees. ACM Transactions on Computational Logic, 15(3):24:1–24:??, September 2014. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Carayol:2015:ERA

[CHS15] Arnaud Carayol, Axel Haddad, and Olivier Serre. Erratum for “Randomization in Automata on Infinite Trees”. ACM Transactions on Computational Logic, 16(4):36:1– 36:??, November 2015. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Chen:2013:CLM

[CJL13] Xiaoping Chen, Jianmin Ji, and Fangzhen Lin. Computing loops with at most one external support rule. ACM Transactions on Computational Logic, 14(1):3:1– 3:??, February 2013. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Cohen:2000:NCP

[CK00] Ernie Cohen and Dexter Kozen. A note on the complexity of propositional Hoare logic. ACM Transactions on Computational Logic, 1(1): 171–174, July 2000. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/tocl/2000-1-1/
p171-cohen/.

Cerna:2020:IAU

[CK20] David Cerna and Temur Kutsia. Idempotent antiunification. ACM Transactions on Computational Logic, 21(2):10:1–10:32, March 2020. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3359060.

Chlebowski:2017:AGE

[CKK17] Szymon Chlebowski, Maciej Komosinski, and Adam Kups. Automated generation of erotetic search scenarios: Classification, optimization, and knowledge extraction. ACM Transactions on Computational Logic, 18(2): 8:1–8:??, June 2017. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Chatzikokolakis:2012:ESG

[CKPP12]  Konstantinos  Chatzikoko-

lakis, Sophia Knight, Catuscia Palamidessi, and Prakash Panangaden. Epistemic strategies and games on concurrent processes. ACM Transactions on Computational Logic, 13(4):28:1–28:??, October 2012. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Chevalier:2008:CRS

[CKRT08] Yannick Chevalier, Ralf Küsters, Michaël Rusinowitch, and Mathieu Turuani. Complexity results for security protocols with Diffie–Hellman exponentiation and commuting public key encryption. ACM Transactions on Computational Logic, 9(4):24:1– 24:??, August 2008. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Comon-Lundh:2010:DSP

[CLCZ10] Hubert Comon-Lundh, Véronique Cortier, and Eugen Zãlinescu. Deciding security properties for cryptographic protocols. application to key cycles. ACM Transactions on Computational Logic, 11(2):9:1– 9:??, January 2010. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Clemente:2019:BRT

[CLLM19] Lorenzo Clemente, Sławomir Lasota, Ranko Lazić, and Filip Mazowiecki. Binary reachability of timed-register pushdown automata and branching vector addition systems. ACM Transactions on Computational Logic, 20 (3):14:1–14:??, July 2019. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3326161.

Ciabattoni:2021:DLP

[CLRT21] Agata Ciabattoni, Tim S.

Lyon, Revantha Ramanayake, and Alwen Tiu. Display to labeled proofs and back again for tense logics. ACM Transactions on Computational Logic, 22(3):20:1–20:31, June 2021. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

https://dl.acm.org/doi/
10.1145/3460492.

Choi:2007:RPR

[CLS07] C. W. Choi, J. H. M. Lee, and P. J. Stuckey. Removing propagation redundant constraints in redundant modeling. ACM Transactions on Computational Logic, 8(4): 23:1–23:??, August 2007. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Ciabattoni:2015:TPO

[CLSZ15] Agata Ciabattoni, Ori Lahav, Lara Spendier, and Anna Zamansky. Taming paraconsistent (and other) logics: an algorithmic approach. ACM Transactions on Computational Logic, 16(1):5:1– 5:??, March 2015. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Cai:2023:LSS

[CLZ23] Shaowei Cai, Bohan Li, and Xindi Zhang. Local search for satisfiability modulo integer arithmetic theories. ACM Transactions on Computational Logic, 24(4):32:1–32:??, October 2023. CODEN ???? ISSN 1529-3785 (print), 1557-945X (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3597495.

Coppola:2006:OOR

[CM06] Paolo Coppola and Simone Martini. Optimizing optimal reduction: a type inference algorithm for elementary affine logic. ACM Transactions on Computational Logic, 7(2):219–260, April 2006. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Chen:2017:OHS

[CM17] Hubie Chen and Moritz Müller. One hierarchy spawns another: Graph deconstructions and the complexity classification of conjunctive queries. ACM Transactions on Computational Logic, 18(4):29:1–29:??, December 2017. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Carvalho:2023:CQC

[CMMZ23] Catarina Carvalho, Florent Madelaine, Barnaby Martin, and Dmitriy Zhuk.

The complexity of quantified constraints: Collapsibility, switchability, and the algebraic formulation. ACM Transactions on Computational Logic, 24(1):5:1–5:??, January 2023. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https://

dl.acm.org/doi/10.1145/
3568397.

Creignou:2012:CRF

[CMVT12] Nadia Creignou, Arne Meier, Heribert Vollmer, and Michael Thomas. The complexity of reasoning for fragments of autoepistemic logic. ACM Transactions on Computational Logic, 13(2):17:1– 17:??, April 2012. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Comon:2003:DCO

[CNNR03] Hubert Comon, Paliath Narendran, Robert Nieuwenhuis, and Michael Rusinowitch. Deciding the confluence of ordered term rewrite systems. ACM Transactions on Computational Logic, 4(1): 33–55, January 2003. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Case:2022:IAR

[CP22] Adam Case and Christopher P. Porter. The intersection of algorithmically random closed sets and effective dimension. ACM Transactions on Computational Logic, 23(4):24:1–24:??, October 2022. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3545114.

Creignou:2016:BMW

[CPRW16] Nadia Creignou, Odile Papini, Stefan Rümmele, and Stefan Woltran. Belief merging within fragments of propositional logic. ACM Transactions on Computational Logic, 17(3):20:1–20:??, July 2016. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Crolard:2009:ELL

[CPV09] Tristan Crolard, Emmanuel Polonowski, and Pierre Valarcher. Extending the loop language with higher-order procedural variables. ACM Transactions on Computational Logic, 10 (4):26:1–26:??, August 2009. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Crafa:2015:LCB

[CR15] Silvia Crafa and Francesco Ranzato. Logical characterizations of behavioral relations on transition systems of probability distributions. ACM Transactions on Computational Logic, 16(1):2:1– 2:??, March 2015. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Ciabattoni:2016:PLS

[CR16] Agata Ciabattoni and Revantha Ramanayake. Power and limits of structural display rules. ACM Transactions on Computational Logic, 17(3): 17:1–17:??, July 2016. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Cohen:2020:NWF

[CR20] Liron Cohen and Reuben N. S. Rowe. Non-well-founded proof theory of transitive closure logic. ACM Transactions on Computational Logic, 21 (4):31:1–31:31, October 2020. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3404889.

Crary:2007:SCE

[Cra07] Karl Crary. Sound and complete elimination of singleton kinds. ACM Transactions on Computational Logic, 8(2):??, April 2007. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Crary:2008:FCC

[CS08] Karl Crary and Susmit Sarkar. Foundational certified code in the Twelf metalogical framework. ACM Transactions on Computational Logic, 9(3):16:1–16:??, June 2008. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Carayol:2020:HGS

[CS20] Arnaud Carayol and Olivier Serre. How good is a strategy in a game with nature? ACM Transactions on Computational Logic, 21(3): 21:1–21:39, May 2020. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3377137.

Cohen:2005:EAA

[CSN05] Sara Cohen, Yehoshua Sagiv, and Werner Nutt. Equivalences among aggregate queries with negation. ACM Transactions on Computational Logic, 6(2):328–360, April 2005. CODEN AT-CLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Creignou:2010:NBC

[CSS10] Nadia Creignou, Henning Schnoor, and Ilka Schnoor. Nonuniform Boolean constraint satisfaction problems with cardinality constraint. ACM Transactions on Computational Logic, 11(4):24:1– 24:??, July 2010. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Cook:2006:SRW

[CT06] Stephen Cook and Neil Thapen. The strength of replacement in weak arithmetic. ACM Transactions on Computational Logic, 7(4): 749–764, October 2006. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Chadha:2010:CGA

[CV10] Rohit Chadha and Mahesh Viswanathan. A counterexample-guided abstractionrefinement framework for Markov decision processes. ACM Transactions on Computational Logic, 12(1):1:1– 1:??, October 2010. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Charatonik:2016:TVL

[CW16] Witold Charatonik and Piotr Witkowski. Two-variable logic with counting and trees. ACM Transactions on Computational Logic, 17(4):31:1– 31:??, November 2016. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Dagnino:2022:MTB

[Dag22] Francesco Dagnino. A metatheory for big-step semantics. ACM Transactions on Computational Logic, 23 (3):20:1–20:50, July 2022. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3522729.

DalLago:2009:CSL

[Dal09a] Ugo Dal Lago. Context semantics, linear logic, and computational complexity. ACM Transactions on Computational Logic, 10(4):25:1– 25:??, August 2009. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

DalLago:2009:GLH

[Dal09b] Ugo Dal Lago. The geometry of linear higher-order recursion. ACM Transactions on Computational Logic, 10(2): 8:1–8:??, February 2009. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Denecker:2001:LPR

[DBM01] Marc Denecker, Maurice Bruynooghe, and Victor Marek. Logic programming revisited: Logic programs as inductive definitions. ACM Transactions on Computational Logic, 2(4):623– 654, October 2001. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Dezani-Ciancaglini:2010:IIT

[DCDGT10] Mariangiola Dezani-Ciancaglini, Roberto Di Cosmo, Elio Giovannetti, and Makoto Tatsuta. On isomorphisms of intersection types. ACM Transactions on Computational Logic, 11(4):25:1–25:??, July 2010. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Dezani-Ciancaglini:2003:CCC

[DCHA03] M. Dezani-Ciancaglini, F. Honsell, and F. Alessi. A complete characterization of complete intersection-type preorders. ACM Transactions on Computational Logic, 4(1): 120–147, January 2003. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Demri:2015:TVS

[DD15] Stéphane Demri and Morgan Deters. Two-variable separation logic and its inner circle. ACM Transactions on Computational Logic, 16(2): 15:1–15:??, March 2015. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Demri:2016:ECS

[DD16] Stephane Demri and Morgan Deters. Expressive completeness of separation logic with two variables and no separating conjunction. ACM Transactions on Computational Logic, 17(2):12:1– 12:??, March 2016. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Doherty:2022:UOW

[DDDW22] Simon Doherty, Sadegh Dalvandi, Brijesh Dongol, and Heike Wehrheim. Unifying operational weak memory verification: an axiomatic approach. ACM Transactions on Computational Logic, 23 (4):27:1–27:??, October 2022. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3545117.

DeBruijn:2011:ENL

[DEPT11] Jos De Bruijn, Thomas Eiter, Axel Polleres, and Hans Tompits. Embedding nonground logic programs into autoepistemic logic for knowledge-base combination. ACM Transactions on Computational Logic, 12(3):20:1–20:??, May 2011. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Degtyarev:2006:MTR

[DFK06] Anatoli Degtyarev, Michael Fisher, and Boris Konev. Monodic temporal resolution. ACM Transactions on Computational Logic, 7(1):108– 150, January 2006. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Dovier:2006:DRS

[DFO06] Agostino Dovier, Andrea Formisano, and Eugenio G.

Omodeo. Decidability results for sets with atoms. ACM Transactions on Computational Logic, 7(2):269– 301, April 2006. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Durand:2007:FOQ

[DG07] Arnaud Durand and Etienne Grandjean. Firstorder queries on structures of bounded degree are computable with constant delay. ACM Transactions on Computational Logic, 8(4):21:1– 21:??, August 2007. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Dowek:2012:PNL

[DG12] Gilles Dowek and Murdoch J.

Gabbay. Permissive-nominal logic: First-order logic over nominal terms and sets. ACM Transactions on Computational Logic, 13(3):20:1– 20:??, August 2012. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Dawar:2004:IFP

[DGK04] Anuj Dawar, Erich Grädel, and Stephan Kreutzer. Inflationary fixed points in modal logic. ACM Transactions on Computational Logic, 5(2): 282–315, April 2004. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

DeBoer:2004:PCT

[DGM04] Frank S. De Boer, Maurizio Gabbrielli, and Maria Chiara Meo. Proving correctness of timed concurrent constraint programs. ACM Transactions on Computational Logic, 5(4): 706–731, October 2004. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

DiGiusto:2012:EPM

[DGM12] Cinzia Di Giusto, Maurizio Gabbrielli, and Maria Chiara Meo. On the expressive power of multiple heads in CHR. ACM Transactions on Computational Logic, 13(1):6:1– 6:??, January 2012. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Danos:2002:PGS

[DH02] Vincent Danos and Russell S. Harmer. Probabilistic game semantics. ACM Transactions on Computational Logic, 3(3):359–382, July 2002. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Daca:2017:FSM

[DHKP17] Przemyslaw Daca, Thomas A. Henzinger, Jan Kretínský, and Tatjana Petrov. Faster statistical model checking for unbounded temporal properties. ACM Transactions on Computational Logic, 18(2): 12:1–12:??, June 2017. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Danicic:2012:CDD

[DHL12] Sebastian Danicic, Robert M.

Hierons, and Michael R. Laurence. Complexity of data dependence problems for program schemas with concurrency. ACM Transactions on Computational Logic, 13(2): 15:1–15:??, April 2012. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Dongol:2016:CUC

[DHS16] Brijesh Dongol, Ian J. Hayes, and Georg Struth. Convolution as a unifying concept: Applications in separation logic, interval calculi, and concurrency. ACM Transactions on Computational Logic, 17(3):15:1–15:??, July 2016. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Durand:2012:HDL

[DK12] Arnaud Durand and Juha Kontinen. Hierarchies in dependence logic. ACM Transactions on Computational Logic, 13(4):31:1–31:??, October 2012. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Durand:2022:TFD

[DKDV22] Arnaud Durand, Juha Kontinen, Nicolas De Rugy-Altherre, and Jouko Väänänen. Tractability frontier of data complexity in team semantics. ACM Transactions on Computational Logic, 23 (1):3:1–3:21, January 2022. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3471618.

Dix:2006:HTP

[DKS06] Jürgen Dix, Sarit Kraus, and V. S. Subrahmanian. Heterogeneous temporal probabilistic agents. ACM Transactions on Computational Logic, 7(1): 151–198, January 2006. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

DeHaan:2017:PCF

[DKS17] Ronald De Haan, Iyad Kanj, and Stefan Szeider. On the parameterized complexity of finding small unsatisfiable subsets of CNF formulas and CSP instances. ACM Transactions on Computational Logic, 18(3):21:1– 21:??, August 2017. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

DeNicola:2004:MLM

[DL04] Rocco De Nicola and Michele Loreti. A modal logic for mobile agents. ACM Transactions on Computational Logic, 5(1):79–128, January 2004. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Demri:2009:LFQ

[DL09] Stéphane Demri and Ranko Lazić. LTL with the freeze quantifier and register automata. ACM Transactions on Computational Logic, 10 (3):16:1–16:??, April 2009. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Demri:2021:EAR

[DLM21] Stéphane Demri, Etienne Lozes, and Alessio Mansutti. The effects of adding reachability predicates in quantifierfree separation logic. ACM Transactions on Computational Logic, 22(2):14:1–14:56, June 2021. CODEN AT-CLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https://

dl.acm.org/doi/10.1145/
3448269.

DeGiacomo:2001:IEG

[DLS01] Giuseppe De Giacomo, Hector J. Levesque, and Sebastian Sardiña. Incremental execution of guarded theories. ACM Transactions on Computational Logic, 2(4): 495–525, October 2001. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Dapic:2017:QCS

[DMM17] Petar Dapić, Petar Marković, and Barnaby Martin. Quantified constraint satisfaction problem on semicomplete digraphs. ACM Transactions on Computational Logic, 18 (1):2:1–2:??, April 2017. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Dima:2018:RPT

[DMP18] Catalin Dima, Bastien Maubert, and Sophie Pinchinat. Relating paths in transition systems: The fall of the modal Mu-calculus. ACM Transactions on Computational Logic, 19(3):23:1–23:??, September 2018. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Desharnais:2006:KAD

[DMS06] Jules Desharnais, Bernhard Möller, and Georg Struth. Kleene algebra with domain. ACM Transactions on Computational Logic, 7(4):798– 833, October 2006. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Deuser:2021:SKA

[DN21] Kaya Deuser and Pavel Naumov. Strategic knowledge acquisition. ACM Transactions on Computational Logic, 22(3):17:1–17:18, June 2021. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3459993.

Donini:2002:DLM

[DNR02] Francesco M. Donini, Daniele Nardi, and Riccardo Rosati. Description logics of minimal knowledge and negation as failure. ACM Transactions on Computational Logic, 3(2): 177–225, April 2002. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Dix:2000:PAP

[DNS00] Jürgen Dix, Mirco Nanni, and V. S. Subrahmanian. Probabilistic agent programs. ACM Transactions on Computational Logic, 1(2):208–246, 2000. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

http://www.acm.org/pubs/
citations/journals/tocl/
2000-1-2/p208-dix/.

Dovier:2008:UAC

[DPR08] Agostino Dovier, Carla Piazza, and Gianfranco Rossi. A uniform approach to constraint-solving for lists, multisets, compact lists, and sets. ACM Transactions on Computational Logic, 9(3): 15:1–15:??, June 2008. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Drabent:2016:CCL

[Dra16] Wlodzimierz Drabent. Correctness and completeness of logic programs. ACM Transactions on Computational Logic, 17(3):18:1–18:??, July 2016. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Dyckhoff:2013:APT

[DST13] Roy Dyckhoff, Mehrnoosh Sadrzadeh, and Julien Truffaut. Algebra, proof theory and applications for an intuitionistic logic of propositions, actions and adjoint modal operators. ACM Transactions on Computational Logic, 14(4):34:1–34:??, November 2013. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Delgrande:2013:MTA

[DSTW13] James Delgrande, Torsten Schaub, Hans Tompits, and Stefan Woltran. A modeltheoretic approach to belief change in answer set programming. ACM Transactions on Computational Logic, 14(2): 14:1–14:??, June 2013. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Denecker:2008:LNI

[DT08] Marc Denecker and Eugenia Ternovska. A logic of nonmonotone inductive definitions. ACM Transactions on Computational Logic, 9(2): 14:1–14:??, March 2008. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Dershowitz:2010:CPP

[DT10] Nachum Dershowitz and Iddo Tzameret. Complexity of propositional proofs under a promise. ACM Transactions on Computational Logic, 11(3):18:1–18:??, May 2010. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Dudenhefner:2021:KSI

[DU21] Andrej Dudenhefner and Paweł Urzyczyn. Kripke semantics for intersection formulas. ACM Transactions on Computational Logic, 22 (3):15:1–15:16, June 2021. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3453481.

Dawar:2022:SCR

[DW22] Anuj Dawar and Gregory Wilsenach. Symmetric circuits for rank logic. ACM Transactions on Computational Logic, 23(1):6:1–6:35, January 2022. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https://

dl.acm.org/doi/10.1145/
3476227.

Danvy:2013:TST

[DZ13] Olivier Danvy and Ian Zerny. Three syntactic theories for combinatory graph reduction. ACM Transactions on Computational Logic, 14(4):29:1– 29:??, November 2013. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Eberhard:2017:ACF

[EEH17a] Sebastian Eberhard, Gabriel Ebner, and Stefan Hetzl. Algorithmic compression of finite tree languages by rigid acyclic grammars. ACM Transactions on Computational Logic, 18(4):26:1–26:??, December 2017. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Eickmeyer:2017:SOI

[EEH17b] Kord Eickmeyer, Michael Elberfeld, and Frederik Harwath. Succinctness of orderinvariant logics on depthbounded structures. ACM Transactions on Computational Logic, 18(4):33:1–33:??, December 2017. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Eisner:2014:SLW

[EFH14] Cindy Eisner, Dana Fisman, and John Havlicek. Safety and liveness, weakness and strength, and the underlying topological relations. ACM Transactions on Computational Logic, 15(2): 13:1–13:??, April 2014. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Eiter:2004:LPA

[EFL^{+}04] Thomas Eiter, Wolfgang Faber, Nicola Leone, Gerald Pfeifer, and Axel Polleres. A logic programming approach to knowledge-state planning: Semantics and complexity. ACM Transactions on Computational Logic, 5(2): 206–263, April 2004. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Eiter:2005:RAE

[EFST05] Thomas Eiter, Michael Fink, Giuliana Sabbatini, and Hans Tompits. Reasoning about evolving nonmonotonic knowledge bases. ACM Transactions on Computational Logic, 6(2):389–440, April 2005. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Eiter:2007:SCC

[EFW07] Thomas Eiter, Michael Fink, and Stefan Woltran. Semantical characterizations and complexity of equivalences in answer set programming. ACM Transactions on Computational Logic, 8(3):17:1–17:??, July 2007. CODEN ATCLA8.

ISSN 1529-3785 (print), 1557-945X (electronic).

Enqvist:2022:TLC

[EG22] Sebastian Enqvist and Valentin Goranko. The temporal logic of coalitional goal assignments in concurrent multiplayer games. ACM Transactions on Computational Logic, 23 (4):21:1–21:??, October 2022. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3517128.

Elberfeld:2016:WFO

[EGT16] Michael Elberfeld, Martin Grohe, and Till Tantau. Where first-order and monadic second-order logic coincide. ACM Transactions on Computational Logic, 17(4):25:1–25:??, November 2016. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Engelfriet:2001:MDS

[EH01] Joost Engelfriet and Hendrik Jan Hoogeboom. MSO definable string transductions and two-way finitestate transducers. ACM Transactions on Computational Logic, 2(2):216–254, 2001. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/tocl/
                                        2001-2-2/p216-engelfriet/

p216-engelfriet.pdf;   http:
//www.acm.org/pubs/citations/
journals/tocl/2001-2-2/
p216-engelfriet/.

Eiter:2011:WFS

[EILS11] Thomas Eiter, Giovambattista Ianni, Thomas Lukasiewicz, and Roman Schindlauer. Well-founded semantics for description logic programs in the Semantic Web. ACM Transactions on Computational Logic, 12(2):11:1–11:??, January 2011. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Echenim:2020:BSR

[EIP20] Mnacho Echenim, Radu Iosif, and Nicolas Peltier. The Bernays–Schönfinkel–Ramsey class of separation logic with uninterpreted predicates. ACM Transactions on Computational Logic, 21(3):19:1– 19:46, May 2020. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3380809.

Echenim:2013:ISN

[EP13] Mnacho Echenim and Nicolas Peltier. Instantiation schemes for nested theories. ACM Transactions on Computational Logic, 14(2):11:1– 11:??, June 2013. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Eiter:2010:FDN

[EŠ10] Thomas Eiter and Mantas Šimkus. FDNC: Decidable nonmonotonic disjunctive logic programs with function symbols. ACM Transactions on Computational Logic, 11(2):14:1–14:??, January 2010. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Egly:2001:PCR

[ET01] Uwe Egly and Hans Tompits. Proof-complexity results for nonmonotonic reasoning. ACM Transactions on Computational Logic, 2(3):340– 387, 2001. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

                                        http://www.acm.org/pubs/
                                        citations/journals/tocl/
                                        2001-2-3/p340-egly/.

East:2006:PCB

[ET06] Deborah East and Miroslaw Truszczyński. Predicatecalculus-based logics for modeling and solving search problems. ACM Transactions on Computational Logic, 7(1): 38–83, January 2006. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Eickmeyer:2020:MCO

[EvdHK^{+}20] Kord Eickmeyer, Jan van den Heuvel, Ken-Ichi Kawarabayashi, Stephan Kreutzer, Patrice Ossona De Mendez, Michał Pilipczuk, Daniel A. Quiroz, Roman Rabinovich, and Sebastian Siebertz. Modelchecking on ordered structures. ACM Transactions on Computational Logic, 21(2): 11:1–11:28, March 2020. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3360011.

Etessami:2012:MCR

[EY12] Kousha Etessami and Mihalis Yannakakis. Model checking of recursive probabilistic systems. ACM Transactions on Computational Logic, 13(2): 12:1–12:??, April 2012. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Feng:2020:MTO

[FCGQ20] Shiguang Feng, Claudia Carapelle, Oliver Fernández Gil, and Karin Quaas. MTL and TPTL for one-counter machines: Expressiveness, model checking, and satisfiability. ACM Transactions on Computational Logic, 21(2):12:1– 12:34, March 2020. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3372789.

Fernandez-Duque:2014:NFA

[FD14] David Fernández-Duque. Nonfinite axiomatizability of dynamic topological logic. ACM Transactions on Computational Logic, 15(1):4:1–4:??, February 2014. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Fisher:2001:CTR

[FDP01] Michael Fisher, Clare Dixon, and Martin Peim. Clausal temporal resolution. ACM Transactions on Computational Logic, 2(1):12–56, 2001. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/tocl/
                                        2001-2-1/p12-fisher/p12-
                                        fisher.pdf;   http://www.
                                        acm.org/pubs/citations/
                                        journals/tocl/2001-2-1/
                                        p12-fisher/.

Feng:2014:SBQ

[FDY14] Yuan Feng, Yuxin Deng, and Mingsheng Ying. Symbolic bisimulation for quantum processes. ACM Transactions on Computational Logic, 15(2): 14:1–14:??, April 2014. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Ferraris:2011:LPP

[Fer11] Paolo Ferraris. Logic programs with propositional connectives and aggregates. ACM Transactions on Computational Logic, 12(4):25:1–25:??, July 2011. CODEN ATCLA8.

ISSN 1529-3785 (print), 1557-945X (electronic).

Fairweather:2018:TNR

[FF18] Elliot Fairweather and Maribel Fernández. Typed nominal rewriting. ACM Transactions on Computational Logic, 19(1):6:1–6:??, February 2018. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Fiorentini:2020:DBU

[FF20] Camillo Fiorentini and Mauro Ferrari. Duality between unprovability and provability in forward refutation-search for intuitionistic propositional logic. ACM Transactions on Computational Logic, 21(3): 22:1–22:47, May 2020. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3372299.

Ferrari:2005:CDP

[FFF05] Mauro Ferrari, Camillo Fiorentini, and Guido Fiorino. On the complexity of the disjunction property in intuitionistic and modal logics. ACM Transactions on Computational Logic, 6(3):519–538, July 2005. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Ferrari:2012:SRI

[FFF12] Mauro Ferrari, Camillo Fiorentini, and Guido Fiorino. Simplification rules for intuitionistic propositional tableaux. ACM Transactions on Computational Logic, 13(2):14:1– 14:??, April 2012. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Ferrari:2015:EDD

[FFF15] Mauro Ferrari, Camillo Fiorentini, and Guido Fiorino. An evaluation-driven decision procedure for G3i. ACM Transactions on Computational Logic, 16(1):8:1–8:??, March 2015. CODEN AT-CLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Fazzinga:2015:CPA

[FFP15] Bettina Fazzinga, Sergio Flesca, and Francesco Parisi. On the complexity of probabilistic abstract argumentation frameworks. ACM Transactions on Computational Logic, 16(3):22:1–22:??, July 2015. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Fomin:2022:PCE

[FGT22] Fedor V. Fomin, Petr A. Golovach, and Dimitrios M. Thilikos. Parameterized complexity of elimination distance to first-order logic properties. ACM Transactions on Computational Logic, 23 (3):17:1–17:35, July 2022. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3517129.

Friedman:2000:FOC

[FHK00] Nir Friedman, Joseph Y.

Halpern, and Daphne Koller. First-order conditional logic for default reasoning revisited. ACM Transactions on Computational Logic, 1(2):175– 207, 2000. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

http://www.acm.org/pubs/
citations/journals/tocl/
2000-1-2/p175-friedman/.

Figueira:2012:DDX

[Fig12] Diego Figueira. Decidability of downward XPath. ACM Transactions on Computational Logic, 13(4):34:1– 34:??, October 2012. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Friedmann:2015:RBI

[FKL15] Oliver Friedmann, Felix Klaedtke, and Martin Lange. Ramsey-based inclusion checking for visibly pushdown automata. ACM Transactions on Computational Logic, 16(4):34:1–34:??, November 2015. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Fuhs:2017:VPP

[FKN17] Carsten Fuhs, Cynthia Kop, and Naoki Nishida. Verifying procedural programs via constrained rewriting induction. ACM Transactions on Computational Logic, 18(2): 14:1–14:??, June 2017. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Fan:2020:RMB

[FL20] Tuan-Fang Fan and Churn-Jung Liau. Reason-maintenance belief logic with uncertain information. ACM Transactions on Computational Logic, 21(1):3:1–3:32, January 2020. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3355608.

Filmus:2015:SSS

[FLM^{+}15] Yuval Filmus, Massimo Lauria, Mladen Miksa, Jakob Nordström, and Marc Vinyals. From small space to small width in resolution. ACM Transactions on Computational Logic, 16(4):28:1–28:??, November 2015. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Feng:2022:VDQ

[FLY22] Yuan Feng, Sanjiang Li, and Mingsheng Ying. Verification of distributed quantum programs. ACM Transactions on Computational Logic, 23(3):19:1–19:40, July 2022. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3517145.

Facchini:2016:IPG

[FMS16] Alessandro Facchini, Filip Murlak, and Michal Skrzypczak. Index problems for game automata. ACM Transactions on Computational Logic, 17(4):24:1–24:??, November 2016. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Filmus:2023:MRS

[FMSV23] Yuval Filmus, Meena Mahajan, Gaurav Sood, and Marc Vinyals. MaxSAT resolution and subcube sums. ACM Transactions on Computational Logic, 24(1):8:1– 8:??, January 2023. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https://

dl.acm.org/doi/10.1145/
3565363.

Felgenhauer:2015:LSP

[FMZV15] Bertram Felgenhauer, Aart Middeldorp, Harald Zankl, and Vincent Van Oostrom. Layer systems for proving confluence. ACM Transactions on Computational Logic, 16(2):14:1–14:??, March 2015. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Fontaine:2015:WIH

[Fon15] Gaëlle Fontaine. Why is it hard to obtain a dichotomy for consistent query answering? ACM Transactions on Computational Logic, 16(1): 7:1–7:??, March 2015. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Fijalkow:2017:MSO

[FP17] Nathanaël Fijalkow and Charles Paperman. Monadic secondorder logic with arbitrary monadic predicates. ACM Transactions on Computational Logic, 18(3):20:1–20:??, August 2017. CODEN AT-CLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Frigeri:2014:FTL

[FPS14] Achille Frigeri, Liliana Pasquale, and Paola Spoletini. Fuzzy time in linear temporal logic. ACM Transactions on Computational Logic, 15(4):30:1– 30:??, September 2014. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Furia:2010:TSC

[FR10] Carlo A. Furia and Matteo Rossi. A theory of sampling for continuous-time metric temporal logic. ACM Transactions on Computational Logic, 12(1):8:1–8:??, October 2010. CODEN ATCLA8.

ISSN 1529-3785 (print), 1557-945X (electronic).

Fichte:2015:BND

[FS15a] Johannes K. Fichte and Stefan Szeider. Backdoors to normality for disjunctive logic programs. ACM Transactions on Computational Logic, 17 (1):7:1–7:??, December 2015. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Furusawa:2015:CDA

[FS15b] Hitoshi Furusawa and Georg Struth. Concurrent dynamic algebra. ACM Transactions on Computational Logic, 16(4):30:1–30:??, November 2015. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Furusawa:2016:TM

[FS16] Hitoshi Furusawa and Georg Struth. Taming multirelations. ACM Transactions on Computational Logic, 17(4): 28:1–28:??, November 2016. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Gurfinkel:2012:RVB

[GC12] Arie Gurfinkel and Marsha Chechik. Robust vacuity for branching temporal logic. ACM Transactions on Computational Logic, 13(1):1:1– 1:??, January 2012. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Gratzer:2022:MPA

[GCK^{+}22] Daniel Gratzer, Evan Cavallo, G. A. Kavvos, Adrien Guatto, and Lars Birkedal. Modalities and parametric adjoints. ACM Transactions on Computational Logic, 23(3):18:1–18:29, July 2022. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3514241.

Gall:2018:OSC

[GF18] Daniel Gall and Thom Frühwirth. An operational semantics for the cognitive architecture ACT-R and its translation to constraint handling rules. ACM Transactions on Computational Logic, 19(3):22:1–22:??, September 2018. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Ghilardi:2023:IRA

[GGKN23] Silvio Ghilardi, Alessandro Gianola, Deepak Kapur, and Chiara Naso. Interpolation results for arrays with length and MaxDiff. ACM Transactions on Computational Logic, 24(4):28:1–28:??, October 2023. CODEN ???? ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3587161.

Ghafari:2012:RPP

[GGKT12] Naghmeh Ghafari, Arie Gurfinkel, Nils Klarlund, and Richard Trefler. Reachability problems in piecewise FIFO systems. ACM Transactions on Computational Logic, 13(1): 7:1–7:??, January 2012. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Giordano:2009:ATC

[GGOP09] Laura Giordano, Valentina Gliozzi, Nicola Olivetti, and Gian Luca Pozzato. Analytic tableaux calculi for KLM logics of nonmonotonic reasoning. ACM Transactions on Computational Logic, 10(3): 18:1–18:??, April 2009. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Giordano:2009:TCP

[GGOS09] Laura Giordano, Valentina Gliozzi, Nicola Olivetti, and Camilla Schwind. Tableau calculus for preference-based conditional logics: PCL and its extensions. ACM Transactions on Computational Logic, 10(3):21:1–21:??, April 2009. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Ghilardi:2020:FPE

[GGS20] Silvio Ghilardi, Maria João Gouveia, and Luigi Santocanale. Fixed-point elimination in the intuitionistic propositional calculus. ACM Transactions on Computational Logic, 21(1):4:1– 4:37, January 2020. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3359669.

Gascon:2011:UMC

[GGSS11] Adrià Gascón, Guillem Godoy, and Manfred Schmidt-Schauss. Unification and matching on compressed terms. ACM Transactions on Computational Logic, 12(4):26:1–26:??, July 2011. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Gottlob:2002:DLD

[GGV02] Georg Gottlob, Erich Grädel, and Helmut Veith. Datalog LITE: a deductive query language with linear time model checking. ACM Transactions on Computational Logic, 3(1): 42–79, January 2002. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Gradel:2003:LSI

[GHJP03] Erich Grädel, Joseph Y. Halpern, Radha Jaghadeesan, and Adolfo Piperno. LICS 2001 special issue. ACM Transactions on Computational Logic, 4(3):295, July 2003. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Geerts:2008:FOC

[GHK08] Floris Geerts, Sofie Haesevoets, and Bart Kuijpers. First-order complete and computationally complete query languages for spatio-temporal databases. ACM Transactions on Computational Logic, 9(2): 13:1–13:??, March 2008. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Gradel:2002:BFB

[GHO02] Erich Grädel, Colin Hirsch, and Martin Otto. Back and forth between guarded and modal logics. ACM Transactions on Computational Logic, 3(3):418–463, July 2002. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Gajarsky:2020:NPF

[GHO^{+}20] Jakub Gajarský, Petr Hlinený, Jan Obdrzálek, Daniel Lokshtanov, and M. S. Ramanujan. A new perspective on FO model checking of dense graph classes. ACM Transactions on Computational Logic, 21(4):28:1– 28:23, October 2020. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https://

dl.acm.org/doi/10.1145/
3383206.

Gutierrez:2021:ENE

[GHPW21] Julian Gutierrez, Paul Harrenstein, Giuseppe Perelli, and Michael Wooldridge. Expressiveness and Nash equilibrium in iterated Boolean games. ACM Transactions on Computational Logic, 22 (2):8:1–8:38, June 2021. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https://

dl.acm.org/doi/10.1145/
3439900.

Goldblatt:2012:WSP

[GJ12] Robert Goldblatt and Marcel Jackson. Well-structured program equivalence is highly undecidable. ACM Transactions on Computational Logic, 13 (3):26:1–26:??, August 2012. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Groote:2017:ACS

[GJKW17] Jan Friso Groote, David N. Jansen, Jeroen J. A. Keiren, and Anton J. Wijs. An O(m log n) algorithm for computing stuttering equivalence and branching bisimulation. ACM Transactions on Computational Logic, 18(2):13:1– 13:??, June 2017. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Goller:2015:CDM

[GJL15] Stefan Göller, Jean-Christoph Jung, and Markus Lohrey. The complexity of decomposing modal and first-order theories. ACM Transactions on Computational Logic, 16(1): 9:1–9:??, March 2015. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Genaim:2008:INS

[GK08] Samir Genaim and Andy King. Inferring non-suspension conditions for logic programs with dynamic scheduling. ACM Transactions on Computational Logic, 9(3):17:1– 17:??, June 2008. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Gnaedig:2009:TRU

[GK09] Isabelle Gnaedig and Hélène Kirchner. Termination of rewriting under strategies. ACM Transactions on Computational Logic, 10(2):10:1– 10:??, February 2009. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Garlik:2018:SSC

[GK18] Michal Garlík and Leszek Aleksander Kołodziejczyk. Some subsystems of constant-depth Frege with parity. ACM Transactions on Computational Logic, 19(4):29:1–29:??, December 2018. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3243126.

Gajarsky:2020:FOI

[GKN^{+}20] Jakub Gajarský, Stephan Kreutzer, Jaroslav NesETril, Patrice Ossona De Mendez, Michał Pilipczuk, Sebastian Siebertz, and Szymon Toruńczyk. First-order interpretations of bounded expansion classes. ACM Transactions on Computational Logic, 21(4):29:1–29:41, October 2020. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

https://dl.acm.org/doi/
10.1145/3382093.

Geck:2019:PCC

[GKNS19] Gaetano Geck, Bas Ketsman, Frank Neven, and Thomas Schwentick. Parallelcorrectness and containment for conjunctive queries with union and negation. ACM Transactions on Computational Logic, 20(3):18:1–18:??, July 2019. CODEN AT-CLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3329120.

Goranko:2018:GTS

[GKR18] Valentin Goranko, Antti Kuusisto, and Raine Rönnholm. Game-theoretic semantics for alternating-time temporal logic. ACM Transactions on Computational Logic, 19(3): 17:1–17:??, September 2018. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Galesi:2010:OSD

[GL10] Nicola Galesi and Massimo Lauria. Optimality of sizedegree tradeoffs for polynomial calculus. ACM Transactions on Computational Logic, 12(1):4:1–4:??, October 2010. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Gaintzarain:2013:LFM

[GL13] Jose Gaintzarain and Paqui Lucio. Logical foundations for more expressive declarative temporal logic programming languages. ACM Transactions on Computational Logic, 14(4):28:1–28:??, November 2013. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Goller:2014:RPR

[GL14] Stefan Göller and Anthony Widjaja Lin. Refining the process rewrite systems hierarchy via ground tree rewrite systems. ACM Transactions on Computational Logic, 15(4):26:1– 26:??, September 2014. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Geneves:2015:EDC

[GLSG15] Pierre Genevès, Nabil Layaïda, Alan Schmitt, and Nils Gesbert. Efficiently deciding μcalculus with converse over finite trees. ACM Transactions on Computational Logic, 16 (2):16:1–16:??, March 2015. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Givan:2002:PTC

[GM02] Robert Givan and David Mcallester. Polynomial-time computation via local inference relations. ACM Transactions on Computational Logic, 3(4):521–541, October 2002. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Gabbrielli:2009:CSC

[GM09a] Maurizio Gabbrielli and Maria Chiara Meo. A compositional semantics for CHR. ACM Transactions on Computational Logic, 10(2):11:1–11:??, February 2009. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Guerrini:2009:PTC

[GM09b] Stefano Guerrini and Andrea Masini. Proofs, tests and continuation passing style. ACM Transactions on Computational Logic, 10(2):12:1– 12:??, February 2009. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Gaboardi:2012:ICP

[GMR12] Marco Gaboardi, Jean-Yves Marion, and Simona Ronchi Della Rocca. An implicit characterization of PSPACE. ACM Transactions on Computational Logic, 13(2):18:1– 18:??, April 2012. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Gelade:2012:DCF

[GMS12] Wouter Gelade, Marcel Marquardt, and Thomas Schwentick. The dynamic complexity of formal languages. ACM Transactions on Computational Logic, 13(3):19:1–19:??, August 2012. CODEN AT-CLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Goncharov:2020:TUT

[GMS20] Sergey Goncharov, Stefan Milius, and Alexandra Silva. Toward a uniform theory of effectful state machines. ACM Transactions on Computational Logic, 21(3):23:1– 23:63, May 2020. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3372880.

Gurevich:2011:LIP

[GN11] Yuri Gurevich and Itay Neeman. Logic of infons: The propositional case. ACM Transactions on Computational Logic, 12(2):9:1–9:??, January 2011. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Gelade:2012:SCI

[GN12] Wouter Gelade and Frank Neven. Succinctness of the complement and intersection of regular expressions. ACM Transactions on Computational Logic, 13(1):4:1– 4:??, January 2012. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Grohe:2023:CDG

[GN23] Martin Grohe and Daniel Neuen. Canonisation and definability for graphs of bounded rank width. ACM Transactions on Computational Logic, 24(1):6:1–6:??, January 2023. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3568025.

Godoy:2004:CTR

[GNT04] Guillem Godoy, Robert Nieuwenhuis, and Ashish Tiwari. Classes of term rewrite systems with polynomial confluence problems. ACM Transactions on Computational Logic, 5(2):321–331, April 2004. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Ghilardi:2008:CCF

[GNZ08] Silvio Ghilardi, Enrica Nicolini, and Daniele Zucchelli.

A comprehensive combination framework. ACM Transactions on Computational Logic, 9(2):8:1–8:??, March 2008. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Greco:2023:LLP

[GP23] Giuseppe Greco and Alessandra Palmigiano. Linear logic properly displayed. ACM Transactions on Computational Logic, 24(2):13:1–13:??, April 2023. CODEN AT-CLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https://

dl.acm.org/doi/10.1145/
3570919.

Gottlob:2010:MDF

[GPW10] Georg Gottlob, Reinhard Pichler, and Fang Wei. Monadic datalog over finite structures of bounded treewidth. ACM Transactions on Computational Logic, 12(1):3:1–3:??, October 2010. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Giacobazzi:2005:MAD

[GRS05] Roberto Giacobazzi, Francesco Ranzato, and Francesca Scozzari. Making abstract domains condensing. ACM Transactions on Computational Logic, 6(1):33–60, January 2005. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Ganty:2021:CAC

[GRV21] Pierre Ganty, Francesco Ranzato, and Pedro Valero. Complete abstractions for checking language inclusion. ACM Transactions on Computational Logic, 22(4):22:1– 22:40, October 2021. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https://

                                        dl.acm.org/doi/10.1145/
                                        3462673.

Grohe:2000:LOI

[GS00] Martin Grohe and Thomas Schwentick. Locality of orderinvariant first-order formulas. ACM Transactions on Computational Logic, 1(1): 112–130, July 2000. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

                                        tronic). URL http://www.
                                        acm.org/pubs/citations/
                                        journals/tocl/2000-1-1/
                                        p112-grohe/.

Grohe:2002:FOT

[GS02] Martin Grohe and Luc Segoufin. On first-order topological queries. ACM Transactions on Computational Logic, 3(3):336–358, July 2002. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Goranko:2009:TBD

[GS09] Valentin Goranko and Dmitry Shkatov. Tableau-based decision procedures for logics of strategic ability in multiagent systems. ACM Transactions on Computational Logic, 11(1):3:1–3:??, October 2009. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Gastin:2013:FSA

[GS13a] Paul Gastin and Nathalie Sznajder. Fair synthesis for asynchronous distributed systems. ACM Transactions on Computational Logic, 14(2): 9:1–9:??, June 2013. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Gebser:2013:TCL

[GS13b] Martin Gebser and Torsten Schaub. Tableau calculi for logic programs under answer set semantics. ACM Transactions on Computational Logic, 14(2):15:1–15:??, June 2013. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Gallo:2022:LBD

[GSM^{+}22] Fabio R. Gallo, Gerardo I.

Simari, Maria Vanina Martinez, Natalia Abad Santos, and Marcelo A. Falappa. Local belief dynamics in network knowledge bases. ACM Transactions on Computational Logic, 23(1):4:1–4:36, January 2022. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3477394.

Guglielmi:2007:SIS

[Gug07] Alessio Guglielmi. A system of interaction and structure. ACM Transactions on Computational Logic, 8(1):??, January 2007. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Guidi:2009:FS

[Gui09] Ferruccio Guidi. The formal system λδ. ACM Transactions on Computational Logic, 11(1):5:1–5:??, October 2009. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Guidi:2022:FSU

[Gui22] Ferruccio Guidi. A formal system for the universal quantification of schematic variables. ACM Transactions on Computational Logic, 23 (1):2:1–2:37, January 2022. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3470646.

Guller:2018:ADG

[Gul18] Dusan Guller. Automated deduction in Gödel logic. ACM Transactions on Computational Logic, 19(3):18:1– 18:??, September 2018. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Gurevich:2000:SAS

[Gur00] Yuri Gurevich. Sequential abstract-state machines capture sequential algorithms. ACM Transactions on Computational Logic, 1(1):77–111, July 2000. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

http://www.acm.org/pubs/
citations/journals/tocl/
2000-1-1/p77-gurevich/.

Griffioen:2004:TNS

[GV04] David Griffioen and Frits Vaandrager. A theory of normed simulations. ACM Transactions on Computational Logic, 5(4):577–610, October 2004. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Gradel:2022:LMS

[GW22] Erich Grädel and Richard Wilke. Logics with multiteam semantics. ACM Transactions on Computational Logic, 23(2):13:1–13:30, April 2022. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3487579.

Hamano:2018:GIM

[Ham18] Masahiro Hamano. Geometry of interaction for MALL via Hughes–Van Glabbeek proofnets. ACM Transactions on Computational Logic, 19(4): 25:1–25:??, December 2018.

CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3234694.

Hannula:2018:HIL

[Han18] Miika Hannula. Hierarchies in inclusion logic with lax semantics. ACM Transactions on Computational Logic, 19(3):16:1–16:??, September 2018. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Hou:2018:MLS

[HCGT18] Zhé Hóu, Ranald Clouston, Rajeev Goré, and Alwen Tiu. Modular labelled sequent calculi for abstract separation logics. ACM Transactions on Computational Logic, 19(2): 13:1–13:??, June 2018. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Heras:2013:CPH

[HCMS13] Jónathan Heras, Thierry Coquand, Anders Mörtberg, and Vincent Siles. Computing persistent homology within Coq/ SSReflect. ACM Transactions on Computational Logic, 14(4):26:1–26:??, November 2013. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Hesselink:2005:EVP

[Hes05] Wim H. Hesselink. Eternity variables to prove simulation of specifications. ACM Transactions on Computational Logic, 6(1):175–201, January 2005. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Hung:2007:PIX

[HGS07] Edward Hung, Lise Getoor, and V. S. Subrahmanian. Probabilistic interval XML. ACM Transactions on Computational Logic, 8(4):24:1– 24:??, August 2007. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Halamish:2015:MDL

[HK15a] Shulamit Halamish and Orna Kupferman. Minimizing deterministic lattice automata. ACM Transactions on Computational Logic, 16(1):1:1– 1:??, March 2015. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Hampson:2015:UPB

[HK15b] Christopher Hampson and Agi Kurucz. Undecidable propositional bimodal logics and one-variable first-order linear temporal logics with counting. ACM Transactions on Computational Logic, 16(3):27:1–27:??, July 2015. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Hertling:2021:ECL

[HK21] Peter Hertling and Gisela Krommes. EXPSPACEcompleteness of the logics K4 × S5 and S4 × S5 and the logic of subset spaces. ACM Transactions on Computational Logic, 22 (4):24:1–24:71, October 2021. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3465384.

Hella:2020:SMI

[HKMV20] Lauri Hella, Antti Kuusisto, Arne Meier, and Heribert Vollmer. Satisfiability of modal inclusion logic: Lax and strict semantics. ACM Transactions on Computational Logic, 21(1):7:1– 7:18, January 2020. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3356043.

Huang:2019:RAC

[HKO19] Xiaowei Huang, Marta Kwiatkowska, and Maciej Olejnik. Reasoning about cognitive trust in stochastic multiagent systems. ACM Transactions on Computational Logic, 20(4): 21:1–21:??, September 2019. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3329123.

Hannula:2018:CPL

[HKVV18] Miika Hannula, Juha Kontinen, Jonni Virtema, and Heribert Vollmer. Complexity of propositional logics in team semantic. ACM Transactions on Computational Logic, 19 (1):2:1–2:??, February 2018. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Herzig:2013:PUO

[HLM13] Andreas Herzig, Jerome Lang, and Pierre Marquis. Propositional update operators based on formula/literal dependence. ACM Transactions on Computational Logic, 14(3): 24:1–24:??, August 2013. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Hernich:2020:DOM

[HLPW20] André Hernich, Carsten Lutz, Fabio Papacchini, and Frank Wolter. Dichotomies in ontology-mediated querying with the guarded fragment. ACM Transactions on Computational Logic, 21(3):20:1– 20:47, May 2020. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3375628.

Hitchcock:2007:ACD

[HLT07] John M. Hitchcock, Jack H.

Lutz, and Sebastiaan A. Terwijn. The arithmetical complexity of dimension and randomness. ACM Transactions on Computational Logic, 8(2):??, April 2007. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Hague:2017:CPA

[HMOS17] Matthew Hague, Andrzej S. Murawski, C.-H. Luke Ong, and Olivier Serre. Collapsible pushdown automata and recursion schemes. ACM Transactions on Computational Logic, 18(3):25:1–25:??, August 2017. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Henzinger:2005:CST

[HMR05] Thomas A. Henzinger, Rupak Majumdar, and Jean-François Raskin. A classification of symbolic transition systems. ACM Transactions on Computational Logic, 6(1): 1–32, January 2005. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Harland:2003:RDB

[HP03] James Harland and David Pym. Resource-distribution via Boolean constraints. ACM Transactions on Computational Logic, 4(1):56–90, January 2003. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Harper:2005:ECF

[HP05] Robert Harper and Frank Pfenning. On equivalence and canonical forms in the LF type theory. ACM Transactions on Computational Logic, 6(1):61–101, January 2005. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Hu:2022:CTV

[HPS22] Jason Z. S. Hu, Brigitte Pientka, and Ulrich Schöpp. A category theoretic view of contextual types: From simple types to dependent types. ACM Transactions on Computational Logic, 23 (4):25:1–25:??, October 2022. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3545115.

Haret:2017:MHF

[HRW17] Adrian Haret, Stefan Rümmele, and Stefan Woltran. Merging in the Horn fragment. ACM Transactions on Computational Logic, 18(1):6:1– 6:??, April 2017. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Hofmann:2010:PPP

[HS10] Martin Hofmann and Ulrich Schöpp. Pure pointer programs with iteration. ACM Transactions on Computational Logic, 11(4):26:1–26:??, July 2010. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Heule:2015:SAC

[HS15] Marijn J. H. Heule and Stefan Szeider. A SAT approach to clique-width. ACM Transactions on Computational Logic, 16(3):24:1–24:??, July 2015. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Herbreteau:2020:WLT

[HSTW20] Frédéric Herbreteau, B. Srivathsan, Thanh-Tung Tran, and Igor Walukiewicz. Why liveness for timed automata is hard, and what we can do about it. ACM Transactions on Computational Logic, 21(3):17:1–17:28, May 2020. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:

//dl.acm.org/doi/abs/10.
1145/3372310.

Horne:2019:MDN

[HTAC19] Ross Horne, Alwen Tiu, Bogdan Aman, and Gabriel Ciobanu. De Morgan dual nominal quantifiers modelling private names in noncommutative logic. ACM Transactions on Computational Logic, 20(4):22:1–22:??, September 2019. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL https://dl.

acm.org/ft_gateway.cfm?
id=3325821.

Hughes:2005:PNU

[HV05] Dominic J. D. Hughes and Rob J. Van Glabbeek. Proof nets for unit-free multiplicativeadditive linear logic. ACM Transactions on Computational Logic, 6(4):784–842, October 2005. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Huang:2018:ESL

[HV18] Xiaowei Huang and Ron Van Der Meyden. An epistemic strategy logic. ACM Transactions on Computational Logic, 19(4):26:1–26:??, December 2018. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3233769.

Heymans:2008:OAS

[HVV08] Stijn Heymans, Davy Van Nieuwenborgh, and Dirk Vermeir. Open answer set programming with guarded programs. ACM Transactions on Computational Logic, 9(4): 26:1–26:??, August 2008. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Horbach:2010:SFD

[HW10] Matthias Horbach and Christoph Weidenbach. Superposition for fixed domains. ACM Transactions on Computational Logic, 11(4):27:1–27:??, July 2010. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Iocchi:2009:RAA

[ILNR09] Luca Iocchi, Thomas Lukasiewicz, Daniele Nardi, and Riccardo Rosati. Reasoning about actions with sensing under qualitative and probabilistic uncertainty. ACM Transactions on Computational Logic, 10 (1):5:1–5:??, January 2009. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Ignjatovic:2004:SAL

[IS04] Aleksandar Ignjatovic and Arun Sharma. Some applications of logic to feasibility in higher types. ACM Transactions on Computational Logic, 5(2):332–350, April 2004. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Impagliazzo:2006:CDF

[IS06] Russell Impagliazzo and Nathan Segerlind. Constantdepth Frege systems with counting axioms polynomially simulate Nullstellensatz refutations. ACM Transactions on Computational Logic, 7(2): 199–218, April 2006. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Japaridze:2006:PCLa

[Jap06a] Giorgi Japaridze. Propositional computability logic I. ACM Transactions on Computational Logic, 7(2): 302–330, April 2006. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Japaridze:2006:PCLb

[Jap06b] Giorgi Japaridze. Propositional computability logic II. ACM Transactions on Computational Logic, 7(2): 331–362, April 2006. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Jones:2009:FCM

[JK09] Neil D. Jones and Lars Kristiansen. A flow calculus of mwp-bounds for complexity analysis. ACM Transactions on Computational Logic, 10 (4):28:1–28:??, August 2009. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Jurdzinski:2011:AAD

[JL11] Marcin Jurdziński and Ranko Lazić. Alternating automata on data trees and XPath satisfiability. ACM Transactions on Computational Logic, 12(3):19:1–19:??, May 2011. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Janhunen:2006:UPD

[JNS^{+}06] Tomi Janhunen, Ilkka Niemelä, Dietmar Seipel, Patrik Simons, and Jia-Huai You. Unfolding partiality and disjunctions in stable model semantics. ACM Transactions on Computational Logic, 7(1): 1–37, January 2006. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Jouannaud:2015:NHO

[JR15] Jean-Pierre Jouannaud and Albert Rubio. Normal higherorder termination. ACM Transactions on Computational Logic, 16(2):13:1–13:??, March 2015. CODEN AT-CLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Kartzow:2013:FOL

[Kar13] Alexander Kartzow. Firstorder logic on higher-order nested pushdown trees. ACM Transactions on Computational Logic, 14(2):8:1–8:??, June 2013. CODEN AT-CLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Kawamura:2009:DR

[Kaw09] Akitoshi Kawamura. Differential recursion. ACM Transactions on Computational Logic, 10(3):22:1–22:??, April 2009. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Khan:2011:LIS

[KB11] Md. Aquil Khan and Mohua Banerjee. Logics for information systems and their dynamic extensions. ACM Transactions on Computational Logic, 12(4):29:1–29:??, July 2011. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Khan:2023:LTI

[KBP23] Md. Aquil Khan, Mohua Banerjee, and Sibsankar Panda. Logics for temporal information systems in rough set theory. ACM Transactions on Computational Logic, 24(1):2:1–2:??, January 2023. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3549075.

Klinkhamer:2019:VLF

[KE19] Alex Klinkhamer and Ali Ebnenasir. On the verification of livelock-freedom and self-stabilization on parameterized rings. ACM Transactions on Computational Logic, 20(3):16:1–16:??, July 2019. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3326456.

Khaniki:2022:PCR

[Kha22] Erfan Khaniki. On proof complexity of resolution over polynomial calculus. ACM Transactions on Computational Logic, 23(3):16:1–16:24, July 2022. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

https://dl.acm.org/doi/
10.1145/3506702.

Kjos-Hanssen:2012:ACE

[KHST12] Bjørn Kjos-Hanssen, Frank Stephan, and Jason Teutsch. Arithmetic complexity via effective names for random sequences. ACM Transactions on Computational Logic, 13 (3):24:1–24:??, August 2012. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Kojima:2017:HLG

[KI17] Kensuke Kojima and Atsushi Igarashi. A Hoare logic for GPU kernels. ACM Transactions on Computational Logic, 18(1):3:1–3:??, April 2017. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Komosinski:2014:IEA

[KKLJU14] Maciej Komosinski, Adam Kups, Dorota Leszczyńska-Jasion, and Mariusz Urbański. Identifying efficient abductive hypotheses using multicriteria dominance relation. ACM Transactions on Computational Logic, 15(4):28:1– 28:??, September 2014. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Kahler:2010:DSP

[KKW10] Detlef Kähler, Ralf Küsters, and Thomas Wilke. Deciding strategy properties of contract-signing protocols. ACM Transactions on Computational Logic, 11(3):17:1– 17:??, May 2010. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Klaedtke:2008:BAS

[Kla08] Felix Klaedtke. Bounds on the automata size for Presburger arithmetic. ACM Transactions on Computational Logic, 9(2):11:1–11:??, March 2008. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Kaminski:2010:CSI

[KM10] Michael Kaminski and Simone Martini. CSL 2008 special issue. ACM Transactions on Computational Logic, 11(4): 21:1–21:??, July 2010. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Kontchakov:2013:TLC

[KNPHZ13] Roman Kontchakov, Yavor Nenov, Ian Pratt-Hartmann, and Michael Zakharyaschev. Topological logics with connectedness over Euclidean spaces. ACM Transactions on Computational Logic, 14(2): 13:1–13:??, June 2013. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Kolaitis:2005:LSI

[Kol05] Phokion G. Kolaitis. LICS 2003 special issue. ACM Transactions on Computational Logic, 6(4):674, October 2005. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Kontinen:2009:LCC

[Kon09] Juha Kontinen. A logical characterization of the counting hierarchy. ACM Transactions on Computational Logic, 10(1):7:1–7:??, January 2009. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Konovalov:2021:GRB

[Kon21] Aleksandr Yu. Konovalov. Generalized realizability and basic logic. ACM Transactions on Computational Logic, 22(4):25:1–25:23, October 2021. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3468856.

Konovalov:2023:GRI

[Kon23] Aleksandr Konovalov. A generalized realizability and intuitionistic logic. ACM Transactions on Computational Logic, 24(2):11:1–11:??, April 2023. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3565367.

Koutsos:2021:DSS

[Kou21] Adrien Koutsos. Decidability of a sound set of inference rules for computational indistinguishability. ACM Transactions on Computational Logic, 22(1):3:1–3:44, January 2021. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

https://dl.acm.org/doi/
10.1145/3423169.

Kozen:2000:HLK

[Koz00] Dexter Kozen. On Hoare logic and Kleene algebra with tests. ACM Transactions on Computational Logic, 1(1):60–76, July 2000. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

http://www.acm.org/pubs/
citations/journals/tocl/
2000-1-1/p60-kozen/.

Khan:2018:SML

[KP18] Md. Aquil Khan and Vineeta Singh Patel. A simple modal logic for reasoning in multigranulation rough set model. ACM Transactions on Computational Logic, 19(4): 30:1–30:??, December 2018. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3274664.

Kupke:2022:CRG

[KPS22] Clemens Kupke, Dirk Pattinson, and Lutz Schröder. Coalgebraic reasoning with global assumptions in arithmetic modal logics. ACM Transactions on Computational Logic, 23(2):11:1–11:34, April 2022. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3501300.

Kaminski:2002:RQA

[KR02] Michael Kaminski and Guy Rey. Revisiting quantification in autoepistemic logic. ACM Transactions on Computational Logic, 3(4):542– 561, October 2002. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Kramer:2015:LII

[Kra15] Simon Kramer. Logic of intuitionistic interactive proofs (formal theory of perfect knowledge transfer). ACM Transactions on Computational Logic, 16(4):35:1–35:??, November 2015. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Krajicek:2021:SCD

[Kra21] Jan Krajícek. Small circuits and dual weak PHP in the universal theory of p-time algorithms. ACM Transactions on Computational Logic, 22(2):11:1–11:4, June 2021. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3446207.

Krotzsch:2013:CHD

[KRH13] Markus Krötzsch, Sebastian Rudolph, and Pascal Hitzler. Complexities of Horn description logics. ACM Transactions on Computational Logic, 14(1):2:1–2:??, February 2013. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Khoussainov:2005:ALO

[KRS05] Bakhadyr Khoussainov, Sasha Rubin, and Frank Stephan. Automatic linear orders and trees. ACM Transactions on Computational Logic, 6(4): 675–700, October 2005. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Keiren:2012:SAB

[KRW12] Jeroen J. A. Keiren, Michel A.

Reniers, and Tim A. C. Willemse. Structural analysis of Boolean equation systems. ACM Transactions on Computational Logic, 13(1): 8:1–8:??, January 2012. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Kazana:2013:EMS

[KS13a] Wojciech Kazana and Luc Segoufin. Enumeration of monadic second-order queries on trees. ACM Transactions on Computational Logic, 14(4):25:1–25:??, November 2013. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Ketema:2013:LUB

[KS13b] Jeroen Ketema and Jakob Grue Simonsen. Least upper bounds on the size of confluence and Church–Rosser diagrams in term rewriting and λ-calculus. ACM Transactions on Computational Logic, 14(4):31:1–31:??, November 2013. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Kraus:2015:HHH

[KS15] Nicolai Kraus and Christian Sattler. Higher homotopies in a hierarchy of univalent universes. ACM Transactions on Computational Logic, 16(2): 18:1–18:??, March 2015. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Krebs:2017:ECA

[KS17] Andreas Krebs and Howard Straubing. An effective characterization of the alternation hierarchy in two-variable logic. ACM Transactions on Computational Logic, 18(4): 30:1–30:??, December 2017. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Kiefer:2022:GIL

[KSS22] Sandra Kiefer, Pascal Schweitzer, and Erkal Selman. Graphs identified by logics with counting. ACM Transactions on Computational Logic, 23 (1):1:1–1:31, January 2022. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3417515.

Kozen:2003:SLP

[KT03] Dexter Kozen and Jerzy Tiuryn. Substructural logic and partial correctness. ACM Transactions on Computational Logic, 4(3):355–378, July 2003. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Kopczynski:2015:VHF

[KT15] Eryk Kopczyński and Tony Tan. On the variable hierarchy of first-order spectra. ACM Transactions on Computational Logic, 16(2):17:1– 17:??, March 2015. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Kieronski:2018:FST

[KT18] Emanuel Kieroński and Lidia Tendera. Finite satisfiability of the two-variable guarded fragment with transitive guards and related variants. ACM Transactions on Computational Logic, 19(2):

8:1–8:??, June 2018. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Kuznetsov:2021:ALU

[Kuz21]  Stepan Kuznetsov.  Ac-

tion logic is undecidable. ACM Transactions on Computational Logic, 22(2):10:1– 10:26, June 2021. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https://

dl.acm.org/doi/10.1145/
3445810.

Kupferman:2001:WAA

[KV01] Orna Kupferman and Moshe Y. Vardi. Weak alternating automata are not that weak. ACM Transactions on Computational Logic, 2(3):408– 429, 2001. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

http://www.acm.org/pubs/
citations/journals/tocl/
2001-2-3/p408-kupferman/.

Korovin:2005:KBC

[KV05a] Konstantin Korovin and Andrei Voronkov. Knuth–Bendix constraint solving is NPcomplete. ACM Transactions on Computational Logic, 6(2): 361–388, April 2005. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Kupferman:2005:LTB

[KV05b] Orna Kupferman and Moshe Y. Vardi. From linear time to branching time. ACM Transactions on Computational Logic, 6(2):273–294, April 2005. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Kutz:2003:LMS

[KWS^{+}03] Oliver Kutz, Frank Wolter, Holger Sturm, Nobu-Yuki Suzuki, and Michael Zakharyaschev. Logics of metric spaces. ACM Transactions on Computational Logic, 4(2):260–294, April 2003. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Lahav:2013:USF

[LA13] Ori Lahav and Arnon Avron.

A unified semantic framework for fully structural propositional sequent systems. ACM Transactions on Computational Logic, 14(4): 27:1–27:??, November 2013. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Lazic:2011:SAA

[Laz11] Ranko Lazić. Safety alternating automata on data words. ACM Transactions on Computational Logic, 12(2):10:1– 10:??, January 2011. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Lahiri:2007:PAI

[LB07] Shuvendu K. Lahiri and Randal E. Bryant. Predicate abstraction with indexed predicates. ACM Transactions on Computational Logic, 9(1): 4:1–4:??, December 2007. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Li:2018:APA

[LDJ^{+}18] Yongjian Li, Kaiqiang Duan, David N. Jansen, Jun Pang, Lijun Zhang, Yi Lv, and Shaowei Cai. An automatic proving approach to parameterized verification. ACM Transactions on Computational Logic, 19(4):27:1– 27:??, December 2018. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL https://dl.
acm.org/ft_gateway.cfm?
id=3232164.

Lescanne:2018:QAL

[Les18] Pierre Lescanne. Quantitative aspects of linear and affine closed Lambda terms. ACM Transactions on Computational Logic, 19(2):9:1– 9:??, June 2018. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Lavrac:2001:ETA

[LF01] Nada Lavrač and Peter A. Flach. An extended transformation approach to inductive logic programming.

ACM Transactions on Computational Logic, 2(4):458– 494, October 2001. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Liau:2005:MLF

[Lia05] Churn-Jung Liau. A modal logic framework for multiagent belief fusion. ACM Transactions on Computational Logic, 6(1):124–174, January 2005. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Liberatore:2000:CCR

[Lib00a] Paolo Liberatore. Compilability and compact representations of revision of Horn knowledge bases. ACM Transactions on Computational Logic, 1(1):131–161, July 2000. CODEN AT-CLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/tocl/2000-1-1/
p131-liberatore/.

Libkin:2000:LCL

[Lib00b] Leonid Libkin. Logics with counting and local properties. ACM Transactions on Computational Logic, 1(1):33–59, July 2000. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

http://www.acm.org/pubs/

                                        citations/journals/tocl/
                                        2000-1-1/p33-libkin/.

Libkin:2001:LCL

[Lib01] Leonid Libkin. Logics capturing local properties. ACM Transactions on Computational Logic, 2(1):135–153, 2001. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/tocl/
                                        2001-2-1/p135-libkin/p135-
                                        libkin.pdf;   http://www.
                                        acm.org/pubs/citations/
                                        journals/tocl/2001-2-1/
                                        p135-libkin/.

Libkin:2003:VIF

[Lib03] Leonid Libkin. Variable independence for first-order definable constraints. ACM Transactions on Computational Logic, 4(4):431–451, October 2003. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Liberatore:2006:CRD

[Lib06] Paolo Liberatore. Complexity results on DPLL and resolution. ACM Transactions on Computational Logic, 7(1): 84–107, January 2006. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Liberatore:2007:WFS

[Lib07] Paolo Liberatore. Where failsafe default logics fail. ACM Transactions on Computational Logic, 8(2):??, April 2007. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Liberatore:2016:BME

[Lib16] Paolo Liberatore. Belief merging by examples. ACM Transactions on Computational Logic, 17(2):9:1–9:??, March 2016. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Liberatore:2023:MIR

[Lib23] Paolo Liberatore. Mixed iterated revisions: Rationale, algorithms, and complexity. ACM Transactions on Computational Logic, 24(3):27:1– 27:??, July 2023. CODEN ???? ISSN 1529-3785 (print), 1557-945X (electronic). URL

https://dl.acm.org/doi/
10.1145/3583071.

Leino:2015:APS

[LL15] K. Rustan M. Leino and Paqui Lucio. An assertional proof of the stability and correctness of Natural Mergesort. ACM Transactions on Computational Logic, 17(1): 6:1–6:22, December 2015. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Laplante:2007:PAM

[LLM^{+}07] Sophie Laplante, Richard Lassaigne, Frédéric Magniez, Sylvain Peyronnet, and Michel de Rougemont. Probabilistic abstraction for model checking: an approach based on property testing. ACM Transactions on Computational Logic, 8(4):20:1–20:??, August 2007. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Luttgen:2002:IBS

[LM02] Gerald Lüttgen and Michael Mendler. The intuitionism behind Statecharts steps. ACM Transactions on Computational Logic, 3(1):1–41, January 2002. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Larose:2022:QRT

[LMM^{+}22] Benoît Larose, Barnaby Martin, Petar Marković, Daniël Paulusma, Siani Smith, and Stanislav Zivný. QCSP on reflexive tournaments. ACM Transactions on Computational Logic, 23(3):14:1–14:22, July 2022. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

https://dl.acm.org/doi/
10.1145/3508069.

Luck:2017:PCS

[LMS17] Martin Lück, Arne Meier, and Irena Schindler. Parametrised complexity of satisfiability in temporal logic. ACM Transactions on Computational Logic, 18(1):1:1–1:??, April 2017. CODEN ATCLA8.

ISSN 1529-3785 (print), 1557-945X (electronic).

Leone:2019:FQA

[LMTV19] Nicola Leone, Marco Manna, Giorgio Terracina, and Pierfrancesco Veltri. Fast query answering over existential rules. ACM Transactions on Computational Logic, 20 (2):12:1–12:??, April 2019. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3308448.

Lopes:2021:RAP

[LNH21] Bruno Lopes, Cláudia Nalon, and Edward Hermann Haeusler. Reasoning about Petri nets: a calculus based on resolution and dynamic logic. ACM Transactions on Computational Logic, 22(2):9:1– 9:22, June 2021. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https://

dl.acm.org/doi/10.1145/
3441655.

Lodha:2019:SAB

[LOS19] Neha Lodha, Sebastian Ordyniak, and Stefan Szeider. A SAT approach to branchwidth. ACM Transactions on Computational Logic, 20 (3):15:1–15:??, July 2019. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3326159.

Lazic:2016:ZHH

[LOW16] Ranko Lazić, Joël Ouaknine, and James Worrell. Zeno, Hercules, and the Hydra: Safety metric temporal logic is Ackermann-complete. ACM Transactions on Computational Logic, 17(3):16:1–16:??, July 2016. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Lellmann:2019:MSC

[LP19] Björn Lellmann and Elaine Pimentel. Modularisation of sequent calculi for normal and non-normal modalities. ACM Transactions on Computational Logic, 20(2): 7:1–7:??, April 2019. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3288757.

Leone:2006:DSK

[LPF^{+}06] Nicola Leone, Gerald Pfeifer, Wolfgang Faber, Thomas Eiter, Georg Gottlob, Simona Perri, and Francesco Scarcello. The DLV system for knowledge representation and reasoning. ACM Transactions on Computational Logic, 7(3):499–562, July 2006. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

LeRoux:2018:MG

[LPR18] Stéphane Le Roux, Arno Pauly, and Jean-François Raskin. Minkowski games. ACM Transactions on Computational Logic, 19(3):20:1– 20:??, September 2018. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Lifschitz:2001:SEL

[LPV01] Vladimir Lifschitz, David Pearce, and Agustín Valverde. Strongly equivalent logic programs. ACM Transactions on Computational Logic, 2(4): 526–541, October 2001. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Lifschitz:2006:WTM

[LR06] Vladimir Lifschitz and Alexander Razborov. Why are there so many loop formulas? ACM Transactions on Computational Logic, 7(2): 261–268, April 2006. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Liberatore:2007:CPA

[LS07] Paolo Liberatore and Marco Schaerf. Compilability of propositional abduction. ACM Transactions on Computational Logic, 8(1):??, January 2007. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Lazic:2015:NCB

[LS15] Ranko Lazić and Sylvain Schmitz. Nonelementary complexities for branching VASS, MELL, and extensions. ACM Transactions on Computational Logic, 16(3):20:1–20:??, July 2015. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Loyer:2004:HBS

[LSS04] Yann Loyer, Nicolas Spyratos, and Daniel Stamate. Hypothesisbased semantics of logic programs in multivalued logics. ACM Transactions on Computational Logic, 5(3): 508–527, July 2004. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Lonc:2003:FPC

[LT03] Zbigniew Lonc and Miroslaw Truszczyński. Fixed-parameter complexity of semantics for logic programs. ACM Transactions on Computational Logic, 4(1):91–119, January 2003. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Lanotte:2009:PBC

[LT09] Ruggero Lanotte and Simone Tini. Probabilistic bisimulation as a congruence. ACM Transactions on Computational Logic, 10(2):9:1– 9:??, February 2009. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Lukasiewicz:2001:PLP

[Luk01] Thomas Lukasiewicz. Probabilistic logic programming with conditional constraints. ACM Transactions on Computational Logic, 2(3):289– 339, 2001. CODEN AT-CLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/tocl/2001-2-3/
p289-lukasiewicz/.

Lutz:2004:NTC

[Lut04] Carsten Lutz. NEXP TIMEcomplete description logics with concrete domains. ACM Transactions on Computational Logic, 5(4):669– 705, October 2004. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Levy:2012:NUH

[LV12] Jordi Levy and Mateu Villaret. Nominal unification from a higher-order perspective. ACM Transactions on Computational Logic, 13(2): 10:1–10:??, April 2012. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Lomuscio:2000:KMS

[LvdMR00] Alessio R. Lomuscio, Ron van der Meyden, and Mark Ryan. Knowledge in multiagent systems: initial configurations and broadcast. ACM Transactions on Computational Logic, 1(2):247–284, 2000. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

http://www.acm.org/pubs/
citations/journals/tocl/
2000-1-2/p247-lomuscio/.

Lasota:2008:ATA

[LW08] Slawomir Lasota and Igor Walukiewicz. Alternating timed automata. ACM Transactions on Computational Logic, 9(2):10:1–10:??, March 2008. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Legay:2010:ORM

[LW10] Axel Legay and Pierre Wolper. On (Omega-)regular model checking. ACM Transactions on Computational Logic, 12 (1):2:1–2:??, October 2010. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Lumsdaine:2015:LUM

[LW15] Peter Lefanu Lumsdaine and Michael A. Warren. The local universes model: an overlooked coherence construction for dependent type theories. ACM Transactions on Computational Logic, 16(3):23:1– 23:??, July 2015. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Larchey-Wendling:2013:NPS

[LWG13] Dominique Larchey-Wendling and Didier Galmiche. Nondeterministic phase semantics and the undecidability of Boolean BI. ACM Transactions on Computational Logic, 14(1):6:1–6:??, February 2013. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Lin:2007:RCA

[LY07] Fangzhen Lin and Jia-Huai You. Recycling computed answers in rewrite systems for abduction. ACM Transactions on Computational Logic, 8(2):??, April 2007. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Lynch:2005:CLR

[Lyn05] James F. Lynch. Convergence law for random graphs with specified degree sequence. ACM Transactions on Computational Logic, 6(4):727– 748, October 2005. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Lahav:2019:PSC

[LZ19] Ori Lahav and Yoni Zohar.

Pure sequent calculi: Analyticity and decision procedure. ACM Transactions on Computational Logic, 20 (3):13:1–13:??, July 2019. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3319501.

Mamouras:2016:HLD

[Mam16] Konstantinos Mamouras. The Hoare logic of deterministic and nondeterministic monadic recursion schemes. ACM Transactions on Computational Logic, 17(2):13:1– 13:??, March 2016. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Madalinska-Bugaj:2012:GQE

[MBN12] Ewa Madalińska-Bugaj and Linh Anh Nguyen. A generalized QSQR evaluation method for Horn knowledge bases. ACM Transactions on Computational Logic, 13 (4):32:1–32:??, October 2012. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Menghi:2020:MCM

[MBRS20] Claudio Menghi, Marcello M. Bersani, Matteo Rossi, and Pierluigi San Pietro. Model checking MITL formulae on timed automata: a logicbased approach. ACM Transactions on Computational Logic, 21(3):26:1–26:44, May 2020. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

https://dl.acm.org/doi/
abs/10.1145/3383687.

McKinley:2013:PNH

[McK13] Richard McKinley. Proof nets for Herbrand’s theorem. ACM Transactions on Computational Logic, 14(1):5:1– 5:??, February 2013. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Mielniczuk:2004:BTF

[Mie04] Pawel Mielniczuk. Basic theory of feature trees. ACM Transactions on Computational Logic, 5(3):385–402, July 2004. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

McDowell:2002:RHO

[MM02] Raymond C. McDowell and Dale A. Miller. Reasoning with higher-order abstract syntax in a logical framework. ACM Transactions on Computational Logic, 3(1): 80–136, January 2002. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

McIver:2007:RQC

[MM07] Annabelle McIver and Carroll Morgan. Results on the quantitative μ-calculus qMμ. ACM Transactions on Computational Logic, 8(1):??, January 2007. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Madelaine:2012:CPF

[MM12] Florent Madelaine and Barnaby Martin. The complexity of positive first-order logic without equality. ACM Transactions on Computational Logic, 13(1):5:1–5:??, January 2012. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Mogavero:2014:RAS

[MMPV14] Fabio Mogavero, Aniello Murano, Giuseppe Perelli, and Moshe Y. Vardi. Reasoning about strategies: On the model-checking problem. ACM Transactions on Computational Logic, 15(4):34:1– 34:??, November 2014. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Mahmood:2023:PCL

[MMS23] Yasir Mahmood, Arne Meier, and Johannes Schmidt. Parameterized complexity of logic-based argumentation in Schaefer’s framework. ACM Transactions on Computational Logic, 24(3):26:1–26:??, July 2023. CODEN ???? ISSN 1529-3785 (print), 1557-945X (electronic). URL

https://dl.acm.org/doi/
10.1145/3582499.

Martini:2021:SLN

[MMZ21] Simone Martini, Andrea Masini, and Margherita Zorzi. From 2-sequents and linear nested sequents to natural deduction for normal modal logics. ACM Transactions on Computational Logic, 22 (3):19:1–19:29, June 2021. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3461661.

More:2012:CCG

[MN12] Sara Miner More and Pavel Naumov. Calculus of cooperation and game-based reasoning about protocol privacy. ACM Transactions on Computational Logic, 13(3):22:1– 22:??, August 2012. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Montenegro:2020:ELT

[MNPS20] Manuel Montenegro, Susana Nieva, Ricardo Peña, and Clara Segura. Extending liquid types to arrays. ACM Transactions on Computational Logic, 21(2):13:1– 13:41, March 2020. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3362740.

Murawski:2006:FVM

[MO06] Andrzej S. Murawski and C.- H. Luke Ong. Fast verification of MLL proof nets via IMLL. ACM Transactions on Computational Logic, 7(3):

473–498, July 2006. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Makino:2012:DII

[MO12] Kazuhisa Makino and Hirotaka Ono. Deductive inference for the interiors and exteriors of Horn theories. ACM Transactions on Computational Logic, 13(3):23:1– 23:??, August 2012. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Metcalfe:2005:SHC

[MOG05] George Metcalfe, Nicola Olivetti, and Dov Gabbay. Sequent and hypersequent calculi for abelian and Łukasiewicz logics. ACM Transactions on Computational Logic, 6(3):578–613, July 2005. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Michaliszyn:2015:DEM

[MOK15] Jakub Michaliszyn, Jan Otop, and Emanuel Kieroński. On the decidability of elementary modal logics. ACM Transactions on Computational Logic, 17(1):2:1–2:??, December 2015. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Moyen:2009:RCG

[Moy09] Jean-Yves Moyen. Resource control graphs. ACM Transactions on Computational Logic, 10(4):29:1–29:??, August 2009. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Momigliano:2003:HOP

[MP03] Alberto Momigliano and Frank Pfenning. Higher-order pattern complement and the strict λ-calculus. ACM Transactions on Computational Logic, 4(4):493–529, October 2003. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Marion:2009:SIS

[MP09] Jean-Yves Marion and Romain Péchoux. Sup-interpretations, a semantic method for static analysis of program resources. ACM Transactions on Computational Logic, 10(4):27:1– 27:??, August 2009. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Matheja:2023:DPG

[MPZ23] Christoph Matheja, Jens Pagel, and Florian Zuleger. A decision procedure for guarded separation logic complete entailment checking for separation logic with inductive definitions. ACM Transactions on Computational Logic, 24(1):1:1–1:??, January 2023. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3534927.

Mesnard:2003:PLT

[MR03] Fred Mesnard and Salvatore Ruggieri. On proving left termination of constraint logic programs. ACM Transactions on Computational Logic, 4(2): 207–259, April 2003. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Molinaro:2014:SSS

[MSS14] Cristian Molinaro, Amy Sliva, and V. S. Subrahmanian. Super-solutions: Succinctly representing solutions in abductive annotated probabilistic temporal logic. ACM Transactions on Computational Logic, 15(3):18:1–18:??, September 2014. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Miller:2005:PTG

[MT05] Dale Miller and Alwen Tiu. A proof theory for generic judgments. ACM Transactions on Computational Logic, 6(4): 749–783, October 2005. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Murawski:2005:AUP

[Mur05] Andrzej S. Murawski. About the undecidability of program equivalence in finitary languages with state.

ACM Transactions on Computational Logic, 6(4):701– 726, October 2005. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Michel:2004:DBI

[MV04] Laurent Michel and Pascal Van Hentenryck. A decomposition-based implementation of search strategies. ACM Transactions on Computational Logic, 5(2): 351–383, April 2004. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Marchioni:2015:LGL

[MW15] Enrico Marchioni and Michael Wooldridge. Lukasiewicz games: a logic-based approach to quantitative strategic interactions. ACM Transactions on Computational Logic, 16(4):33:1–33:??, November 2015. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Mastroeni:2017:APS

[MZ17] Isabella Mastroeni and Damiano Zanardini. Abstract program slicing: an abstract interpretation-based approach to program slicing. ACM Transactions on Computational Logic, 18(1): 7:1–7:??, April 2017. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Nguyen:2012:CPD

[NC12] Phuong Nguyen and Stephen Cook. The complexity of proving the Discrete Jordan Curve Theorem. ACM Transactions on Computational Logic, 13(1):9:1–9:??, January 2012. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Nalon:2019:MRP

[NDH19] Cláudia Nalon, Clare Dixon, and Ullrich Hustadt. Modal resolution: Proofs, layers, and refinements. ACM Transactions on Computational Logic, 20(4):23:1–23:??, September 2019. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

                                        tronic). URL https://dl.
                                        acm.org/ft_gateway.cfm?
                                        id=3331448.

Nielson:2011:MAC

[NNN11] Flemming Nielson, Sebastian Nanz, and Hanne Riis Nielson. Modal abstractions of concurrent behavior. ACM Transactions on Computational Logic, 12(3):18:1–18:??, May 2011. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Nordstrom:2012:RSP

[Nor12] Jakob Nordström. On the relative strength of pebbling and resolution. ACM Transactions on Computational Logic, 13(2):16:1–16:??, April 2012. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Nanevski:2008:CMT

[NPP08] Aleksandar Nanevski, Frank Pfenning, and Brigitte Pientka. Contextual modal type theory. ACM Transactions on Computational Logic, 9(3): 23:1–23:??, June 2008. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Novakovic:2015:PSC

[NS15] Novak Novaković and Lutz Straßburger. On the power of substitution in the calculus of structures. ACM Transactions on Computational Logic, 16(3):19:1–19:??, July 2015. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Neider:2018:CSP

[NSM18] Daniel Neider, Shambwaditya Saha, and P. Madhusudan. Compositional synthesis of piece-wise functions by learning classifiers. ACM Transactions on Computational Logic, 19(2):10:1–10:??, June 2018. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Neven:2018:FSM

[NSST18] Frank Neven, Nicole Schweikardt, Frederic Servais, and Tony Tan. Finite-state map-reduce computation and relational algebra queries. ACM Transactions on Computational Logic, 19(2):14:1–14:??, June 2018. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Neven:2004:FSM

[NSV04] Frank Neven, Thomas Schwentick, and Victor Vianu. Finite state machines for strings over infinite alphabets. ACM Transactions on Computational Logic, 5(3):403–435, July 2004. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Namjoshi:2010:CCR

[NT10] Kedar S. Namjoshi and Richard J. Trefler. On the completeness of compositional reasoning methods. ACM Transactions on Computational Logic, 11(3):16:1–16:??, May 2010. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Naumov:2017:IFU

[NT17] Pavel Naumov and Jia Tao. Information flow under budget constraints. ACM Transactions on Computational Logic, 18(4):32:1–32:??, December 2017. CODEN AT-CLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Newton:2019:TNA

[NV19] Jim Newton and Didier Verna. A theoretical and numerical analysis of the worst-case size of reduced ordered binary decision diagrams. ACM Transactions on Computational Logic, 20 (1):6:1–6:??, February 2019. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3274279.

Olivetti:2007:SCT

[OPS07] Nicola Olivetti, Gian Luca Pozzato, and Camilla B. Schwind. A sequent calculus and a theorem prover for standard conditional logics. ACM Transactions on Computational Logic, 8(4):22:1– 22:??, August 2007. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Oetsch:2021:BUE

[OSTW21] Johannes Oetsch, Martina Seidl, Hans Tompits, and Stefan Woltran. Beyond uniform equivalence between answerset programs. ACM Transactions on Computational Logic, 22(1):2:1–2:46, January 2021. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3422361.

Paulson:2000:MUI

[Pau00] Lawrence C. Paulson. Mechanizing UNITY in Isabelle. ACM Transactions on Computational Logic, 1(1):3–32, July 2000. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/tocl/
                                        2000-1-1/p3-paulson/p3-
                                        paulson.pdf;   http://www.
                                        acm.org/pubs/citations/
                                        journals/tocl/2000-1-1/
                                        p3-paulson/.

Paulson:2006:DFE

[Pau06] Lawrence C. Paulson. Defining functions on equivalence classes. ACM Transactions on Computational Logic, 7(4): 658–675, October 2006. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Poza:2014:CRS

[PDHR14] María Poza, César Domínguez, Jónathan Heras, and Julio Rubio. A certified reduction strategy for homological image processing. ACM Transactions on Computational Logic, 15(3):23:1–23:??, September 2014. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Pientka:2009:HOT

[Pie09] Brigitte Pientka. Higher-order term indexing using substitution trees. ACM Transactions on Computational Logic, 11(1):6:1–6:??, October 2009. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Pitts:2020:THE

[Pit20] Andrew M. Pitts. Typal heterogeneous equality types. ACM Transactions on Computational Logic, 21(3):25:1– 25:10, May 2020. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3379447.

Peppas:2012:MMB

[PKW12] Pavlos Peppas, Costas D.

Koutras, and Mary-Anne Williams. Maps in multiple belief change. ACM Transactions on Computational Logic, 13(4):30:1–30:??, October 2012. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Platzer:2015:DGL

[Pla15] André Platzer. Differential game logic. ACM Transactions on Computational Logic, 17(1):1:1–1:??, December 2015. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Platzer:2017:DHG

[Pla17] André Platzer. Differential hybrid games. ACM Transactions on Computational Logic, 18(3):19:1–19:??, August 2017. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Pedicini:2007:PPE

[PQ07] Marco Pedicini and Francesco Quaglia. PELCR: Parallel environment for optimal lambda-calculus reduction. ACM Transactions on Computational Logic, 8(3): 14:1–14:??, July 2007. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Padmanabha:2023:DFF

[PR23] Anantha Padmanabha and R. Ramanujam. A decidable fragment of first order modal logic: Two variable term modal logic. ACM Transactions on Computational Logic, 24(4):29:1–29:??, October 2023. CODEN ???? ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3593584.

Praveen:2013:DTH

[Pra13] M. Praveen. Does treewidth help in modal satisfiability? ACM Transactions on Computational Logic, 14(3):18:1– 18:??, August 2013. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Ponomaryov:2017:PDL

[PS17] Denis Ponomaryov and Mikhail Soutchanski. Progression of decomposed local-effect action theories. ACM Transactions on Computational Logic, 18(2):16:1–16:??, June 2017. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Pakusa:2018:DCF

[PSS18] Wied Pakusa, Svenja Schalthöfer, and Erkal Selman. Definability of Cai–Fürer–Immerman problems in choiceless polynomial time. ACM Transactions on Computational Logic, 19 (2):7:1–7:??, June 2018. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Place:2020:AST

[PZ20] Thomas Place and Marc Zeitoun. Adding successor: a transfer theorem for separation and covering. ACM Transactions on Computational Logic, 21(2):9:1– 9:45, March 2020. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL https://dl.
acm.org/doi/abs/10.1145/
3356339.

Podder:2022:ZOL

[PZ22] Moumanti Podder and Maksim Zhukovskii. Zero–one laws for existential first-order sentences of bounded quantifier depth. ACM Transactions on Computational Logic, 23(2):12:1–12:27, April 2022. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3489466.

Quaas:2019:SDW

[QS19] Karin Quaas and Mahsa Shirmohammadi. Synchronizing data words for register automata. ACM Transactions on Computational Logic, 20 (2):11:1–11:??, April 2019. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:

                                        //dl.acm.org/ft_gateway.
                                        cfm?id=3309760.

Rabinovich:2007:CL

[Rab07] Alexander Rabinovich. On compositionality and its limitations. ACM Transactions on Computational Logic, 8 (1):??, January 2007. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Rabe:2015:LTM

[Rab15] Florian Rabe. Lax theory morphisms. ACM Transactions on Computational Logic, 17(1):5:1–5:??, December 2015. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Rabe:2018:MTR

[Rab18]  Florian Rabe.  A modu-

lar type reconstruction algorithm. ACM Transactions on Computational Logic, 19(4): 24:1–24:??, December 2018.

CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3234693.

Ratschan:2004:CAS

[Rat04] Stefan Ratschan. Convergent approximate solving of firstorder constraints by approximate quantifiers. ACM Transactions on Computational Logic, 5(2):264–281, April 2004. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Ratschan:2006:ESQ

[Rat06] Stefan Ratschan. Efficient solving of quantified inequality constraints over the real numbers. ACM Transactions on Computational Logic, 7(4): 723–748, October 2006. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Reiter:2001:KBP

[Rei01] Ray Reiter. On knowledgebased programming with sensing in the situation calculus. ACM Transactions on Computational Logic, 2(4): 433–457, October 2001. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Radcliffe:2017:UNF

[RMV17] Nicholas R. Radcliffe, Luis F. T. Moraes, and Rakesh M. Verma. Uniqueness of normal forms for shallow term rewrite systems. ACM Transactions on Computational Logic, 18(2):17:1–17:??, June 2017. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Rabe:2013:LRL

[RS13] Florian Rabe and Kristina Sojakova. Logical relations for a logical framework. ACM Transactions on Computational Logic, 14(4):32:1– 32:??, November 2013. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Riguzzi:2014:TEL

[RS14] Fabrizio Riguzzi and Terrance Swift. Terminating evaluation of logic programs with finite three-valued models. ACM Transactions on Computational Logic, 15(4): 32:1–32:??, September 2014. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Rybina:2001:DPT

[RV01] Tatiana Rybina and Andrei Voronkov. A decision procedure for term algebras with queues. ACM Transactions on Computational Logic, 2(2):155–181, 2001. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

http://www.acm.org/pubs/
articles/journals/tocl/

2001-2-2/p155-rybina/p155-
rybina.pdf;   http://www.
acm.org/pubs/citations/
journals/tocl/2001-2-2/
p155-rybina/.

Rondogiannis:2005:MMS

[RW05] Panos Rondogiannis and William W. Wadge. Minimum model semantics for logic programs with negationas-failure. ACM Transactions on Computational Logic, 6(2): 441–467, April 2005. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Sietsma:2013:CKE

[SA13] Floor Sietsma and Krzysztof R.

Apt. Common knowledge in email exchanges. ACM Transactions on Computational Logic, 14(3):23:1–23:??, August 2013. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Sakama:2005:IAS

[Sak05] Chiaki Sakama. Induction from answer sets in nonmonotonic logic programs. ACM Transactions on Computational Logic, 6(2):203– 231, April 2005. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Salibra:2003:TIO

[Sal03] Antonino Salibra. Topological incompleteness and order incompleteness of the lambda calculus. ACM Transactions on Computational Logic, 4(3): 379–401, July 2003. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Sangiorgi:2017:ECU

[San17] Davide Sangiorgi. Equations, contractions, and unique solutions. ACM Transactions on Computational Logic, 18 (1):4:1–4:??, April 2017. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Saurin:2010:TSC

[Sau10] Alexis Saurin. Typing streams in the Λμ-calculus. ACM Transactions on Computational Logic, 11(4):28:1–28:??, July 2010. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Shakarian:2013:UGA

[SBSM13] Paulo Shakarian, Matthias Broecheler, V. S. Subrahmanian, and Cristian Molinaro. Using generalized annotated programs to solve social network diffusion optimization problems. ACM Transactions on Computational Logic, 14(2):10:1–10:??, June 2013. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Son:2006:DDK

[SBTM06] Tran Cao Son, Chitta Baral, Nam Tran, and Sheila Mcilraith. Domain-dependent knowledge in answer set planning. ACM Transactions on Computational Logic, 7(4): 613–657, October 2006. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Schweikardt:2005:AFO

[Sch05] Nicole Schweikardt. Arithmetic, first-order logic, and counting quantifiers. ACM Transactions on Computational Logic, 6(3):634–671, July 2005. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Schwartz:2015:DRS

[Sch15] Daniel G. Schwartz. Dynamic reasoning systems. ACM Transactions on Computational Logic, 16(4):32:1– 32:??, November 2015. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Simari:2013:PAQ

[SDSS13] Gerardo I. Simari, John P.

Dickerson, Amy Sliva, and V. S. Subrahmanian. Parallel abductive query answering in probabilistic logic programs. ACM Transactions on Computational Logic, 14(2): 12:1–12:??, June 2013. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Schellhorn:2014:SCP

[SDW14] Gerhard Schellhorn, John Derrick, and Heike Wehrheim. A sound and complete proof technique for linearizability of concurrent data structures. ACM Transactions on Computational Logic, 15(4):31:1– 31:??, September 2014. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Seiller:2018:IGN

[Sei18] Thomas Seiller. Interaction graphs: Non-deterministic automata. ACM Transactions on Computational Logic, 19(3):21:1–21:??, September 2018. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Sergot:2001:CTN

[Ser01] Marek Sergot. A computational theory of normative positions. ACM Transactions on Computational Logic, 2(4): 581–622, October 2001. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Su:2017:PJL

[SFL17] Che-Ping Su, Tuan-Fang Fan, and Churn-Jung Liau. Possibilistic justification logic: Reasoning about justified uncertain beliefs. ACM Transactions on Computational Logic, 18(2):15:1–15:??, June 2017. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Shoham:2007:GBF

[SG07] Sharon Shoham and Orna Grumberg. A game-based framework for CTL counterexamples and 3-valued abstraction-refinement. ACM Transactions on Computational Logic, 9(1):1:1–1:??, December 2007. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Stone:2006:EES

[SH06] Christopher A. Stone and Robert Harper. Extensional equivalence and singleton types. ACM Transactions on Computational Logic, 7(4): 676–722, October 2006. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Schmidt:2007:ATP

[SH07] Renate A. Schmidt and Ullrich Hustadt. The axiomatic translation principle for modal logic. ACM Transactions on Computational Logic, 8(4):19:1–19:??, August 2007. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Shapirovsky:2022:SPS

[Sha22] Ilya Shapirovsky. Satisfiability problems on sums of Kripke frames. ACM Transactions on Computational Logic, 23(3):15:1–15:25, July 2022. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3508068.

Shen:2008:RRL

[She08] Yi-Dong Shen. Reasoning with recursive loops under the PLP framework. ACM Transactions on Computational Logic, 9(4):27:1–27:??, August 2008. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Sakama:2008:CAS

[SI08] Chiaki Sakama and Katsumi Inoue. Coordination in answer set programming. ACM Transactions on Computational Logic, 9(2):9:1– 9:??, March 2008. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Simmons:2014:SF

[Sim14] Robert J. Simmons. Structural focalization. ACM Transactions on Computational Logic, 15(3):21:1–21:??, September 2014. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Simonsen:2015:CRS

[Sim15] Jakob Grue Simonsen. A confluent rewriting system having no computable, onestep, normalizing strategy.

ACM Transactions on Computational Logic, 16(2):10:1– 10:??, February 2015. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Schockaert:2012:FEL

[SJV12] Steven Schockaert, Jeroen Janssen, and Dirk Vermeir. Fuzzy equilibrium logic: Declarative problem solving in continuous domains. ACM Transactions on Computational Logic, 13(4):33:1– 33:??, October 2012. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Schneider-Kamp:2009:ATP

[SKGST09] Peter Schneider-Kamp, Jürgen Giesl, Alexander Serebrenik, and René Thiemann. Automated termination proofs for logic programs by term rewriting. ACM Transactions on Computational Logic, 11 (1):2:1–2:??, October 2009. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Sojakova:2016:ETP

[Soj16]  Kristina Sojakova.  The

equivalence of the torus and the product of two circles in homotopy type theory. ACM Transactions on Computational Logic, 17(4):29:1– 29:??, November 2016. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Schroder:2009:PBR

[SP09] Lutz Schröder and Dirk Pattinson. PSPACE bounds for rank-1 modal logics. ACM Transactions on Computational Logic, 10(2):13:1– 13:??, February 2009. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Santo:2011:CMS

[SP11] José Espírito Santo and Luís Pinto. A calculus of multiary sequent terms. ACM Transactions on Computational Logic, 12(3):22:1–22:??, May 2011. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Skarlatidis:2015:PEC

[SPAV15] Anastasios Skarlatidis, Georgios Paliouras, Alexander Artikis, and George A. Vouros. Probabilistic event calculus for event recognition. ACM Transactions on Computational Logic, 16(2):11:1– 11:??, February 2015. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Son:2014:FNU

[SPNS14] Tran Cao Son, Enrico Pontelli, Ngoc-Hieu Nguyen, and Chiaki Sakama. Formalizing negotiations using logic programming. ACM Transactions on Computational Logic, 15(2):12:1–12:??, April 2014. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Shakarian:2011:APT

[SPSS11] Paulo Shakarian, Austin Parker, Gerardo Simari, and Venkatramana V. S. Subrahmanian. Annotated probabilistic temporal logic. ACM Transactions on Computational Logic, 12(2):14:1–14:??, January 2011. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Sofronie-Stokkermans:2007:UBD

[SS07] Viorica Sofronie-Stokkermans.

On unification for bounded distributive lattices. ACM Transactions on Computational Logic, 8(2):??, April 2007. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Shakarian:2012:APT

[SSS12] Paulo Shakarian, Gerardo I.

Simari, and V. S. Subrahmanian. Annotated probabilistic temporal logic: Approximate fixpoint implementation. ACM Transactions on Computational Logic, 13(2): 13:1–13:??, April 2012. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Schirrmacher:2023:FOL

[SSV23] Nicole Schirrmacher, Sebastian Siebertz, and Alexandre Vigny. First-order logic with connectivity operators. ACM Transactions on Computational Logic, 24(4):30:1– 30:??, October 2023. CODEN ???? ISSN 1529-3785 (print), 1557-945X (electronic). URL

                                        https://dl.acm.org/doi/
                                        10.1145/3595922.

Schmidt:2014:UTD

[ST14] Renate A. Schmidt and Dmitry Tishkovsky. Using tableau to decide description logics with full role negation and identity. ACM Transactions on Computational Logic, 15(1):7:1–7:??, February 2014. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Sebastiani:2015:OMT

[ST15] Roberto Sebastiani and Silvia Tomasi. Optimization modulo theories with linear rational costs. ACM Transactions on Computational Logic, 16(2): 12:1–12:??, February 2015. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Stone:2005:DMG

[Sto05] Matthew Stone. Disjunction and modular goal-directed proof search. ACM Transactions on Computational Logic, 6(3):539–577, July 2005. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Subramani:2004:OLT

[Sub04]  K. Subramani.  Optimal

length tree-like resolution refutations for 2SAT formulas. ACM Transactions on Computational Logic, 5(2): 316–320, April 2004. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Schubert:2016:HHP

[SUWC16] Aleksy Schubert, Pawel Urzyczyn, and Daria Walukiewicz-Chrzaszcz. How hard is positive quantification? ACM Transactions on Computational Logic, 17(4):30:1–30:??, November 2016. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Seidl:2008:FOV

[SV08] Helmut Seidl and Kumar Neeraj Verma. Flat and one-variable clauses: Complexity of verifying cryptographic protocols with single blind copying. ACM Transactions on Computational Logic, 9(4):28:1– 28:??, August 2008. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Samer:2010:DLS

[SV10] Marko Samer and Helmut Veith. On the distributivity of LTL specifications. ACM Transactions on Computational Logic, 11(3):20:1– 20:??, May 2010. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Schroder:2018:CFC

[SV18] Lutz Schröder and Yde Venema. Completeness of flat coalgebraic fixpoint logics. ACM Transactions on Computational Logic, 19(1):4:1– 4:??, February 2018. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Shen:2003:DAC

[SYY^{+}03] Yi-Dong Shen, Jia-Huai You, Li-Yan Yuan, Samuel S. P. Shen, and Qiang Yang. A dynamic approach to characterizing termination of general logic programs. ACM Transactions on Computational Logic, 4(4):417–430, October 2003. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Szeider:2011:MSO

[Sze11] Stefan Szeider. Monadic second order logic on graphs with local cardinality constraints. ACM Transactions on Computational Logic, 12(2):12:1– 12:??, January 2011. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Takagi:2023:SAL

[Tak23]  Tsubasa Takagi.  Seman-

tic analysis of a linear temporal extension of quantum logic and its dynamic aspect. ACM Transactions on Computational Logic, 24(3):19:1– 19:??, July 2023. CODEN ???? ISSN 1529-3785 (print), 1557-945X (electronic). URL

https://dl.acm.org/doi/
10.1145/3576926.

Tan:2013:GRP

[Tan13] Tony Tan. Graph reachability and pebble automata over infinite alphabets. ACM Transactions on Computational Logic, 14(3):19:1–19:??, August 2013. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Tan:2014:ETV

[Tan14] Tony Tan. Extending twovariable logic on data trees with order on data values and its automata. ACM Transactions on Computational Logic, 15(1):8:1–8:??, February 2014. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Tubella:2018:SPS

[TG18] Andrea Aler Tubella and Alessio Guglielmi. Subatomic proof systems: Splittable systems. ACM Transactions on Computational Logic, 19(1): 5:1–5:??, February 2018. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Tiu:2010:PSS

[TM10] Alwen Tiu and Dale Miller. Proof search specifications of bisimulation and modal logics for the π-calculus. ACM Transactions on Computational Logic, 11(2):13:1– 13:??, January 2010. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Tomovic:2020:FOL

[TOD20] S. Tomović, Z. Ognjanović, and D. Doder. A first-order logic for reasoning about knowledge and probability. ACM Transactions on Computational Logic, 21(2):16:1– 16:30, March 2020. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

                                        tronic). URL https://dl.
                                        acm.org/doi/abs/10.1145/
                                        3359752.

Tripakis:2009:CTB

[Tri09] Stavros Tripakis. Checking timed Büchi automata emptiness on simulation graphs. ACM Transactions on Computational Logic, 10(3):15:1– 15:??, April 2009. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Trybus:2016:RRB

[Try16]  Adam Trybus.  Rational

region-based affine logic of the real plane. ACM Transactions on Computational Logic, 17(3):21:1–21:??, July 2016. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Tao:2015:CFS

[TSH15] Jia Tao, Giora Slutzki, and Vasant Honavar. A conceptual framework for secrecypreserving reasoning in knowledge bases. ACM Transactions on Computational Logic, 16(1):3:1–3:??, March 2015. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Toran:2023:NVG

[TW23] Jacobo Torán and Florian Wörz. Number of variables for graph differentiation and the resolution of graph isomorphism formulas. ACM Transactions on Computational Logic, 24(3):23:1–23:??, July 2023. CODEN ???? ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

/dl.acm.org/doi/10.1145/
3580478.

Tucker:2002:ACA

[TZ02] J. V. Tucker and J. I.

Zucker. Abstract computability and algebraic specification. ACM Transactions on Computational Logic, 3(2): 279–333, April 2002. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Tucker:2004:AVC

[TZ04] J. V. Tucker and J. I. Zucker. Abstract versus concrete computation on metric partial algebras. ACM Transactions on Computational Logic, 5(4): 611–668, October 2004. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Urban:2011:MML

[UCB11] Christian Urban, James Cheney, and Stefan Berghofer. Mechanizing the metatheory of LF. ACM Transactions on Computational Logic, 12 (2):15:1–15:??, January 2011. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Udrea:2010:AR

[URS10] Octavian Udrea, Diego Reforgiato Recupero, and V. S. Subrahmanian. Annotated RDF. ACM Transactions on Computational Logic, 11 (2):10:1–10:??, January 2010. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

VanBakel:2018:CNP

[Van18a] Steffen Van Bakel. Characterisation of normalisation properties for λμ using strict negated intersection types. ACM Transactions on Computational Logic, 19(1):3:1– 3:??, February 2018. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

VanDenBerg:2018:PCP

[Van18b] Benno Van Den Berg. Path categories and propositional identity types. ACM Transactions on Computational Logic, 19(2):15:1–15:??, June 2018. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Vukmirovic:2023:SIE

[VBH23] Petar Vukmirović, Jasmin Blanchette, and Marijn J. H. Heule. SAT-inspired eliminations for superposition. ACM Transactions on Computational Logic, 24(1):7:1– 7:??, January 2023. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https://

dl.acm.org/doi/10.1145/
3565366.

vandenBerg:2012:TSM

[vdBG12] Benno van den Berg and Richard Garner. Topological and simplicial models of identity types. ACM Transactions on Computational Logic, 13(1):3:1–3:??, January 2012. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Verbaeten:2001:TPL

[VDS01] Sofie Verbaeten, Danny De Schreye, and Konstantinos Sagonas. Termination proofs for logic programs with tabling. ACM Transactions on Computational Logic, 2(1):57–92, 2001. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL

http://www.acm.org/pubs/
articles/journals/tocl/
2001-2-1/p57-verbaeten/
p57-verbaeten.pdf;   http:
//www.acm.org/pubs/citations/
journals/tocl/2001-2-1/
p57-verbaeten/.

Vennekens:2006:SOA

[VGD06] Joost Vennekens, David Gilis, and Marc Denecker. Splitting an operator: Algebraic modularity results for logics with fixpoint semantics. ACM Transactions on Computational Logic, 7(4):765– 797, October 2006. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). See erratum [VGD07].

Vennekens:2007:ESO

[VGD07] Joost Vennekens, David Gilis, and Marc Denecker. Erratum to splitting an operator: Algebraic modularity results for logics with fixpoint semantics. ACM Transactions on Computational Logic, 8(1):??, January 2007. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). See [VGD06].

Verma:2005:NDT

[VH05] Rakesh Verma and Ara Hayrapetyan. A new decidability technique for ground term rewriting systems with applications. ACM Transactions on Computational Logic, 6(1):102–123, January 2005. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

vanHentenryck:2000:SSO

[vHPP00] Pascal van Hentenryck, Laurent Perron, and Jean-François Puget. Search and strategies in OPL. ACM Transactions on Computational Logic, 1(2):285– 320, 2000. CODEN AT-CLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL http://www.
acm.org/pubs/citations/
journals/tocl/2000-1-2/
p285-van_hentenryck/.

VanDenDries:2009:AC

[VM09] Lou Van Den Dries and Yiannis N. Moschovakis. Arithmetic complexity. ACM Transactions on Computational Logic, 10(1):2:1–2:??, January 2009. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Voronkov:2001:HOP

[Vor01]  Andrei Voronkov.  How

to optimize proof-search in modal logics: new methods of proving redundancy criteria for sequent calculi. ACM Transactions on Computational Logic, 2(2):182– 215, 2001. CODEN ATCLA8.

ISSN 1529-3785 (print), 1557-945X (electronic). URL

                                        http://www.acm.org/pubs/
                                        articles/journals/tocl/
                                        2001-2-2/p182-voronkov/
                                        p182-voronkov.pdf;   http:
                                        //www.acm.org/pubs/citations/
                                        journals/tocl/2001-2-2/
                                        p182-voronkov/.

VanDerMeyden:2020:UCM

[VP20] Ron Van Der Meyden and Manas K. Patra. Undecidable cases of model checking probabilistic temporal–epistemic logic. ACM Transactions on Computational Logic, 21 (4):32:1–32:26, October 2020. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/
                                        3409250.

VandenBussche:2007:PTI

[VV07] Jan Van den Bussche and Stijn Vansummeren. Polymorphic type inference for the named nested relational calculus. ACM Transactions on Computational Logic, 9(1): 3:1–3:??, December 2007. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Verbitsky:2019:TBA

[VZ19] Oleg Verbitsky and Maksim Zhukovskii. Tight bounds on the asymptotic descriptive complexity of subgraph isomorphism. ACM Transactions on Computational Logic, 20(2):9:1–9:??, April 2019. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https:

//dl.acm.org/ft_gateway.
cfm?id=3303881.

Wittocx:2013:CPF

[WDB13] Johan Wittocx, Marc Denecker, and Maurice Bruynooghe. Constraint propagation for first-order logic and inductive definitions. ACM Transactions on Computational Logic, 14(3):17:1–17:??, August 2013. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Wang:2023:WAS

[WEZL23] Yisong Wang, Thomas Eiter, Yuanlin Zhang, and Fangzhen Lin. Witnesses for answer sets of logic programs. ACM Transactions on Computational Logic, 24(2):15:1– 15:??, April 2023. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic). URL https://

dl.acm.org/doi/10.1145/
3568955.

Wrona:2019:CMI

[Wro19] Michał Wrona. The complexity of minimal inference problem for conservative constraint languages. ACM Transactions on Computational Logic, 20(2):8:1– 8:??, April 2019. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (elec-

tronic). URL https://dl.

                                        acm.org/ft_gateway.cfm?
                                        id=3301410.

Wadler:2003:MEM

[WT03] Philip Wadler and Peter Thiemann. The marriage of effects and monads. ACM Transactions on Computational Logic, 4(1):1–32, January 2003. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Wen:2016:MPT

[WWSL16] Lian Wen, Kewen Wang, Yi-Dong Shen, and Fangzhen Lin. A model for phase transition of random answer-set programs. ACM Transactions on Computational Logic, 17(3):22:1–22:??, July 2016. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Wang:2015:DLO

[WWT15] Zhe Wang, Kewen Wang, and Rodney Topor. DL-Lite ontology revision based on an alternative semantic characterization. ACM Transactions on Computational Logic, 16(4): 31:1–31:??, November 2015. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Wang:2005:CCW

[WZ05] Kewen Wang and Lizhu Zhou. Comparisons and computation of well-founded semantics for disjunctive logic programs. ACM Transactions on Computational Logic, 6(2): 295–327, April 2005. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Wolter:2008:UUA

[WZ08] Frank Wolter and Michael Zakharyaschev. Undecidability of the unification and admissibility problems for modal and description logics. ACM Transactions on Computational Logic, 9(4):25:1– 25:??, August 2008. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Ying:2009:AQP

[YFDJ09] Mingsheng Ying, Yuan Feng, Runyao Duan, and Zhengfeng Ji. An algebra of quantum processes. ACM Transactions on Computational Logic, 10(3):19:1–19:??, April 2009. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Ying:2014:MCL

[YLYF14] Mingsheng Ying, Yangjia Li, Nengkun Yu, and Yuan Feng. Model-checking linear-time properties of quantum systems. ACM Transactions on Computational Logic, 15(3): 22:1–22:??, September 2014. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Yorsh:2007:LCH

[YRSW07] Greta Yorsh, Thomas Reps, Mooly Sagiv, and Reinhard Wilhelm. Logical characterizations of heap abstractions. ACM Transactions on Computational Logic, 8(1):??, January 2007. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Yorke-Smith:2009:CCR

[YSG09] Neil Yorke-Smith and Carmen Gervet. Certainty closure: Reliable constraint reasoning with incomplete or erroneous data. ACM Transactions on Computational Logic, 10(1): 3:1–3:??, January 2009. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Zanardini:2014:IFS

[ZG14] Damiano Zanardini and Samir Genaim. Inference of fieldsensitive reachability and cyclicity. ACM Transactions on Computational Logic, 15(4):33:1–33:??, September 2014. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Zhang:2006:LPB

[Zha06] Yan Zhang. Logic programbased updates. ACM Transactions on Computational Logic, 7(3):421–472, July 2006. CODEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Zhang:2014:RCB

[ZHD14] Lan Zhang, Ullrich Hustadt, and Clare Dixon. A resolution calculus for the branchingtime temporal logic CTL. ACM Transactions on Computational Logic, 15(1):10:1– 10:??, February 2014. CO-DEN ATCLA8. ISSN 1529-3785 (print), 1557-945X (electronic).

Zhang:2017:ELP

[ZZ17] Heng Zhang and Yan Zhang.

Expressiveness of logic programs under the general stable model semantics. ACM Transactions on Computational Logic, 18(2):9:1–9:??, June 2017. CODEN AT-CLA8. ISSN 1529-3785 (print), 1557-945X (electronic).